### v1

# Perseus: Removing Energy Bloat from Large Model Training

Abstract
Training large AI models on numerous GPUs consumes a massive amount of energy. We observe that not all energy consumed during training directly contributes to end-to-end training throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
In this work, we identify two independent sources of energy bloat in large model training, intrinsic and extrinsic, and propose Perseus, a unified optimization framework that mitigates both. Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient iterative graph cut-based algorithm and schedules energy consumption of its forward and backward computations across time to remove intrinsic and extrinsic energy bloat. Evaluation on large models like GPT-3 and Bloom shows that Perseus reduces energy consumption of large model training by up to 30%, enabling savings otherwise unobtainable before.
1 Introduction
As deep neural networks (DNNs) continue to grow in model and dataset size [29, 25], the energy consumption of large model training is increasing as well. For instance, training GPT-3 [11] reportedly consumed 1,287 megawatt-hour (MWh) in 2021 [48]. The recent surge in the development of generative AI (GenAI) and large language models (LLMs) shows no signs of deceleration in AI energy demands.
Optimizing the energy consumption of large model training is, therefore, crucial to minimizing the operational costs associated with AI, with additional benefits. For example, low-carbon electricity, which is preferred especially in the context of sustainability commitments [1, 4, 6, 7], is still very much a scarce and contended resource [50]. Reducing energy consumption allows more workloads to fit within the limited supply of low-carbon electricity [10]. Furthermore, reducing energy consumption without slowdown reduces power draw, either allowing more aggressive datacenter power oversubscription or conversely reducing power throttling for existing workloads [35, 54, 46].
Despite many recent works on increasing the throughput of large model training [26, 41, 19, 43, 73, 34, 27], reining in the growing energy consumption remains an open challenge [55, 48]. In this paper, we aim to identify and remove energy bloat – the portion of energy consumption that can be removed without throughput loss – in large model training. We identify two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2). Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning [41, 19, 23, 73], but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation. When all stages do not have the same computation intensity, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput. Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3). Root causes behind such slowdowns are varied, including power/thermal throttling [75, 47, 35, 54, 46], I/O bottlenecks in the storage/network [39, 71], hardware/software failures [61, 27], etc., and the likelihood of their presence increases with the scale and duration of training [22, 28, 64]. All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput. Extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
In this work, we propose Perseus that relies on a unified optimization framework to remove both intrinsic and extrinsic energy bloat during large model training (§3). At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, which enables it to minimize intrinsic energy bloat under normal operating conditions and to mitigate extrinsic energy bloat arising from straggler pipelines. Existing works fall short on both fronts. EnvPipe [13] is limited only to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction. Zeus [68], in contrast, ignores intrinsic bloat altogether as it only considers single-GPU training, which also leads it to identify suboptimal “iteration time–energy” frontiers at best.
We prove that optimally solving our optimization problem is in fact NP-hard and propose an efficient algorithm that provides an exact solution for a relaxed problem (§4). At its core, our solution represents an iteration of large model training as a directed acyclic graph (DAG). Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies. Importantly, each node is annotated with its planned execution time and energy consumption, which we call the energy schedule. Perseus’s optimization algorithm starts from the energy schedule that consumes the least possible amount of energy, and iteratively speeds up the execution time of the schedule while incurring the least possible energy increase, until no more speed up is possible. This generates all energy schedules on the “iteration time–energy” Pareto frontier. Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
While solving the optimization problem depends on accurate fine-grained computation energy profiling, modern GPUs provide energy consumption information only in coarse-grained time intervals. Perseus introduces an online profiling technique that can accurately profile the energy consumption of fine-grained computations with low overhead (§5).
Evaluation on a variety of large models including GPT-3 [11], BERT [15], T5 [51], Bloom [66] and a scaled-up version of Wide-ResNet [69] shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown (§6).
Overall, we make the following contributions in this paper:
  • We identify intrinsic and extrinsic energy bloat in large model training, fundamentally caused by computation time imbalance at different levels.
  • We propose Perseus, an open-source unified optimization framework that aims to remove both types of energy bloat using an efficient graph cut-based algorithm.
  • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat.
2 Motivation
First, we provide necessary background regarding large model training (§2.1). Then, we introduce intrinsic (§2.2) and extrinsic (§2.3) energy bloat present in large model training.
2.1 Large Model Training
The rapid increase in DNN model size in recent years made data, tensor, and pipeline parallelism [26, 41, 43, 19, 73, 34] essential ingredients for large model training. Especially, pipeline parallelism partitions a large model into multiple stages and assigns each stage to different GPUs. The training batch is split into multiple microbatches, and their forward and backward computations are pipelined through those stages. Then, in order to scale out to more GPUs, such pipelines are replicated multiple times to perform data parallel training, where each pipeline computes a subset of the training batch and synchronizes gradients with one other at the end of each iteration. Due to the synchronization barrier in the end, the next iteration can only begin when all pipelines have finished.
2.2 Intrinsic Energy Bloat
Figure 1: One training iteration of GPT-3 1.3B with 4 pipeline stages (partitioned with minimum imbalance ratio) and 6 microbatches on NVIDIA A100 GPUs drawn to scale. For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2. The critical path of computation is traced with a blue line. Colors are average power consumption. The blue background indicates that GPUs consume power when blocking on communication.
Model Size Imbalance Ratio
4 stages 8 stages
GPT-3 [11] 3B 1.13 1.25
7B 1.11 1.23
13B 1.08 1.17
175B 1.02 1.03
Bloom [66] 3B 1.13 1.25
7B 1.13 1.25
176B 1.05 1.10
BERT [15] 0.1B 1.33 2.00
0.3B 1.17 1.33
T5 [51] 0.2B 1.19 1.50
0.7B 1.05 1.11
2.9B 1.06 1.16
Wide-ResNet50 [69] 0.8B 1.23 1.46
Wide-ResNet101 [69] 1.5B 1.09 1.25
Table 1: Imbalance ratio between the longest and the shortest stages for various models on NVIDIA A100 GPUs. 1.00 would mean perfect balance.
We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 0(a). In addition to the familiar bubbles found in any pipeline schedule [41], we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage. Such gaps exist because the computation time of each pipeline stage is not perfectly balanced. Partitioning stages in a balanced manner is an important problem in distributed execution planning [26, 41, 19, 73], but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
To understand the amount of possible pipeline stage imbalance, we exhaustively enumerated all possible pipeline partitions [^1] For Transformer-based models, we partition at the granularity of Transformer layers. For Wide-ResNet, we partition at the granularity of bottleneck layers, which are three Convolution layers wrapped with a Skip connection. to find the one with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest. Table 1 lists the minimum imbalance ratio for various models, which shows that perfect balance is difficult to achieve. Appendix A explains the partitioning details of each model and where imbalance comes from.
Given some amount of stage imbalance, as we have seen in Figure 0(a), not all forward and backward computations are on the critical path of computation. This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy. We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation so that they do not affect the critical path (Figure 0(b)). Although seemingly simple, we prove that this problem is NP-hard by reduction from the 0/1 Knapsack problem (Appendix B). Visualizations for other models are in Appendix C.
2.3 Extrinsic Energy Bloat
During large model training, numerous replicas of the same pipeline run in a data parallel fashion in order to train on more data with more GPUs, and every pipeline must synchronize gradients at the end of each iteration before moving on to the next. When one pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes. Figure 1(a) illustrates this situation. Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time are wasteful. We call this extrinsic energy bloat, because unlike intrinsic energy bloat, the cause of energy bloat is extrinsic to the pipeline.
Perseus finds the energy-optimal iteration time for the non-straggler pipeline, again by precisely slowing down computations in the pipeline (Figure 1(b)). However, unlike other problem settings that do not consider energy consumption, we show in Section 3.1 that fully utilizing all the slack time created by the straggler may not be energy-optimal! Perseus can automatically make such decisions and implements the energy-optimal slowdown.
Figure 2: Toy example illustrating one straggler pipeline running together with a non-straggler pipeline using data parallelism. The non-straggler wastes energy by running as fast as possible. Perseus can precisely slow down non-straggler pipelines to reduce such extrinsic energy bloat.
Stragglers arise from numerous sources, with many introducing sizable energy bloat according to recent reports. Thermal or power throttling in a datacenter may lower the GPU’s frequency to prevent hardware or the power supply unit from being damaged, resulting in 10–50% slowdown [75, 47, 35, 54, 46]. I/O bottlenecks in the storage or network, especially during data loading for image, video, or recommendation models, can be longer than GPU computation by up to 4×4$\times$4 × [39, 71], thereby acting like a persistent (virtual) straggler pipeline. Recent failure-resilient large model training frameworks [61, 27] deploy heterogeneous pipelines with different number of stages or amount of compute after failures, introducing non-uniform iteration times across pipelines until recovery. As the scale of large model training jobs increase, the probability of the job encountering one or more causes of stragglers increases [22, 28, 64].
Stragglers due to events like thermal/power throttling, consistent I/O bottlenecks, and fault-tolerant planning can persist across many iterations [46, 39, 27]. A training system can identify such stragglers and calculate the extent of slowdown by comparing iteration times across data-parallel pipelines. In this work, we assume that such knowledge is available and focus on how to plan time and energy consumption across time and allow quick adaptation given such information.
3 Perseus Overview
Perseus is an energy optimization system for large model training. In this section, we present our unified optimization framework that aims to remove both types of energy bloat (§3.1) and walk through the workflow of Perseus (§3.2).
3.1 Optimization Approach
Figure 3: Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be. The first case is when there are no stragglers. $T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier. The black dot is the default mode of execution where all computations run at the maximum speed. The green dot is the energy-optimal iteration time of the non-straggler pipeline.
Intuitively, slowing down computations selectively in a training pipeline without affecting its critical path will keep the same iteration time while reducing its energy consumption (§2.2). Furthermore, when stragglers emerge, slowing down computations in a non-straggler pipeline without making it a straggler itself will reduce energy consumption even more (§2.3). We formalize these two intuitions into a unified optimization framework and derive a universal prescription for a non-straggler pipeline’s energy-optimal iteration time.
By controlling the execution speed of individual computations in the pipeline, our goal is to minimize the energy consumption of a pipeline, and we can safely slow down a pipeline’s iteration time until the straggler’s iteration time $T^{\prime}$:
$\min_{F}$ $\quad\textrm{Energy}(F)$ (1)
s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
where $F$ is the set of GPU frequencies to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively. Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane. However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier, a monotonically decreasing tradeoff curve.
Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t. the “iteration time–energy” Pareto frontier. $T_{min}$ is the shortest iteration time on the frontier, which coincides with the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption. Together, $T_{\min}$ and $T^{*}$ bookend the Pareto frontier.
  1. 1.
    When there are no stragglers (Figure 2(a)), we simply operate on the point on the Pareto frontier with the same iteration time, which reduces intrinsic energy bloat.
  2. 2.
    When a moderately slow straggler is detected (Figure 2(b)), we additionally reduce extrinsic energy bloat by slowing down non-straggler pipelines until $T^{\prime}$, using up all the slack time created by the straggler.
  3. 3.
    Finally, when the straggler’s iteration time goes beyond the minimum-energy point $T^{*}$ on the frontier (Figure 2(c)), we stop at $T^{*}$ instead of $T^{\prime}$. This is because slowing down beyond $T^{*}$ will instead increase energy.
The three cases can be merged into one universal prescription for the pipeline’s energy-optimal iteration time:
$T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$ (2)
Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier, we know the solution of Equation 1 for any feasible value of $T^{\prime}$. Then our system will be able to react quickly to emerging stragglers by simply looking up the set of frequencies $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$. We present our algorithm to efficiently characterize the entire “iteration time–energy” Pareto frontier in Section 4.
3.2 Perseus Architecture
Energy Schedule.
Perseus represents each iteration of the training pipeline as a static directed acyclic graph (DAG), where nodes are forward and backward computations and edges are dependencies in between computations. Each node on the computation DAG is annotated with its planned time and energy consumption, which we call the energy schedule. Perseus realizes an energy schedule by executing each computation with a specific GPU frequency.
System Components.
Figure 4 shows the high-level architecture of Perseus. Perseus is split into a framework-independent server and a framework-integrated client. The server runs our time–energy optimization algorithm, produces all energy schedules on the “iteration time–energy” Pareto frontier, and caches them in a table for fast lookup. The client profiles computations of the pipelines online during training, and realizes the energy schedule by changing the GPU’s frequency during runtime.
Figure 4: Perseus architecture and workflow.
Training Lifecycle Using Perseus.
In Perseus, a training job is specified by its computation DAG. When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy consumption of each forward and backward computation in each stage while training is running, and sends the profile to the server.
As training proceeds, 2 the server asynchronously characterizes the “iteration time–energy” Pareto frontier of the training job (§4), after which 3 the Pareto-optimal energy schedule corresponding to the shortest possible iteration time $T_{\min}$ is deployed to the client. Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework’s pipeline execution engine (§5). Pareto-optimal energy schedules are saved in a lookup table format index by $T^{\prime}$.
During training, 4 the hardware infrastructure (e.g., data center rack power manager) notifies the Perseus server of a straggler pipeline and how much slowdown it will experience. The Perseus server can 5 quickly react to this by looking up the Pareto-optimal energy schedule corresponding to that straggler iteration time, and deploying it to the clients.
4 Algorithm Design
In this section, we describe our algorithm to obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail. We start by formulating the problem (§4.1). Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3). Finally, we extend our algorithm to support tensor parallelism, single-choice operations, and diverse pipeline schedules (§4.4).
4.1 Problem Formulation
Expression for Energy Consumption.
The energy consumption of a pipeline is not only from computation; it is the sum of three parts: (1) Computation; (2) Blocking on communication between computations; and (3) Blocking on communication until the straggler pipeline finishes. Formally,
$\sum_{i}{e_{i}(f_{i})}+P_{\textrm{blocking}}(N\cdot T-\sum_{i}{t_{i}(f_{i})})+P_{\textrm{blocking}}\cdot N\cdot(T^{\prime}-T)$ (3)
[…]
where $P_{\textrm{blocking}}$ is the power consumption of the GPU when it is blocking on communication, $N$ is the number of pipeline stages, and $t_{i}(f_{i})$ and $e_{i}(f_{i})$ are the time and energy consumption of computation $i$ with frequency $f_{i}$, respectively.
Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline depends on the straggler’s iteration time $T^{\prime}$ in 2.222This makes sense because if the straggler slows down more and more, the non-straggler will consume more energy waiting for the straggler to finish. Specifically, the Pareto frontier for any $T^{\prime}$ is the Pareto frontier of $\textrm{Time}(F)$ vs. 1 shifted upwards by 2. Here, 1 does not depend on $T^{\prime}$. Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs. 1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$ as shifting the frontier upwards does not affect iteration time nor the frequency assignments that lead to each iteration time. Thus, we define
$\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$ (4)
and characterize the Pareto frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
Finding the Pareto Frontier.
Finding one point on the “iteration time–energy” Pareto frontier with iteration time $T^{\prime}$ is equivalent to solving the following optimization problem:
$\min_{F}$ $\quad\textrm{Energy}(F)$ (5)
s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
We call this problem Pipeline Energy Minimization (PEM).
Theorem 1.
Pipeline Energy Minimization is NP-hard.
Proof.
Reduction from 0/1 Knapsack. Details in Appendix B. ∎
The complete Pareto frontier can be obtained by solving PEM for all $T^{\prime}\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is in turn because the possible choices of GPU frequencies are discrete. However, if choices were continuous, the problem is exactly and efficiently solvable [57]. This is akin to integer linear programs becoming tractable when relaxed to linear programs.
We fit the exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements to make the choices continuous. We choose the exponential function due to its inherent flexibility and natural fit to data. We show in Section 6 that this relaxation produces high-quality approximate solutions that lead to actual savings. After relaxation, the problem’s optimization variables are the time and energy consumption planned for each computation in the pipeline, or the energy schedule.
4.2 Algorithm Overview
Iteratively Discovering the Pareto Frontier.
Figure 5: Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
While the relaxed Pipeline Energy Minimization problem is no longer NP-Hard, solving the problem for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient. Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate the next energy schedule adjacent to it on the Pareto frontier? Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, and every energy schedule we encounter will be Pareto-optimal.
[…]
We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) is incorrect. That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
TL;DR.
1
2
Input: DAG $\mathcal{G}$ of computations $i\in\mathcal{G}$
Amount of time to reduce in one iteration $\tau$
Iteration time with all max frequencies $T_{\min}$
Output: Set of Pareto-optimal energy schedules $\mathcal{P}$
3
4
$\triangleright$ Begin with the minimum energy schedule
5 $p\leftarrow$ Minimum energy for all computations $\mathcal{P}\leftarrow\{p\}$ while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
       $\triangleright$ Reduce time by $\tau$ with minimal energy increase (§4.3)
6       $p\leftarrow$ ReduceTime($\mathcal{G}$, $p$, $\tau$) $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
7 end while
$\triangleright$ Every schedule in $\mathcal{P}$ is Pareto-optimal
8 return $\mathcal{P}$
9
Algorithm 1 Iteratively Discovering the Pareto Frontier.
Algorithm 1 provides an overview of our optimization process. First, the energy schedule with the minimum energy consumption is constructed by planning every computation to run with minimum energy (line 1). Starting from there, the iteration time of the schedule is iteratively reduced by unit time $\tau$ while incurring minimal energy increase (line 1; Section 4.3). This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process is Pareto-optimal.
4.3 Reducing Time with Minimum Energy Increase
Figure 6: A simplified example of how to reduce iteration time by unit time $\tau$. Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2). Then the Critical DAG ( 3) is obtained by considering every and only the computations on the critical path. Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$. Cut A and Cut B are two examples of valid s-t cut ( 4). Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
In this section, we describe our core subroutine ReduceTime (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
Node- and Edge-Centric Computation DAGs.
Originally, Perseus’s representation of the computation DAG is node-centric, which has forward and backward computations as nodes and their dependencies as edges. As a setup for subsequent steps, we convert this into an edge-centric computation DAG where computations are edges and dependencies are nodes (i.e., all incoming edges must complete before any outgoing edge can begin). This conversion can be done by splitting each node on the original computation DAG into two nodes and connecting the two with an edge annotated with the computation on the original node.
Removing Non-Critical Computations.
Our goal is to reduce the execution time of the computation DAG by $\tau$, which is equivalent to reducing the length of all critical paths by $\tau$. Since computations that are not on any critical path (i.e., non-critical computations) do not affect the length of the critical path, we simply remove them from the computation DAG.
Finding Computations to Speed Up.
Which computations on the DAG should we speed up in order to reduce the length of all critical paths by $\tau$? The key observation is that any s-t cut on the computation DAG represents a way to reduce the execution time of the DAG by $\tau$. Specifically, by speeding up the computations on all cut edges by $\tau$, the entire computation DAG can be sped up exactly by $\tau$.
Figure 6 shows two examples of this. We have already 1 transformed an energy schedule into 2 a computation DAG where computations are represented by edges, and 3 removed all non-critical computations. Here, 4 shows two valid s-t cuts: Cut A and Cut B. 5A speeds up the Backward computation cut by Cut A from $3\tau$ to $2\tau$, and the iteration time of the energy schedule was reduced by $\tau$. Similarly, 5B speeds up the Forward computation and Backward computation defined by Cut B from $5\tau$ to $4\tau$ and from $7\tau$ to $6\tau$ respectively, and the iteration time of the energy schedule was also reduced by $\tau$. Especially, in the second case, iteration time was only reduced because computations on two parallel critical paths were sped up together.
Solving with Minimum Cut.
We have shown with examples that any s-t cut represents a way to reduce the duration of the computation DAG by $\tau$. Then, a natural question is, which cut brings the smallest possible energy increase?
We can precisely map the capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of the exponential function fit for that computation and defining it to be the edge’s flow capacity. Then, our problem reduces to minimum cut, which we can solve with maximum flow. Appendix D has further details.
Converting to GPU Frequencies.
After finding the minimum capacity (minimum energy increase) cut, we will modify the durations of the computations involved in the cut, which results in a new energy schedule. Finally, we will convert the energy schedule into GPU frequencies that can be realized by the Perseus client. For each computation $i$, we convert its execution time $t$ to the slowest GPU frequency that will execute faster than $t$. This is because when computations are tightly packed by our algorithm, while slightly speeding up a computation is acceptable, slowing down any computation on the critical path will directly slow down the entire DAG, increasing intrinsic energy bloat.
Time Complexity Analysis.
Our optimization algorithm has polynomial runtime. Let $N$ and $M$ respectively denote the number of stages and microbatches. Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$. While for general DAGs the total number of steps is known to be exponential to the size of the DAG [58], we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$. See Appendix E for proof.
In reality, commonly used number of stages ($N$) are 4 to 8 as too many increase pipeline bubble ratio [43, 14]. The number of microbatches ($M$) is typically around $4N$ [26, 59], but recently with high data parallel degree, far less have been reported even for high-performance settings [14]. As such, the algorithm runtime is negligible in practical scenarios (§6.5), given that large model training time is typically several weeks [48].
1
2
Input: DAG $\mathcal{G}$ with computations $i$
Current energy schedule $p$
Amount of iteration time to reduce $\tau$
Output: Pareto-optimal schedule with reduced time $p^{\prime}$
3
4
Function $\mathrm{ReduceTime}$($\mathcal{G}$, $p$, $\tau$):
       $\triangleright$ Construct edge-centric computation DAG ( 2)
       $\mathcal{G}\leftarrow$ Split nodes into two and connect with edge $\mathcal{G}\leftarrow$ Crash bipartite node set to one node $\triangleright$ Remove non-critical computations ( 3)
5       Annotate earliest & latest start times for $\forall i\in\mathcal{G}$ for $i\in\mathcal{G}$ do
             $\triangleright$ Critical computations should have zero slack
6             if $i$ has different earliest and latest start then
7                   Remove $i$ from $\mathcal{G}$
8             end if
9
10       end for
      $\triangleright$ Find set of computations to modify ( 4)
       $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$) $\triangleright$ Modify computation durations ( 5)
       $I\leftarrow\{i\mid i\textrm{ in }S-T\textrm{ cut}\}$ Modify duration of $\forall i\in I$ by $\tau$triangleright Assign frequencies from planned computation times
11       $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned return $p^{\prime}$
12
13
14
Algorithm 2 ReduceTime: Reducing the execution time of a computation DAG by $\tau$ with minimal energy increase.
4.4 Extensions
In this section, we present extensions to our optimization algorithm useful for planning large model training.
Tensor Parallelism.
It is trivial to extend our algorithm to tensor parallelism, another essential ingredient of large model training. The observation is that tensor parallel techniques split operations in equal sizes. Therefore, GPUs that execute different portions of the same operation consume the same amount of time and energy. This allows Perseus to operate only on one tensor parallelism copy for each stage, decide the energy schedule for that copy, and replicate it to other copies in the same stage. We show that Perseus works well for hybrid parallelism in Section 6.4.
Single-Choice Operations.
Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latencies. For instance, especially on fast GPUs like A100, loading and copying input data into VRAM or communication over slower links can take considerable latency. However, the time and energy consumption of these operations are not affected by the GPU’s frequency. Perseus can take the latency of such operations into account during planning by abstracting them as a node in the DAG with only one frequency choice.
Other Pipeline Schedules.
There are various schedules for pipeline parallel training, including GPipe [26], 1F1B [42], interleaved 1F1B [56], and early recomputation 1F1B [34]. As long as the computations on the schedule can be expressed as a DAG, Perseus can optimize its energy consumption without modification. If there is stage imbalance, we believe that any pipeline schedule will have intrinsic energy bloat.
5 Implementation
The Perseus server is implemented in 1,500 lines of Python [3]. The client is implemented in 300 lines of Python as a library that can be imported into training frameworks.
As a reference, we have integrated the Perseus client with Merak [34], whose open source implementation takes the best of Megatron-LM [5] for high-performance tensor parallelism for transformer models and DeepSpeed [2] for its generic pipeline execution engine. Activation recomputation [12] is enabled by default to allow large batch sizes to fit in GPUs.
Online Time/Energy Profiler.
While accurate timing measurement is well-supported, modern GPUs (NVIDIA A40, A100, and H100) update their cumulative energy consumption counter [8] only once every 100 ms, which is similar to the timescale of one microbatch computation. This makes energy measurement noisy, and no ground truth is available.
At the very moment the cumulative energy counter increases, its value should be accurate. Therefore, Perseus polls the energy counter in the background and records the timestamp of when the counter increased and its value. Then, collected (time, energy) points are interpolated with a linear line. The main training process only records the start and end timing of each computation, allowing Perseus to match timestamps with the interpolated energy curve to calculate each computation’s energy consumption.
The GPU’s frequency is scanned from the highest to the lowest at iteration granularity. As an optimization, when computation energy increases for five consecutive frequencies, profiling is terminated. Beyond that point, frequencies will consume more time and energy, making them suboptimal.
Asynchronous Frequency Controller.
The controller lives in a separate process that communicates with the main training process through a pipe, because setting the GPU’s frequency through NVML [8] should not block computation on the main process. While pipeline execution engine implementations differ widely, many have separate code blocks for forward and backward. Therefore, the controller exposes one method, set_frequency, which takes either "forward" or "backward" and sets the GPU’s frequency as planned.
6 Evaluation
We evaluate Perseus on five workloads and compare it against EnvPipe and Zeus. Our key findings are the following.
  • Perseus can effectively reduce intrinsic and extrinsic energy bloat. Training on real GPUs shows up to 28.5% energy savings using Perseus (§6.2).
  • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings. We also observe that there is a tradeoff between energy savings and scale (§6.3).
  • Energy bloat reduction was possible because Perseus can enumerate efficient energy schedules on the “iteration time–energy” Pareto frontier (§6.4).
  • Perseus reduces energy bloat with low overhead (§6.5).
6.1 Experimental Setup
Testbed.
We run our evaluation workloads in a GPU cluster, where each node is equipped with an AMD EPYC 7513 CPU, 512 GB DRAM, and four NVIDIA A40-48G GPUs. For A100 results, we use the node provided by Chameleon Cloud [30], equipped with two Intel Xeon Platinum 8380 CPUs, 512 GB DRAM, and four NVIDIA A100-80G PCIe GPUs.
Workloads and experiment parameters.
We evaluate Perseus with various workloads spanning from GPT-3 [11], Bloom [66], BERT [15], T5 [51], to Wide-ResNet [69]. We use model variants with 1.3B to 6.7B parameters to run the models in our testbed, and scale them up to 176B parameters in large-scale emulation. We chose the microbatch size and number of microbatches that yield the highest throughput given the global batch size. We use the minimum imbalance stage partitioning method described in Section 2.2 for all workloads. Appendix A lists complete model configurations, parameters, and stage partitioning details.
Baselines.
We mainly compare against two prior works:
  • EnvPipe [13] reduces only intrinsic energy bloat while trying to minimize slowdown in pipeline execution. We compare Perseus’s amount of energy bloat reduction against that of EnvPipe (§6.2, §6.3).
  • Zeus [68] characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
6.2 Reducing Energy Bloat
We start with overall energy bloat reduction – both intrinsic (§6.2.1) and extrinsic (§6.2.2) – achieved by Perseus and EnvPipe by running various models on our testbed. We compare Perseus against EnvPipe, which reduces energy consumption while trying to minimize slowdown. Both solutions use the same amount of GPU resources.
6.2.1 Intrinsic Energy Bloat Reduction
Model Energy Savings (%) Slowdown (%)
Perseus EnvPipe Perseus EnvPipe
GPT-3 1.3B 13.2 8.8 0.1 0.1
BERT 1.3B 12.9 8.0 0.5 0.0
T5 3B 10.6 7.4 1.3 3.4
Bloom 3B 11.7 8.9 0.2 0.2
Wide-ResNet 1.5B 3.2 3.7 2.3 4.1
[…]
Model Energy Savings (%) Slowdown (%)
Perseus EnvPipe Perseus EnvPipe
GPT-3 2.7B 21.1 21.7 0.2 5.6
BERT 1.3B 15.7 16.5 0.0 9.7
T5 3B 28.5 19.3 0.0 0.0
Bloom 3B 22.4 19.9 0.0 0.0
Wide-ResNet 1.5B 20.4 16.5 0.2 0.5
[…]
Table 2: Intrinsic energy reduction and slowdown comparison between Perseus and EnvPipe.
Table 2 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the “iteration time–energy” frontier) and that by EnvPipe.
We make two observations regarding Perseus. First, models show varying amounts of energy savings because their stage imbalance vary (Table 1). For instance, unlike other models, Wide-ResNet 1.5B on A100 after minimum imbalance stage partitioning has nearly perfect stage balance, leaving little room for non-critical computations to slow down.
Second, A40 demonstrates more energy savings compared to A100. This is because the dynamic clock frequency range of A100 (210–1410 MHz) is smaller than that of A40 (210–1740 MHz). Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency. However, we expect the upcoming NVIDIA H100 GPU to have better savings because its maximum frequency is 1755 MHz for the PCIe version and 1980 MHz for SXM [44].
EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest, which is not always true. Additionally, it sometimes considerably degrades iteration time because it is not aware of single-choice operations inside the pipeline (§4.4) and can slow down some computations too much.
6.2.2 Intrinsic + Extrinsic Energy Bloat Reduction
Model Method Energy Savings (%) given $T^{\prime}/T$
# Params 1.05 1.1 1.2 1.3 1.4 1.5
GPT-3 Perseus 14.7 15.9 15.5 15.0 14.6 14.3
1.3B EnvPipe 8.7 8.5 8.3 8.1 7.9 7.7
Bloom Perseus 13.6 15.6 15.2 14.7 14.3 14.0
3B EnvPipe 8.8 8.7 8.4 8.2 8.0 7.8
BERT Perseus 14.9 16.9 16.4 15.9 15.5 15.0
1.3B EnvPipe 7.9 7.8 7.5 7.3 7.1 6.9
T5 Perseus 15.3 18.0 17.9 17.4 16.9 16.5
3B EnvPipe 8.4 8.2 8.0 7.8 7.6 7.4
Wide-ResNet Perseus 9.4 12.7 12.6 12.3 12.0 11.6
1.5B EnvPipe 4.9 4.8 4.7 4.5 4.4 4.3
[…]
Model Method Energy Savings (%) given $T^{\prime}/T$
# Params 1.05 1.1 1.2 1.3 1.4 1.5
GPT-3 Perseus 24.5 26.0 25.9 25.2 24.6 24.0
2.7B EnvPipe 22.9 22.6 22.0 21.4 20.9 20.4
Bloom Perseus 25.5 26.4 25.9 25.2 24.6 24.0
3B EnvPipe 19.6 19.3 18.8 18.3 17.8 17.4
BERT Perseus 20.0 22.6 24.1 23.4 22.8 22.2
1.3B EnvPipe 19.2 18.9 18.3 17.8 17.4 16.9
T5 Perseus 27.9 27.3 26.2 25.2 24.3 23.4
3B EnvPipe 18.4 18.0 17.3 16.6 16.0 15.4
Wide-ResNet Perseus 24.3 26.2 26.3 25.7 25.0 24.4
1.5B EnvPipe 16.4 16.2 15.8 15.4 15.0 14.6
[…]
Table 3: Energy savings of Perseus given varying degrees of straggler slowdown ($T^{\prime}/T$). EnvPipe cannot provide additional savings because it does not consider and cannot be applied to optimize extrinsic bloat.
When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its time–energy frontier yields for longer iteration times. Table 3 shows the amount of energy savings given varying degrees of straggler slowdown. As described in Section 3.1, non-stragglers’ iteration time is set to be $\min(T^{*},T^{\prime})$, and energy is reduced from (1) slowing down the pipeline itself and (2) blocking on communication for a shorter amount of time, waiting for the straggler. Slowing down the non-stragglers beyond $T^{*}$ increases total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
The percentage of energy savings increases for lower $T^{\prime}$ values and slowly wanes when $T^{\prime}>T^{*}$. This reduction is due to longer blocking time. That is, the absolute amount of energy savings in Joules is the largest at $T^{*}$, and constant afterward. However, energy consumption during blocking time ($T^{\prime}-T^{*}$) increases for larger $T^{\prime}$ values, lowering the percentage of energy savings for $T^{\prime}>T^{*}$.
Finally, the point of maximum energy savings is different for each model. This is because different models have different $T^{*}$ values, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
6.3 Large-Scale Emulation
Because we do not have access to a GPU cluster required to run huge models like GPT-3 175B, we use emulation grounded on fine-grained profiling for large-scale evaluation.
Emulation Methodology.
Total
# GPUs
# Pipelines
# Microbatches
per Pipelines
Global
Batch Size
1024 16 96 1536
2048 32 48
4096 64 24
8192 128 12
Table 4: Experiment parameters for GPT-3 175B and Bloom 176B. Each pipeline has tensor parallel degree 8 and 8 pipeline stages.
We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B in bfloat16 to construct the time and energy profile of each stage. Then, we run our optimization algorithm to obtain a theoretical “iteration time–energy” frontier, and use it to report emulated savings. We evaluate the changes in the amount of energy savings in a strong scaling setup presented in Table 4. We do not consider weak scaling, where per-pipeline batch is constant and increasing the global batch size proportionally, because varying the global batch size can affect the model quality [31, 21]. We find that, compared to actual measurements of smaller scale models on A100, the emulator always underestimates the actual percentage of energy savings by 20.2% on average (see Appendix F for details). Meaning, the savings reported by our emulation can likely be considered an empirical lower bound for actual savings. We used A100 SXM GPUs for the A100 results in this section, which is more representative of large scale scenarios (Appendix G provides PCIe results).
Results Summary.
Figure 7: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1,024 GPUs.
Figure 7 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1,024 GPUs as a representative. EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it reduces less than Perseus. In contrast, Perseus reduces per-iteration energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
Intrinsic Bloat Reduction.
Model GPU Type
Energy Savings (%)
given # Microbatches
12 24 48 96
GPT-3 175B A100 15.20 14.19 13.62 13.32
A40 11.81 10.22 9.34 8.88
Bloom 176B A100 10.47 7.06 5.23 4.28
A40 6.97 4.49 3.12 2.41
Table 5: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
Table 5 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches. For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases. This is fundamentally due to the ratio of microbatches in the pipeline’s warm-up and flush phase (beginning and end) vs. steady state phase (middle). Microbatches in the former phase are able to slow down until their minimum energy frequency, yielding large energy savings. However, microbatches in the latter (middle of the pipeline) cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings. When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction becomes more and more dominated by the average energy savings of steady state microbatches.
Extrinsic Bloat Reduction.
Figure 8: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines. When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$, whose location is denoted by a star. More blocking time after $T^{*}$ decreases overall energy savings. Please note the different Y-axes.
We simulate stragglers in training large models with hybrid parallelism in various strong scaling configurations (Table 4), where a straggler is slower than others in varying degrees from 1.05 to 1.50. Other non-stragglers, after finishing their computation, must wait until the straggler finishes computation, where extrinsic energy bloat emerges. We measure the amount of energy bloat and Perseus’ extrinsic energy savings.
Figure 8 shows the changes of extrinsic energy bloat reduction given varying degrees of straggler slowdown and various strong scale configurations. The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward, is consistent with what was observed in Section 6.2.2.
An interesting observation here is that there is a tradeoff between scale and energy savings: more pipelines have less percentage of energy savings or less amount of energy savings per pipeline. It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline that cannot be optimized while all the other pipelines optimize their energy consumption. However, this holds only in weak-scaling configuration, i.e., per-pipeline batch size is constant (increasing the global batch size proportionally to the number of pipelines) and it is not the case in strong-scaling configuration, where the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed. With less number of microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration, which cannot be eliminated by intrinsic energy bloat reduction, increases, resulting in less energy savings.
6.4 Iteration Time–Energy Frontier Comparison
Figure 9: Iteration time–energy frontiers for GPT-3, achieved by Perseus and the two baselines derived from Zeus [68]. Perseus Pareto-dominates all other approaches. The dotted vertical line is the iteration time of running all GPUs at their maximum power limit, which is the default mode of operation. Please note the different Y-axes.
The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm. Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus [68], which is an energy optimization framework for a single-GPU training with the training time–energy frontier. We implemented two Zeus-based baselines to make it work in parallelization configurations and generated “iteration time–energy” frontiers.
  1. 1.
    ZeusGlobal: Scans one global power limit for all stages.
  2. 2.
    ZeusPerStage: Sets one power limit per stage that balances forward computation time.
We run training and measurement under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) hybrid parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40. Figure 9 shows the frontiers of Perseus and Zeus for different sizes of GPT-3 under three parallelization configurations. Frontiers for other models are in Appendix H.
Perseus Pareto-dominates Zeus. ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage equally, unable to reduce intrinsic energy bloat. While ZeusPerStage can balance the forward computation time of each stage, it is unaware of the critical path of the DAG, slowing down critical computations. In contrast, Perseus can precisely slow down non-critical computations, tightly packing computation.
6.5 Overhead of Perseus
Profiling.
Our online profiling method (§5) introduces extra training time by running some iterations at a lower frequency. For our A100 workloads, the average percentage of slowdown for such iterations was 8.2%, and the average extra training time introduced by profiling was 13 minutes. This is negligible compared to how long large model training can take.
Algorithm Runtime.
The average time spent on running Perseus’s optimization algorithm (§4) across the five A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes). For our largest scale emulation experiment (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds. While the runtime of the algorithm will increase with larger DAGs for larger models, we believe the overhead be justified because training time also increases with the scale of the training job. Finding the energy-optimal iteration time and energy schedule given the straggler’s iteration time for extrinsic energy bloat reduction is instant.
7 Related Works
Large Model Training.
Many recent works focus on enabling and accelerating large model training using 3D parallelism (data, tensor, and pipeline parallelism). GPipe [26] and PipeDream [41] were the first to introduce pipeline parallelism and explicitly discussed the difficulty of perfectly balancing computation time across stages. Megatron-LM [56, 43] is a Transformer-based large model training framework that provides efficient manually designed execution plans. Later on, modern training frameworks that focus on large scale training [53, 59, 36, 37, 40, 34] were introduced to support various models at scale. DeepSpeed [53] introduced ZeRO redundancy optimizer that shards model states, which is recently widely used for large model training [52, 72]. Alpa [73] and GSPMD[70] are automatic parallelization framework for general DNNs. Finally, some recent works have looked into fault-tolerant large model training as well [61, 27, 9]. Unfortunately, energy consumption is not an optimization metric for any of the major large model training frameworks.
DNN Training and Energy Consumption.
A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training, including those that present observations and estimations [48, 60, 33, 16, 38] and those that propose optimization methods for training time, energy consumption, and carbon footprint [74, 63, 68, 67, 32, 13].
Zeus [68] is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on simple single-GPU training. EnvPipe [13], on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown. However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings and an infinite loop during optimization. Perseus Pareto-dominates both Zeus and EnvPipe by viewing large model training as a computation DAG and introducing a principled optimization algorithm. Perseus is also the first to introduce the notion of extrinsic energy bloat in large model training and optimize both simultaneously.
8 Conclusion
We presented Perseus, an energy optimization system for large model training. Perseus builds on top of observation that there are fundamental computation imbalance at different levels in large model training that causes intrinsic and extrinsic energy bloat. We introduced a principled graph cut-based algorithm that simultaneously reduces both.
Perseus advances the state-of-the-art of DNN training energy optimization by establishing a new time–energy Pareto frontier for large model training. The integration of Perseus into the training workflow has strong implications for the future of AI development. Importantly, reducing energy bloat leads to practically no latency and throughput degradation, which has the potential to greatly enhance the sustainability of distributed training in the midst of recent proliferation of LLMs and GenAI.
Acknowledgements
We thank Chameleon Cloud [30] for providing A100 nodes.
Appendix A Workload Details
A.1 Minimum Imbalance Pipeline Partitioning
Model Size Imbalance Ratio Minimum Imbalance Ratio Partition
4 stages 8 stages 4 stages 8 stages
GPT-3 [11] 1B 1.17 1.33 [0, 6, 12, 19, 25] [0, 4, 7, 10, 13, 16, 19, 22, 25]
3B 1.13 1.25 [0, 8, 16, 25, 33] [0, 5, 9, 13, 17, 21, 25, 29, 33]
7B 1.11 1.23 [0, 8, 16, 24, 33] [0, 4, 8, 12, 16, 20, 24, 28, 33]
13B 1.08 1.17 [0, 10, 20, 30, 41] [0, 5, 10, 15, 20, 25, 30, 35, 41]
175B 1.02 1.03 [0, 24, 48, 72, 97] [0, 12, 24, 36, 48, 60, 72, 84, 97]
Bloom [66] 3B 1.13 1.25 [0, 9, 17, 25, 31] [0, 6, 11, 16, 22, 28, 31]
7B 1.13 1.25 [0, 9, 17, 25, 31] [0, 6, 11, 16, 22, 28, 31]
176B 1.05 1.10 [0, 18, 36, 54, 71] [0, 9, 18, 27, 36, 45, 54, 63, 71]
BERT [15] 0.1B 1.33 2.00 [0, 4, 7, 10, 13] [0, 2, 3, 4, 6, 8, 10, 12, 13]
0.3B 1.17 1.33 [0, 7, 13, 19, 25] [0, 3, 6, 9, 12, 15, 18, 22, 25]
1.3B 1.17 1.33 [0, 7, 13, 19, 25] [0, 4, 7, 10, 13, 16, 19, 22, 25]
T5 [51] 0.2B 1.19 1.50 [0, 9, 15, 20, 25] [0, 5, 9, 13, 15, 17, 19, 22, 25]
0.7B 1.05 1.11 [0, 16, 29, 39, 49] [0, 8, 16, 24, 29, 34, 39, 44, 49]
2.9B 1.06 1.16 [0, 15, 28, 38, 49] [0, 7, 15, 23, 28, 33, 38, 43, 49]
Wide-ResNet50 [69] 0.8B 1.23 1.46 [0, 5, 9, 14, 18] [0, 3, 5, 7, 9, 11, 13, 15, 18]
Wide-ResNet101 [69] 1.5B 1.09 1.25 [0, 8, 17, 26, 35] [0, 4, 8, 12, 16, 21, 26, 31, 35]
[…]
Model Size Imbalance Ratio Minimum Imbalance Ratio Partition
4 stages 8 stages 4 stages 8 stages
GPT-3 [11] 1B 1.15 1.31 [0, 6, 12, 18, 25] [0, 3, 6, 9, 12, 15, 18, 21, 25]
3B 1.11 1.21 [0, 8, 16, 24, 33] [0, 4, 8, 12, 16, 20, 24, 28, 33]
7B 1.08 1.17 [0, 8, 16, 24, 33] [0, 4, 8, 12, 16, 20, 24, 28, 33]
13B 1.07 1.14 [0, 10, 20, 30, 41] [0, 5, 10, 15, 20, 25, 30, 35, 41]
175B 1.01 1.02 [0, 24, 48, 72, 97] [0, 12, 24, 36, 48, 60, 72, 84, 97]
Bloom [66] 3B 1.13 1.25 [0, 9, 17, 25, 31] [0, 5, 9, 13, 17, 21, 25, 29, 31]
7B 1.13 1.25 [0, 9, 17, 25, 31] [0, 5, 9, 13, 17, 21, 25, 29, 31]
176B 1.03 1.06 [0, 18, 36, 54, 71] [0, 9, 18, 27, 36, 45, 54, 63, 71]
BERT [15] 0.1B 1.33 2.00 [0, 4, 7, 10, 13] [0, 1, 2, 4, 6, 8, 10, 12, 13]
0.3B 1.17 1.33 [0, 7, 13, 19, 25] [0, 4, 7, 10, 13, 16, 19, 22, 25]
1.3B 1.17 1.33 [0, 7, 13, 19, 25] [0, 3, 6, 9, 12, 15, 18, 22, 25]
T5 [51] 0.2B 1.20 1.50 [0, 9, 15, 20, 25] [0, 5, 9, 13, 15, 17, 19, 22, 25]
0.7B 1.06 1.12 [0, 16, 29, 39, 49] [0, 8, 16, 24, 29, 34, 39, 44, 49]
2.9B 1.07 1.17 [0, 15, 28, 38, 49] [0, 8, 15, 23, 28, 33, 38, 43, 49]
Wide-ResNet50 [69] 0.8B 1.13 1.72 [0, 5, 9, 14, 18] [0, 3, 5, 7, 9, 11, 13, 15, 18]
Wide-ResNet101 [69] 1.5B 1.08 1.25 [0, 8, 17, 26, 35] [0, 4, 8, 12, 16, 21, 26, 31, 35]
[…]
Table 6: Imbalance ratio between the longest and the shortest stages for various models. 1.00 would mean perfect balance. Partitions for $N$ stages is a list of $N+1$ numbers, where the numbers represent layer indices. For instance, [0, 6, 12, 19, 25] for GPT-3 1.3B means there are 6, 6, 7, and 5 Transformer layers in each stage, and the final stage also has the language model head.
We partition layers of a model into $N$ stages such that the imbalance ratio, defined as the ratio of the longest stage forward latency to the shortest, is minimized. We only consider forward computation time as backward computations are typically proportional to forward computation latency. For Transformer-based models, we define layer as one Transformer layer. For Wide-ResNet, we define layer as one Bottleneck layer, which is three convolution layers wrapped with a skip connection. Due to P2P communication overhead and implementation challenges, many planners and frameworks do not support partitioning in the middle of skip connections. We call this minimum imbalance pipeline partitioning, and throughout the paper, every workload we use is partitioned as such.
Table 6 shows the computation time ratios of the heaviest stage to the lightest stage for 4 and 8 pipeline stages. More pipeline stages generally increases imbalance due to the coarse-grained nature of tensor operations. That is, the relative size of each layer becomes smaller and smaller compared to the total amount of computation allocated to each stage, and imbalance increases.
GPT-3, Bloom, and BERT.
Arguably, these are one of the most homogeneous large models because they are a stack of identical Transformer [62] encoder or decoder layers. However, the very last layer is the language modeling head, which maps the final features to probabilities over the entire vocabulary. The vocabulary size of GPT-3 is 50k, Bloom 251k, and BERT 31k, which results in a very large linear layer for the last stage. This leads to varying amounts of imbalance and different minimum imbalance partitioning results for each models.
T5.
This is also based on Transformer layers, but the first half of the layers are encoders, while the later half are decoders (which corresponds to the original Transformer [62] model’s architecture). However, the decoder layers as an extra cross attention layer in the middle, making it computationally heavier. Finally, T5 also ends with a language model head with 32k vocabulary size. However, minimum imbalance partitioning still balances T5 to a reasonable degree, although it cannot be perfectly balanced.
Wide-ResNet.
For Wide-ResNet, in order to make it suitable for large model training evaluation, we used the variant with width factor 8. Wide-ResNet is a collection of Bottleneck layers with three convolutions wrapped with a skip connection, and there are four different sizes of Bottleneck layers laid out sequentially. Therefore, even with minimum imbalance partitioning, it is difficult to perfectly balance stages.
A.2 Experiment Parameters
Model # Parameters Global Batch Size Microbatch Size # Microbatches
gpt3-6.7b 6.7 B 1024 4 128
Table 7: Experiment Parameters for 3D parallelism experiments on A40 GPUs. Microbatch size is per-pipeline, and there are two data parallel copies of the same pipeline. Thus, global batch size should be calculated as the product of microbatch size and the number of microbatches times two.
Model # Parameters Global Batch Size Microbatch Size # Microbatches
gpt3-2.7b 2.7 B 1024 4 256
bert-huge-uncased 1.3 B 256 8 32
t5-3b 3.0 B 128 4 32
bloom-3b 3.0 B 512 4 128
wide-resnet101 (width factor 8) 1.5 B 1536 32 48
Table 8: Experiment Parameters for eight-stage pipeline parallelism experiments on A40 GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
Model # Parameters Global Batch Size Microbatch Size # Microbatches
gpt3-xl 1.3 B 512 4 128
bert-huge-uncased 1.3 B 256 8 32
t5-3b 3.0 B 128 4 32
bloom-3b 3.0 B 512 4 128
wide-resnet101 (width factor 8) 1.5 B 1536 64 24
Table 9: Experiment Parameters for Pipeline Parallelism Experiments on A100 PCIe GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
Tables 78, and 9 list model variant names and experiment parameters for our experiments. Model names and configurations for GPT-3 were taken as is from the original model publication [11]. Especially, model names and configurations for BERT and T5 were directly taken from the Huggingface Hub pretrained model zoo, except for the huge variant of BERT, which we created to have hidden dimension 2048. Wide-ResNet was based on Torch Vision [45] but scaled up following the model’s original publication [69] using its width factor parameter. The unit time parameter $\tau$ was set to 1 ms for all experiments.
Appendix B Pipeline Energy Minimization is NP-Hard
The original Pipeline Energy Minimization problem is stated in formal terms:
$\min_{F}\quad$ $\textrm{Energy}(F)$ (6)
s.t. $\textrm{Time}(F)\leq T^{\prime}$
where $F$ is the frequency assignment for each pipeline computation $i\in\mathcal{G}$, $T^{\prime}$ is the straggler pipeline’s iteration time. The decision problem corresponding to Equation 6 asks whether it is possible to find frequency assignment $F$ such that the total energy consumption is minimized while the iteration time of pipeline $\mathcal{G}$ is no longer than the straggler’s iteration time. We denote this problem pem.
In the following, we show that a simplification of pem is NP-hard by reduction from the 0/1 Knapsack problem, which makes pem NP-Hard.
B.1 One Stage Two Frequencies Simplification
A simplification of pem is considering the case where there is only one pipeline stage and two frequencies to choose from.
For each pipeline computation $i\in\{1,2,\ldots,n\}$, we can set the GPU frequency to either the lowest value or the highest, denoted as $[L,H]$ respectively. Choosing different frequencies will lead to different execution time and energy consumption. That is, if $i$ is chosen to execute at frequency $L$, it will take $t_{i}(L)$ time and $e_{i}(L)$ energy. On the other hand, if $i$ executes at frequency $H$, it takes $t_{i}(H)$ time and $e_{i}(H)$ energy. The time and energy consumption of $i$ are rational numbers, as they are rounded up to $\tau$.
Our goal is to minimize the energy consumption of executing all computations while satisfying the time constraint. Specifically, given a time deadline $T^{\prime}$, we want to pick a subset of operations $J\subseteq\{1,2,\ldots,n\}$ and assign them to execute at the lowest frequency $L$ and execute the rest of the operations with the highest frequency $H$, such that the total time needed to execute all operations is smaller than or equal to the deadline:
$\sum^{n}_{i=1}(X_{i}t_{i}(L)+(1-X_{i})t_{i}(H))\leq T^{\prime}$
where $X_{i}$ is a 0/1 indicator variable where $X_{i}=1$ if $i\in J$ and $X_{i}=0$ otherwise. Under this time constraint, the goal is to minimize the total energy consumption of executing all computations:
$\sum^{n}_{i=1}(X_{i}e_{i}(L)+(1-X_{i})e_{i}(H)).$
Formally, we denote this problem as
[…]
where $T_{L}=[t_{1}(L),\ldots,t_{n}(L)]$, $T_{H}=[t_{1}(H),\ldots,t_{n}(H)]$ are execution time vectors for low and high frequency respectively, $E_{L}=[e_{1}(L),\ldots,e_{n}(L)]$, $E_{H}=[e_{1}(H),\ldots,e_{n}(H)]$ are energy consumption vectors for low and high frequency respectively, and $EC$ the target energy consumption.
pem-1d returns true if and only if there exists a subset of operations $J\subseteq\{1,2,\ldots,n\}$ such that $\sum^{n}_{i=1}(X_{i}t_{i}(L)+(1-X_{i})t_{i}(H))\leq T^{\prime}$ and $\sum^{n}_{i=1}(X_{i}e_{i}(L)+(1-X_{i})e_{i}(H))\leq EC$.
B.2 0/1 Knapsack Problem
Consider two length $n$ arrays containing positive integer weights $W=(w_{1},w_{2},\ldots,w_{n})$ and values $V=(v_{1},v_{2},\ldots,v_{n})$ where the $i$th item has weight $w_{i}\in\mathbb{Q}^{+}$ and value $v_{i}\in\mathbb{Q}^{+}$, and a knapsack with weight capacity $C\in\mathbb{Q}^{+}$.
The goal is to pick a subset of items $S\subseteq\{1,2,\ldots,n\}$, such that the total weight of the chosen items is less than or equal to the weight capacity: $\sum_{i\in S}w_{i}\leq C$. Under this constraint, the goal is to maximize the total value of items in the knapsack: $\sum_{i\in S}v_{i}$.
Formally, we denote the decision problem of 0/1 Knapsack as
$\textsc{knapsack}(W[1,\ldots,n],V[1,\ldots,n],C,P)$
where $P\in\mathbb{Q}$ is the target value.
knapsack returns true if and only if there exists a subset of items $S\subseteq\{1,2,\ldots,n\}$ such that $\sum_{i\in S}w_{i}\leq C$ and $\sum_{i\in S}v_{i}\geq P$.
It is well known that knapsack is NP-Hard.
B.3 NP-Hardness Proof
Theorem 2.
pem-1d is NP-Hard.
Proof.
[…]
  1. 1.
    Construct $n$ computations and empty vectors $T_{L},T_{H},E_{L},E_{H}$.
  2. 2.
    For $\forall i$, set $t_{i}(L)=w_{i}$ and append to $T_{L}$.
  3. 3.
    For $\forall i$, set $t_{i}(H)=0$ and append to $T_{H}$.
  4. 4.
    For $\forall i$, set $e_{i}(L)=-v_{i}$ and append to $E_{L}$.
  5. 5.
    For $\forall i$, set $e_{i}(H)=0$ and append to $E_{H}$.
  6. 6.
    Set $T^{\prime}$ = $C$ and $EC$ = $-P$.
  7. 7.
    Output $(T=(T_{L},T_{H}),E=(E_{L},E_{H}),T^{\prime},EC)$
Correctness Analysis
[…]
If $(W[1,\ldots,n],V[1,\ldots,n],C,P)\notin\textsc{knapsack}$, there does not exist a subset $S$ such that $\sum_{i\in S}w_{i}\leq C$ and $\sum_{i\in S}v_{i}\geq P$. There are two possibilities: either a subset $S$ that satisfies the weight constraint does not exist at all ($w_{i}>C,\forall i$) or none of the subsets $S$ that satisfy the weight constraint satisfy $\sum_{i\in S}v_{i}\geq P$. For the first possibility, this means all the computations must select the high frequency as the low frequency does not satisfy the time constraint. Then total energy consumption is 0, which is larger than $EC=-P$ since $P\in\mathbb{Q}^{+}$. For the second possibility, for all subsets $S$, $\sum_{i\in S}v_{i}<P$, which means that for all subsets of computations $\sum^{n}_{i=1}(X_{i}e_{i}(L)+(1-X_{i})e_{i}(H))=\sum_{i\in S}-v_{i}>-P=EC$, so none of them satisfy the energy constraint.
Efficiency Analysis
Step 1–5 each takes $O(n)$ time. Step 6 takes $O(1)$ time. Finally, step 7 takes $O(n)$ time.
Therefore, the function $f$ takes $O(n)$ time, which is polynomial time w.r.t the input size.
Appendix C Visualizations for Intrinsic Energy Bloat
Figure 10: Visualization of Perseus’s solution for four stage pipeline workloads on NVIDIA A100 PCIe GPUs. For each workload, the left is running every computation at maximum frequency, and the right is Perseus’s energy schedule that reduces only intrinsic energy bloat without inflating iteration time. Note that these are not real workloads we run in Section 6; real workloads have far more microbatches.
Figure 10 shows the timeline of running one training iteration of BERT 1.3B, T5 3B, Bloom 3B, and Wide-ResNet101 1.5B for maximum frequency and Perseus-optimized energy schedule, respectively. For visualization purposes, we set the number of microbatches to 6. Real evaluation workloads have more microbatches. Energy schedule found by Perseus successfully tunes down frequency for all models without slowing down the iteration time, tightly packing computations over time and reducing intrinsic energy bloat.
Appendix D Full Details of ReduceTime
For the sake of presentation, we made a simplifying statement in the body of the paper that computations only speed up by $\tau$. However, since we aim to speed up all critical paths by precisely $\tau$, speeding up more than one computation from a critical path allows other computations on that critical path to be slowed down. We always take such slowdown opportunity because it will decrease energy consumption.
In the following, we describe the procedure of annotating edges with flow capacities and then solving the problem with maximum flow.
D.1 Generating Capacity DAG
On the computation DAG, we first remove all the computations that are not on any of the critical paths and construct the Critical DAG. We would like to find a set of edges $I^{+}$ to speed up by $\tau$ and $I^{-}$ to slow down by $\tau$ on the Critical DAG, such that the total energy consumption increases minimally. This can be described as the following problem:
$\min_{I^{+},\leavevmode\nobreak\ I^{-}}\quad\sum_{i\in I^{+}}{e_{i}^{+}}-\sum_{i\in I^{-}}{e_{i}^{-}},$ (7)
where $e_{i}(t_{i})$ is the exponential function fit to Pareto-optimal (time, energy) measurments for computation $i$, $e_{i}^{+}=e_{i}(t_{i}-\tau)-e_{i}(t_{i})$ is the extra amount of energy needed to speed up $i$ by $\tau$ and $e_{i}^{-}=e_{i}(t_{i})-e_{i}(t_{i}+\tau)$ is the energy saved from slowing down $i$ by $\tau$.
An important fact that leads us to the solution is that (i) the problem of finding the set of edges to modify such that energy increase is minimized (i.e., solving Equation 7) coincides with (ii) finding the minimum cut of a DAG whose lower and upper bound flow capacities are defined as
$(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$ (8)
This equivalence was discovered by Phillips and Dessouky [49, 24].
Thus, we construct the Capacity DAG from the Critical DAG by annotating its edges with flow capacities given by Equation 8. The capacity of an $S-T$ cut ($S$ is the set of nodes on the source side and $T$ the sink side) on the Capacity DAG with $S\rightarrow T$ edges $I^{+}$ and $T\rightarrow S$ edges $I^{-}$ is identical to the objective in Equation 7. Then, we use the Edmonds-Karp maximum flow algorithm [17] to find the minimum cut of the Capacity DAG. Finally, after the minimum cut has been identified from the Capacity DAG, edges in $I^{+}$ are sped up by $\tau$ and those in $I^{-}$ are slowed down by $\tau$, ultimately reducing the length of every critical path exactly by $\tau$ with the smallest possible energy increase.
D.2 Max Flow Algorithm on the Capacity DAG
A characteristic of our Capacity DAG that precludes the direct application of well-known max flow algorithms is that some edges also have flow lower bounds, asserting that at least a certain amount of flow must pass through in the edge. However, the Max Flow Min Cut theorem by Ford and Fulkerson holds for the case of non-zero flow lower bounds (See Chapter 1, Section 9 of [20]), allowing us to find the minimum cut (which is equivalent to the minimum energy modification set) using any maximum flow algorithm. We adopt an approach that adds dummy source/sink nodes to create a DAG that has zero flow lower bounds, finds the maximum flow on the new DAG, and extracts the flow so that it corresponds to a flow in the original DAG [18]. The algorithm is given in Algorithm 3.
1
2
Input: Directed Acyclic Graph $G=(V,E)$
Source node $s\in V$ and sink node $t\in V$
Lower and upper bounds $l(e)$, $u(e)$ for $\forall e\in E$
Output: A maximum feasible flow of $G$ if it exists
3
4
Construct a new graph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding new source and sink nodes $s^{\prime}$ and $t^{\prime}$, edges from $s^{\prime}$ to each node in $V$, edges from each node in $V$ to $t^{\prime}$, and an edge from $t$ to $s$triangleright Define the capacity $c^{\prime}(e)$ of each edge $e\in E^{\prime}$
5: for $v\in V$ do
6       $c^{\prime}(s^{\prime}v)\leftarrow\sum_{u\in V}l(uv)$ $c^{\prime}(vt^{\prime})\leftarrow\sum_{w\in V}l(vw)$
7 end for
8for $uv\in E$ do
9       $c^{\prime}(uv)\leftarrow u(uv)-l(uv)$
10 end for
$c^{\prime}(ts)\leftarrow\infty$triangleright Find the max flow on G’
$f^{\prime}\leftarrow$ EdmondsKarp ($G^{\prime}$, $c^{\prime}$) $\triangleright$ $G$ has a feasible flow if and only if $G^{\prime}$ has a saturating flow
11 if FlowValue ($f^{\prime}$) $\neq\sum_{v\in V}{c^{\prime}(s^{\prime}v)}$ then
12       return nil
13 end if
$\triangleright$ Convert $f^{\prime}$ to a feasible flow $f$ in $G$
14 for $uv\in E$ do
15       $f(uv)\leftarrow f^{\prime}(uv)+l(uv)$
16 end for
$\triangleright$ Construct residual graph and improve $f$ to max flow
17 for $uv\in E$ do
18       $c(uv)\leftarrow u(uv)-f(uv)$ $c(vu)\leftarrow f(vu)-l(vu)$
19 end for
20return EdmondsKarp ($G$, $c$)
21
Algorithm 3 Maximum Flow with Lower Bounds
Appendix E Proof for Polynomial Runtime
Perseus enumerates the entire “iteration time–energy” Pareto frontier one by one, and the runtime of one iteration is polynomial time with respect to the number of stages $N$ and the number of microbatches $M$. Thus, determining whether the entire algorithm runs in polynomial time reduces to whether the worst case number of iterations is polynomial with respect to $N$ and $M$. While for general DAGs the maximum number of points on the Pareto frontier can be exponential with respect to the size of the DAG [58], here we prove that under mild assumptions for DAGs that represent pipeline schedules, the number of iterations is $O(N+M)$. The assumptions are valid for all pipeline schedules known to the authors, including GPipe [26] and 1F1B [42].
Theorem 3.
For DAGs that represent pipeline schedules, the number of iterations needed is $O(N+M)$.
Proof.
Since we always reduce iteration time by $\tau$, the number of iterations is
$\frac{t_{\max}-t_{\min}}{\tau}$
where $t_{\max}$ and $t_{\min}$ are the maximum and minimum possible iteration time, respectively.
Assume that any pipeline schedule representing one iteration of training has a prologue, a steady state, and an epilogue. The prologue is when the pipeline starts from an empty state and is gradually filled with pipeline computations, while the epilogue is when the pipeline is drained to reach an empty state. It is easy to see that the number of pipeline computations on the critical path of both the prologue and epilogue is $O(N)$, as deeper pipelines (larger $N$) take longer to fill. On the other hand, the steady state of the pipeline is when the pipeline is completely filled, and the number of pipeline computations in any simple path through the steady state of the DAG is $O(M)$. Therefore, the total number of pipeline computations in the critical path of the entire DAG is $O(N+M)$.
$t_{\max}$ and $t_{\min}$ can be constructed by multiplying the number of computations with the average execution time of an computation. Computations are executed with frequencies $f_{\min}$ and $f_{\max}$, respectively, and thus the multipliers of $N$ and $M$ do not cancel out when $t_{\max}-t_{\min}$ is evaluated. Therefore, $t_{\max}-t_{\min}$, and hence $(t_{\max}-t_{\min})/\tau$, is $O(N+M)$. ∎
Appendix F Underestimations of the Emulator
Model Leftmost point Rightmost point
gpt3-xl 20.5% 23.6%
bert-huge-uncased 26.7% 35.3%
t5-3b 19.1% 19.1%
bloom-3b 9.3% 10.7%
wide-resnet101 (width factor 8) 17.5% 19.8%
Table 10: By how much the emulator underestimates the percentage of energy savings for four-stage pipeline workloads on A100 PCIe. Leftmost point represents intrinsic energy bloat, whereas the rightmost point extrinsic.
Our emulator, used to present energy saving results for large scale workloads (GPT-3 175B and Bloom 176B), always underestimates energy savings compared to actual measurements, and argue that the emulated results are empirically a lower bound of actual savings. The characteristics of time and energy changes depends more on the type of the GPU than the parallelization setting as different GPU generations show drastically different sensitivity in terms of time and energy when frequency is varied. Therefore, we compare results from the A100 PCIe GPU.
Table 10 shows the percentage of energy savings underestimation for two points on the “iteration time–energy” frontier:
  • Leftmost point: The point with the shortest iteration time, corresponding to how much the emulator underestimates the amount of possible intrinsic energy bloat reduction.
  • Rightmost point: The point with the longest iteration time, corresponding to how much the emulator underestimates the amount of possible energy savings brought about by slowing down the entire pipeline in the case of a straggler.
We believe that underestimation is happening due to our simplifying assumption that $P_{\mathrm{blocking}}$, the power consumption of the GPU while blocking on communication, is constant regardless of the GPU’s frequency. In reality, $P_{\mathrm{blocking}}$ would be an increasing function of frequency. Therefore, when the emulator assumes a constant $P_{\mathrm{blocking}}$ when calculating the energy consumption of a pipeline that is overall running on a reduced frequency, it will be overestimating its total energy consumption, leading to underestimations.
Appendix G Large-Scale Emulation
Section 6.3 used A100 SXM GPUs, which we momentarily had access to, in order to profile model layers and run large scale emulation. This is because we believe that SXM models are more representative of real large scale GPU clusters that train large models. Nonetheless, in this section, we provide the same set of results from A100 PCIe GPUs, as well, along with the same A40 GPUs results. All other experiment methodology and parameters are identical.
Results Summary.
Figure 11: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1024 GPUs.
Figure 11 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1024 GPUs as a representative.
Intrinsic Bloat Reduction.
Model GPU Type
Energy Savings (%)
given # Microbatches
12 24 48 96
GPT-3 175B A100 9.02 9.77 10.17 10.39
A40 11.81 10.22 9.34 8.88
Bloom 176B A100 6.39 4.99 4.21 3.80
A40 6.97 4.49 3.12 2.41
Table 11: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
Table 11 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches. GPT-3 175B on A100 is unique and has the opposite trend compared to others: intrinsic energy bloat reduction increases with the number of microbatches. The opposite trends stem from the amount of stage imbalance on each GPU, and the amount of slowdown the minimum-energy frequency incurs. What is unique about GPT-3 175B on A100 is that the amount of imbalance is large whereas the minimum-energy frequency incurs little slowdown for forward and backward computations. Therefore, only for GPT-3 175B on A100, microbatches in the middle of the pipeline are able to fully slow down to their potential. Thus, when such microbatches are added to the pipeline, the percentage of intrinsic energy bloat reduction increases.
Extrinsic Energy Bloat.
Figure 12: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines. When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$. More blocking time after $T^{*}$ decreases overall energy savings. Please note the different Y-axes.
Figure 12 presents extrinsic energy bloat results in the strong scaling configuration, where a straggler is slower than others in varying degrees of slowdown from 1.05 to 1.50.
Appendix H Iteration Time–Energy Frontiers
(a) BERT
(b) T5
(c) Bloom
(d) Wide-ResNet
Figure 13: Eight stage pipeline parallelism on A40.
Figure 13 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
T5 shows an interesting frontier due to the hardware topology of our A40 machine setup: Each node has four GPUs and NVLink connects GPUs 0 and 1, and 2 and 3; GPUs 1 and 2 must communicate through the NUMA interconnect; Finally, different nodes are connected with Infiniband only adjacent to GPUs 0 and 1 (data to and from GPUs 2 and 3 must also go through the NUMA interconnect). The implication of this heterogeneous GPU interconnect is that if more than one P2P communications that need to go through the NUMA interconnect happen at the same time, contention happens and both of data transfers slow down significantly. However, Perseus’s plan reduces this contention, overall decreasing iteration time noticeably. Yet, contention and noisy communication latencies still exist, leading to a noisy frontier.
For many ZeusPerStage lines, energy fluctuates significantly when iteration time increases due to ZeusPerStage being unaware of critical paths. Balancing the forward computation time between stages could even let the modified stages take over the critical path. As a result, the iteration time increases, which increases energy bloat, and more energy is spent on blocking on communication (§4.1). When the decreased energy on computation fails to cover the increased energy on P2P communication, total energy increases.
(a) BERT
(b) T5
(c) Bloom
(d) Wide-ResNet
Figure 14: 4 stage pipeline parallelism on A100 PCIe.
Figure 14 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs. Wide-ResNet has a noisy frontier because the variability in microbatch loading time introduces noise in the end-to-end iteration time when computations are tightly packed by Perseus. This was not pronounced in A40 GPUs because compared to A100 PCIe, computation is slower, but data loading time is similar. Thus, the noise in data loading time becomes more noticeable in A100 PCIe.
; bottom:-0.4ex;">Exml[LOGO] body>

### v2

@@
-Perseus: Removing Energy Bloat from Large Model Training
-Abstract
-Training large AI models on numerous GPUs consumes a massive amount of energy.
-We observe that not all energy consumed during training directly contributes to end-to-end training throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
-In this work, we identify two independent sources of energy bloat in large model training, intrinsic and extrinsic, and propose Perseus, a unified optimization framework that mitigates both.
-Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient iterative graph cut-based algorithm and schedules energy consumption of its forward and backward computations across time to remove intrinsic and extrinsic energy bloat.
-Evaluation on large models like GPT-3 and Bloom shows that Perseus reduces energy consumption of large model training by up to 30%, enabling savings otherwise unobtainable before.
-1 Introduction
-As deep neural networks (DNNs) continue to grow in model and dataset size [29, 25], the energy consumption of large model training is increasing as well.
-For instance, training GPT-3 [11] reportedly consumed 1,287 megawatt-hour (MWh) in 2021 [48].
-The recent surge in the development of generative AI (GenAI) and large language models (LLMs) shows no signs of deceleration in AI energy demands.
-Optimizing the energy consumption of large model training is, therefore, crucial to minimizing the operational costs associated with AI, with additional benefits.
-For example, low-carbon electricity, which is preferred especially in the context of sustainability commitments [1, 4, 6, 7], is still very much a scarce and contended resource [50].
-Reducing energy consumption allows more workloads to fit within the limited supply of low-carbon electricity [10].
-Furthermore, reducing energy consumption without slowdown reduces power draw, either allowing more aggressive datacenter power oversubscription or conversely reducing power throttling for existing workloads [35, 54, 46].
-Despite many recent works on increasing the throughput of large model training [26, 41, 19, 43, 73, 34, 27], reining in the growing energy consumption remains an open challenge [55, 48].
-In this paper, we aim to identify and remove energy bloat – the portion of energy consumption that can be removed without throughput loss – in large model training.
-We identify two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
+Reducing Energy Bloat in Large Model Training
+Abstract.
+Training large AI models on numerous GPUs consumes a massive amount of energy, making power delivery one of the largest limiting factors in building and operating datacenters for AI workloads.
+However, we observe that not all energy consumed during training directly contributes to end-to-end throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
+In this work, we identify two independent sources of energy bloat in large model training and propose Perseus, a training system that mitigates both.
+To do this, Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient graph cut-based algorithm and schedules the energy consumption of computations across time to remove both types of energy bloat.
+Evaluation on large models including GPT-3 and Bloom shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification, enabling energy reduction – and therefore cost savings – otherwise unattainable before.
+1. Introduction
+As deep neural networks (DNNs) continue to grow in model and dataset size (scalinglaws-arxiv20, ; chinchilla-neurips22, ), the energy consumption of large model training is increasing as well.
+For instance, training GPT-3 (gpt3, ) reportedly consumed 1.3 GWh (patterson2021carbon, ).
+Then, this was recently dwarfed by a 200B model trained by Amazon consuming about 11.9 GWh (hamilton2024constraint, ), which is enough to power more than 1,000 average US households for a year (us-household, ).
+Such energy-intensive large model training workloads not only inflate datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ).
+Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ; oobleck-sosp23, ), energy optimization still remains an open challenge (greenai-cacm20, ; patterson2021carbon, ).
+While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute.
+As such, a recent line of work demonstrated that software plays a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; enemy-hpca23, ; envpipe-atc23, ).
+In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing the average power draw.
+To that end, we identify energy bloat – the portion of energy consumption that can be removed without slowdown – in software systems for large model training.
+We find two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
 Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2).
-Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning [41, 19, 23, 73], but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation.
-When all stages do not have the same computation intensity, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput.
+Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning (pipedream-sosp19, ; dapple-ppopp21, ; autopipe-icml21, ; alpa-osdi22, ), but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation.
+When all stages do not have the same computation time, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput.
 Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
 Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3).
-Root causes behind such slowdowns are varied, including power/thermal throttling [75, 47, 35, 54, 46], I/O bottlenecks in the storage/network [39, 71], hardware/software failures [61, 27], etc., and the likelihood of their presence increases with the scale and duration of training [22, 28, 64].
+Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ), hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ), etc., and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ).
 All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput.
-Extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
-In this work, we propose Perseus that relies on a unified optimization framework to remove both intrinsic and extrinsic energy bloat during large model training (§3).
-At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, which enables it to minimize intrinsic energy bloat under normal operating conditions and to mitigate extrinsic energy bloat arising from straggler pipelines.
+Thus, extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
+In this work, we propose Perseus, which formulates a unified optimization framework to remove both intrinsic and extrinsic energy bloat from large model training (§3).
+At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers.
 Existing works fall short on both fronts.
-EnvPipe [13] is limited only to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction.
-Zeus [68], in contrast, ignores intrinsic bloat altogether as it only considers single-GPU training, which also leads it to identify suboptimal “iteration time–energy” frontiers at best.
-We prove that optimally solving our optimization problem is in fact NP-hard and propose an efficient algorithm that provides an exact solution for a relaxed problem (§4).
-At its core, our solution represents an iteration of large model training as a directed acyclic graph (DAG).
+EnvPipe (envpipe-atc23, ) is limited to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction.
+Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its “iteration time–energy” frontier suboptimal for large models.
+We show that characterizing the “iteration time–energy” Pareto frontier is NP-hard not only to solve but also to approximate within any constant factor.
+Given this impasse, we propose an efficient algorithm that optimally solves a relaxed problem instead (§4).
+At its core, Perseus represents one training iteration as a directed acyclic graph (DAG).
 Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies.
-Importantly, each node is annotated with its planned execution time and energy consumption, which we call the energy schedule.
-Perseus’s optimization algorithm starts from the energy schedule that consumes the least possible amount of energy, and iteratively speeds up the execution time of the schedule while incurring the least possible energy increase, until no more speed up is possible.
-This generates all energy schedules on the “iteration time–energy” Pareto frontier.
+Perseus’s optimization algorithm then efficiently generates all energy schedules, defined as the planned time and energy consumption of each DAG node (computation), that are on the “iteration time–energy” Pareto frontier.
+In doing so, instead of finding each point on the frontier individually, Perseus employs a graph cut-based algorithm that iteratively crawls up the frontier from the bottom, generating the next Pareto-optimal energy schedule by carefully tweaking its neighbor energy schedule on the frontier.
 Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
-While solving the optimization problem depends on accurate fine-grained computation energy profiling, modern GPUs provide energy consumption information only in coarse-grained time intervals.
-Perseus introduces an online profiling technique that can accurately profile the energy consumption of fine-grained computations with low overhead (§5).
-Evaluation on a variety of large models including GPT-3 [11], BERT [15], T5 [51], Bloom [66] and a scaled-up version of Wide-ResNet [69] shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown (§6).
+Perseus is split into a server and a client library (§5).
+The client library can be integrated into any large model training framework as well as any accelerator that has an energy measurement and execution speed control interface, giving the server an abstract view of training computations and their time and energy consumption.
+The server produces Pareto-optimal energy schedules, which is realized by the client.
+Evaluation on large models including GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and scaled-up Wide-ResNet (wideresnet-bmvc16, ) shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing both energy consumption and average power draw (§6).
 Overall, we make the following contributions in this paper:
   • We identify intrinsic and extrinsic energy bloat in large model training, fundamentally caused by computation time imbalance at different levels.
-  • We propose Perseus, an open-source unified optimization framework that aims to remove both types of energy bloat using an efficient graph cut-based algorithm.
-  • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat.
-2 Motivation
-First, we provide necessary background regarding large model training (§2.1). Then, we introduce intrinsic (§2.2) and extrinsic (§2.3) energy bloat present in large model training.
-2.1 Large Model Training
-The rapid increase in DNN model size in recent years made data, tensor, and pipeline parallelism [26, 41, 43, 19, 73, 34] essential ingredients for large model training.
-Especially, pipeline parallelism partitions a large model into multiple stages and assigns each stage to different GPUs.
-The training batch is split into multiple microbatches, and their forward and backward computations are pipelined through those stages.
-Then, in order to scale out to more GPUs, such pipelines are replicated multiple times to perform data parallel training, where each pipeline computes a subset of the training batch and synchronizes gradients with one other at the end of each iteration.
-Due to the synchronization barrier in the end, the next iteration can only begin when all pipelines have finished.
-2.2 Intrinsic Energy Bloat
-Figure 1: One training iteration of GPT-3 1.3B with 4 pipeline stages (partitioned with minimum imbalance ratio) and 6 microbatches on NVIDIA A100 GPUs drawn to scale.
+  • We propose Perseus, a software-only energy optimization system that removes energy bloat through a unified optimization framework and a graph cut-based algorithm.
+  • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat, bringing hardware-agnostic energy savings.
+2. Motivation
+First, we provide necessary background regarding large model training (§2.1).
+Then, we introduce intrinsic (§2.2) and extrinsic (§2.3) energy bloat present in large model training, and discuss opportunities for energy reduction (§2.4).
+2.1. Large Model Training
+Large model training today is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ; llama3-blog, ; deepspeed-github, ; megatronlm-github, ; gpt-neox-github, ).
+Especially, pipeline parallelism partitions a large model and its training batch into multiple stages and microbatches respectively, and pipelines forward and backward computations through stages.
+Then, such pipelines are replicated multiple times to perform data parallel training.
+Since all pipelines must synchronize gradients at the end of each iteration, computations for the next iteration can only begin when all pipelines have finished and exchanged gradients.
+2.2. Intrinsic Energy Bloat
+Figure 1.
+One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs drawn to scale.
 For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2.
 The critical path of computation is traced with a blue line.
-Colors are average power consumption.
-The blue background indicates that GPUs consume power when blocking on communication.
-Model Size Imbalance Ratio
+Colors show power consumption.
+Especially, the blue background indicates that GPUs consume power when blocking on communication.
+Other models are visualized in Appendix A.
+Model # Parameters Imbalance Ratio
 4 stages 8 stages
-GPT-3 [11] 3B 1.13 1.25
+GPT-3 (gpt3, ) 3B 1.13 1.25
 7B 1.11 1.23
@@
 175B 1.02 1.03
-Bloom [66] 3B 1.13 1.25
+Bloom (bloom-arxiv22, ) 3B 1.13 1.25
 7B 1.13 1.25
 176B 1.05 1.10
-BERT [15] 0.1B 1.33 2.00
+BERT (bert, ) 0.1B 1.33 2.00
 0.3B 1.17 1.33
-T5 [51] 0.2B 1.19 1.50
+T5 (t5-jmlr20, ) 0.2B 1.19 1.50
 0.7B 1.05 1.11
 2.9B 1.06 1.16
-Wide-ResNet50 [69] 0.8B 1.23 1.46
-Wide-ResNet101 [69] 1.5B 1.09 1.25
-Table 1: Imbalance ratio between the longest and the shortest stages for various models on NVIDIA A100 GPUs. 1.00 would mean perfect balance.
+Wide-ResNet50 (wideresnet-bmvc16, ) 0.8B 1.23 1.46
+Wide-ResNet101 (wideresnet-bmvc16, ) 1.5B 1.09 1.25
+Table 1. Latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
 We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a).
-In addition to the familiar bubbles found in any pipeline schedule [41], we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage.
+In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage.
 Such gaps exist because the computation time of each pipeline stage is not perfectly balanced.
-Partitioning stages in a balanced manner is an important problem in distributed execution planning [26, 41, 19, 73], but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
-To understand the amount of possible pipeline stage imbalance, we exhaustively enumerated all possible pipeline partitions [^1] For Transformer-based models, we partition at the granularity of Transformer layers.
+Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
+To understand the amount of possible pipeline stage imbalance, we exhaustively searched for the pipeline partition with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest.
+[^1] For Transformer-based models, we partition at the granularity of Transformer layers.
 For Wide-ResNet, we partition at the granularity of bottleneck layers, which are three convolution layers wrapped with a skip connection.
-to find the one with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest.
 Table 1 lists the minimum imbalance ratio for various models, which shows that perfect balance is difficult to achieve.
-Appendix A explains the partitioning details of each model and where imbalance comes from.
-Given some amount of stage imbalance, as we have seen in Figure 0(a), not all forward and backward computations are on the critical path of computation.
+See Appendix B for partitioning details and sources of imbalance.
+Given stage imbalance, as we have seen in Figure 1(a), not all forward and backward computations are on the critical path of computation.
 This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy.
-We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation so that they do not affect the critical path (Figure 0(b)).
-Although seemingly simple, we prove that this problem is NP-hard by reduction from the 0/1 Knapsack problem (Appendix B).
-Visualizations for other models are in Appendix C.
-2.3 Extrinsic Energy Bloat
-During large model training, numerous replicas of the same pipeline run in a data parallel fashion in order to train on more data with more GPUs, and every pipeline must synchronize gradients at the end of each iteration before moving on to the next.
-When one pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes.
-Figure 1(a) illustrates this situation.
-Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time are wasteful.
-We call this extrinsic energy bloat, because unlike intrinsic energy bloat, the cause of energy bloat is extrinsic to the pipeline.
-Perseus finds the energy-optimal iteration time for the non-straggler pipeline, again by precisely slowing down computations in the pipeline (Figure 1(b)).
-However, unlike other problem settings that do not consider energy consumption, we show in Section 3.1 that fully utilizing all the slack time created by the straggler may not be energy-optimal!
-Perseus can automatically make such decisions and implements the energy-optimal slowdown.
-Figure 2: Toy example illustrating one straggler pipeline running together with a non-straggler pipeline using data parallelism.
-The non-straggler wastes energy by running as fast as possible.
-Perseus can precisely slow down non-straggler pipelines to reduce such extrinsic energy bloat.
-Stragglers arise from numerous sources, with many introducing sizable energy bloat according to recent reports.
-Thermal or power throttling in a datacenter may lower the GPU’s frequency to prevent hardware or the power supply unit from being damaged, resulting in 10–50% slowdown [75, 47, 35, 54, 46].
-I/O bottlenecks in the storage or network, especially during data loading for image, video, or recommendation models, can be longer than GPU computation by up to 4×4$\times$4 × [39, 71], thereby acting like a persistent (virtual) straggler pipeline.
-Recent failure-resilient large model training frameworks [61, 27] deploy heterogeneous pipelines with different number of stages or amount of compute after failures, introducing non-uniform iteration times across pipelines until recovery.
-As the scale of large model training jobs increase, the probability of the job encountering one or more causes of stragglers increases [22, 28, 64].
-Stragglers due to events like thermal/power throttling, consistent I/O bottlenecks, and fault-tolerant planning can persist across many iterations [46, 39, 27].
-A training system can identify such stragglers and calculate the extent of slowdown by comparing iteration times across data-parallel pipelines.
-In this work, we assume that such knowledge is available and focus on how to plan time and energy consumption across time and allow quick adaptation given such information.
-3 Perseus Overview
+We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation without lengthening the critical path (Figure 1(b)).
+Although seemingly simple, this problem is not only NP-hard to solve, but also NP-hard to even approximate to any constant factor (svensson2012hardness, ).
+2.3. Extrinsic Energy Bloat
+Numerous replicas of the same pipeline run in a data parallel fashion in large model training.
+Because every pipeline must synchronize gradients at the end, if a pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)).
+Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time is wasteful.
+We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the pipeline.
+To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time and precisely slow down computations in the non-straggler keeping that iteration time (Figure 2(b)).
+Figure 2.
+Two data parallel pipelines run, and the first one became a straggler (e.g., due to the second GPU temporarily overheating).
+The non-straggler pipeline wastes energy by running as fast as possible.
+Such extrinsic energy bloat can be removed by precisely slowing down non-straggler pipelines.
+Stragglers arise from numerous sources.
+Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ), acting like a persistent straggler pipeline.
+Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ) deploy heterogeneous pipelines, introducing non-uniform iteration times.
+With increasing job and infrastructure scale, the probability of encountering stragglers increases (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ; megascale-nsdi24, ; tpuv4-nsdi24, ).
+Stragglers are often times known to and anticipated by the training infrastructure, because they were created by the infrastructure itself (e.g., throttling, non-compute bottlenecks, fault-tolerant planning).
+Therefore, in this work, we focus on how to plan time and energy across time and allow quick adaptation assuming that such information is available.
+2.4. Potential Benefits of Reducing Energy Bloat
+To gauge potential energy savings, we measure the energy savings of slowing down every computation in the pipeline to their minimum-energy frequencies.
+This will slow down the pipeline’s iteration time, but can act as an upper bound for energy savings.
+For our experimental workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively.
+We show in Section 6.2 that Perseus can realize most of the potential savings with negligible slowdown, even through our problem is NP-hard.
+Figure 3.
+Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be.
+$T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier.
+The black dot is when all computations run at the maximum speed, which wastes energy.
+The green dot is the energy-optimal iteration time of the non-straggler pipeline.
+Solid orange dots make up the “iteration time–energy” Pareto frontier, and the orange dotted line shows that energy increases beyond $T^{*}$.
+3. Perseus Overview
 Perseus is an energy optimization system for large model training.
 In this section, we present our unified optimization framework that aims to remove both types of energy bloat (§3.1) and walk through the workflow of Perseus (§3.2).
-3.1 Optimization Approach
-Figure 3: Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be.
-The first case is when there are no stragglers.
-$T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier.
-The black dot is the default mode of execution where all computations run at the maximum speed.
-The green dot is the energy-optimal iteration time of the non-straggler pipeline.
+3.1. Unified Optimization Framework
 Intuitively, slowing down computations selectively in a training pipeline without affecting its critical path will keep the same iteration time while reducing its energy consumption (§2.2).
@@
 We formalize these two intuitions into a unified optimization framework and derive a universal prescription for a non-straggler pipeline’s energy-optimal iteration time.
-By controlling the execution speed of individual computations in the pipeline, our goal is to minimize the energy consumption of a pipeline, and we can safely slow down a pipeline’s iteration time until the straggler’s iteration time $T^{\prime}$:
-$\min_{F}$ $\quad\textrm{Energy}(F)$ (1)
+Our goal is to minimize a pipeline’s energy consumption by controlling the execution speed of each computation in the pipeline.
+In doing so, we can slow down a pipeline’s iteration time at most until the straggler’s iteration time $T^{\prime}$:
+(1) $\min_{F}$ $\quad\textrm{Energy}(F)$
 s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
-where $F$ is the set of GPU frequencies to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively.
+where $F$ is the set of GPU frequencies [^2] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms (AMD GPUs via AMDSMI (amdsmi, )), which is much shorter than typical large model computation latencies.
+Locking the GPU’s frequency provides deterministic computation latency (nvml, ; clockwork-osdi20, ), making it suitable for tightly planning and packing execution over time.
+to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively.
 Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane.
-However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier, a monotonically decreasing tradeoff curve.
+However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier.
+Now, let us assume we have a fully characterized “iteration time–energy” Pareto frontier with $T_{\min}$ and $T^{*}$ bookending the frontier (described in Section 4).
+$T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption.
 Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t.
-the “iteration time–energy” Pareto frontier.
-$T_{min}$ is the shortest iteration time on the frontier, which coincides with the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption.
-Together, $T_{\min}$ and $T^{*}$ bookend the Pareto frontier.
-  1. 1.
-    When there are no stragglers (Figure 2(a)), we simply operate on the point on the Pareto frontier with the same iteration time, which reduces intrinsic energy bloat.
-  2. 2.
-    When a moderately slow straggler is detected (Figure 2(b)), we additionally reduce extrinsic energy bloat by slowing down non-straggler pipelines until $T^{\prime}$, using up all the slack time created by the straggler.
-  3. 3.
-    Finally, when the straggler’s iteration time goes beyond the minimum-energy point $T^{*}$ on the frontier (Figure 2(c)), we stop at $T^{*}$ instead of $T^{\prime}$.
-This is because slowing down beyond $T^{*}$ will instead increase energy.
+the Pareto frontier:
+  1. (1)
+    When there are no stragglers (Figure 3(a)), we simply operate on the point on the Pareto frontier with iteration time $T_{\min}$, which reduces intrinsic energy bloat.
+  2. (2)
+    When a moderately slow straggler is detected (Figure 3(b)), we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
+  3. (3)
+    Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier (Figure 3(c)).
+In this case, we only slow down non-stragglers until $T^{*}$, because going past $T^{*}$ will instead increase energy.
 The three cases can be merged into one universal prescription for the pipeline’s energy-optimal iteration time:
-$T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$ (2)
-Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier, we know the solution of Equation 1 for any feasible value of $T^{\prime}$.
-Then our system will be able to react quickly to emerging stragglers by simply looking up the set of frequencies $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$.
-We present our algorithm to efficiently characterize the entire “iteration time–energy” Pareto frontier in Section 4.
-3.2 Perseus Architecture
+(2) $T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$
+Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier by finding all Pareto-optimal $F$s, Perseus will be able to react instantly to emerging stragglers by looking up $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$.
+Finally, we note that unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal.
+Being too fast or too slow can both waste energy.
+3.2. Perseus Architecture
 Energy Schedule.
-Perseus represents each iteration of the training pipeline as a static directed acyclic graph (DAG), where nodes are forward and backward computations and edges are dependencies in between computations.
+Perseus represents each iteration of the training pipeline as a static directed acyclic graph (DAG), where nodes are forward and backward computations in each stage and edges are dependencies between computations.
 Each node on the computation DAG is annotated with its planned time and energy consumption, which we call the energy schedule.
@@
 Figure 4 shows the high-level architecture of Perseus.
-Perseus is split into a framework-independent server and a framework-integrated client.
-The server runs our time–energy optimization algorithm, produces all energy schedules on the “iteration time–energy” Pareto frontier, and caches them in a table for fast lookup.
-The client profiles computations of the pipelines online during training, and realizes the energy schedule by changing the GPU’s frequency during runtime.
-Figure 4: Perseus architecture and workflow.
+Perseus is split into a framework- and hardware-agnostic server and a framework-integrated and device-specific client.
+The server pre-characterizes the “iteration time–energy” Pareto frontier (§4) and caches them for fast lookup.
+The client profiles pipeline computations online during training and realizes the energy schedule by changing the GPU’s frequency during runtime (§5).
+Figure 4. Perseus architecture and workflow.
 Training Lifecycle Using Perseus.
-In Perseus, a training job is specified by its computation DAG.
-When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy consumption of each forward and backward computation in each stage while training is running, and sends the profile to the server.
-As training proceeds, 2 the server asynchronously characterizes the “iteration time–energy” Pareto frontier of the training job (§4), after which 3 the Pareto-optimal energy schedule corresponding to the shortest possible iteration time $T_{\min}$ is deployed to the client.
-Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework’s pipeline execution engine (§5).
-Pareto-optimal energy schedules are saved in a lookup table format index by $T^{\prime}$.
-During training, 4 the hardware infrastructure (e.g., data center rack power manager) notifies the Perseus server of a straggler pipeline and how much slowdown it will experience.
-The Perseus server can 5 quickly react to this by looking up the Pareto-optimal energy schedule corresponding to that straggler iteration time, and deploying it to the clients.
-4 Algorithm Design
-In this section, we describe our algorithm to obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail.
-We start by formulating the problem (§4.1).
+In Perseus, a training job is primarily specified by its computation DAG for one training iteration.
+When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy of each forward and backward computation on each frequency while training is running.
+Upon receiving profile results, 2 the server characterizes the “iteration time–energy” Pareto frontier (§4), and Pareto-optimal energy schedules are saved in a lookup table indexed by $T^{\prime}$.
+Then, 3 the energy schedule corresponding to the shortest iteration time is deployed to the client.
+Energy schedules are realized by the client’s Asynchronous Frequency Controller, integrated into the training framework (§5).
+During training, 4 the training infrastructure (e.g., datacenter rack power/temperature manager) notifies the Perseus server of a straggler and its degree of slowdown.
+The server then 5 quickly reacts to this by looking up the Pareto-optimal energy schedule corresponding to the anticipated straggler iteration time, and deploys it to the client.
+4. Characterizing the Pareto Frontier
+In this section, we describe our algorithm to efficiently obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail.
+We first formulate the problem, show that it is NP-hard, and provide a relaxation (§4.1).
 Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3).
-Finally, we extend our algorithm to support tensor parallelism, single-choice operations, and diverse pipeline schedules (§4.4).
-4.1 Problem Formulation
+Finally, we extend our algorithm to support 3D/hybrid parallelism, constant-time operations, and diverse pipeline schedules (§4.4).
+4.1. Problem Formulation
 Expression for Energy Consumption.
-The energy consumption of a pipeline is not only from computation; it is the sum of three parts: (1) Computation; (2) Blocking on communication between computations; and (3) Blocking on communication until the straggler pipeline finishes.
-Formally,
-$\sum_{i}{e_{i}(f_{i})}+P_{\textrm{blocking}}(N\cdot T-\sum_{i}{t_{i}(f_{i})})+P_{\textrm{blocking}}\cdot N\cdot(T^{\prime}-T)$ (3)
+The energy consumption of a pipeline is not only from computation; it is the sum of three parts: (1) Computation; (2) Blocking on communication between computations; and (3) Blocking on communication until the straggler pipeline finishes:
+(3) $\sum_{i}{e_{i}(f_{i})}+P_{\textrm{blocking}}(N\cdot T-\sum_{i}{t_{i}(f_{i})})+P_{\textrm{blocking}}\cdot N\cdot(T^{\prime}-T)$
 […]
 where $P_{\textrm{blocking}}$ is the power consumption of the GPU when it is blocking on communication, $N$ is the number of pipeline stages, and $t_{i}(f_{i})$ and $e_{i}(f_{i})$ are the time and energy consumption of computation $i$ with frequency $f_{i}$, respectively.
-Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline depends on the straggler’s iteration time $T^{\prime}$ in 2.222This makes sense because if the straggler slows down more and more, the non-straggler will consume more energy waiting for the straggler to finish.
-Specifically, the Pareto frontier for any $T^{\prime}$ is the Pareto frontier of $\textrm{Time}(F)$ vs.
-1 shifted upwards by 2.
-Here, 1 does not depend on $T^{\prime}$.
+As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the “iteration time–energy” Pareto frontier ($\textrm{Time}(F)$ vs.
+1+ 2) at $T_{\textrm{opt}}$ and find $F_{\textrm{opt}}$ where the two lines intersect.
+Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline only depends on the straggler’s iteration time $T^{\prime}$ in 2, which is only a constant upwards shift of the frontier.
 Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs.
-1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$ as shifting the frontier upwards does not affect iteration time nor the frequency assignments that lead to each iteration time.
+1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$.
 Thus, we define
-$\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$ (4)
+(4) $\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$
 and characterize the Pareto frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
 Finding the Pareto Frontier.
-Finding one point on the “iteration time–energy” Pareto frontier with iteration time $T^{\prime}$ is equivalent to solving the following optimization problem:
-$\min_{F}$ $\quad\textrm{Energy}(F)$ (5)
-s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
+Finding one point on the Pareto frontier with iteration time $T$ is equivalent to solving the following optimization problem:
+(5) $\min_{F}$ $\quad\textrm{Energy}(F)$
+s.t. $\quad\textrm{Time}(F)\leq T$
 We call this problem Pipeline Energy Minimization (PEM).
-Theorem 1.
+Theorem 4.1.
 Pipeline Energy Minimization is NP-hard.
 Proof.
-Reduction from 0/1 Knapsack. Details in Appendix B. ∎
-The complete Pareto frontier can be obtained by solving PEM for all $T^{\prime}\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
-One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is in turn because the possible choices of GPU frequencies are discrete.
-However, if choices were continuous, the problem is exactly and efficiently solvable [57].
+Reduction from Knapsack. Details in Appendix C. ∎
+The complete Pareto frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
+One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is because the possible choices of GPU frequencies are discrete.
+However, if choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ).
 This is akin to integer linear programs becoming tractable when relaxed to linear programs.
-We fit the exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements to make the choices continuous.
-We choose the exponential function due to its inherent flexibility and natural fit to data.
-We show in Section 6 that this relaxation produces high-quality approximate solutions that lead to actual savings.
-After relaxation, the problem’s optimization variables are the time and energy consumption planned for each computation in the pipeline, or the energy schedule.
-4.2 Algorithm Overview
-Iteratively Discovering the Pareto Frontier.
-Figure 5: Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
-While the relaxed Pipeline Energy Minimization problem is no longer NP-Hard, solving the problem for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient.
-Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate the next energy schedule adjacent to it on the Pareto frontier?
-Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, and every energy schedule we encounter will be Pareto-optimal.
+The transform from the original problem to the relaxed version is done by fitting a continuous exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements for each forward and backward computation.
+We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realizes most of the opportunity for savings.
+Solving the relaxed problem returns the time and energy consumption planned for each computation in the pipeline, or the energy schedule.
+Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$ (Section 4.3).
+4.2. Iteratively Discovering the Pareto Frontier
+Figure 5. Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
+Although our relaxed problem is no longer NP-hard, solving it for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient.
+Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate its neighbor energy schedule on the Pareto frontier?
+Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, discovering every Pareto-optimal energy schedule.
 […]
-We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) is incorrect.
+We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) would be incorrect.
 That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
-TL;DR.
 1
 2
+3
 Input: DAG $\mathcal{G}$ of computations $i\in\mathcal{G}$
@@
 Output: Set of Pareto-optimal energy schedules $\mathcal{P}$
-3
 4
+5
+6
 $\triangleright$ Begin with the minimum energy schedule
-5 $p\leftarrow$ Minimum energy for all computations $\mathcal{P}\leftarrow\{p\}$ while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
+7 $p\leftarrow$ Minimum energy for all computations
+8 $\mathcal{P}\leftarrow\{p\}$
+9
+10while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
        $\triangleright$ Reduce time by $\tau$ with minimal energy increase (§4.3)
-6       $p\leftarrow$ ReduceTime($\mathcal{G}$, $p$, $\tau$) $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
+11       $p\leftarrow$ GetNextPareto($\mathcal{G}$, $p$, $\tau$)
+12       $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
+13
 14 end while
+15
 $\triangleright$ Every schedule in $\mathcal{P}$ is Pareto-optimal
 16 return $\mathcal{P}$
-9
+17
+18
 Algorithm 1 Iteratively Discovering the Pareto Frontier.
@@
 This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process is Pareto-optimal.
-4.3 Reducing Time with Minimum Energy Increase
-Figure 6: A simplified example of how to reduce iteration time by unit time $\tau$.
+4.3. Finding the Neighbor Pareto-Optimal Schedule
+Figure 6.
+A simplified example of how to reduce iteration time by unit time $\tau$.
 Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2).
@@
 Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
-In this section, we describe our core subroutine ReduceTime (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
+In this section, we describe our core subroutine GetNextPareto (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
 Node- and Edge-Centric Computation DAGs.
@@
 As a setup for subsequent steps, we convert this into an edge-centric computation DAG where computations are edges and dependencies are nodes (i.e., all incoming edges must complete before any outgoing edge can begin).
-This conversion can be done by splitting each node on the original computation DAG into two nodes and connecting the two with an edge annotated with the computation on the original node.
+This conversion can be done by splitting each node into two and connecting the two with an edge annotated with the computation on the original node.
 Removing Non-Critical Computations.
-Our goal is to reduce the execution time of the computation DAG by $\tau$, which is equivalent to reducing the length of all critical paths by $\tau$.
-Since computations that are not on any critical path (i.e., non-critical computations) do not affect the length of the critical path, we simply remove them from the computation DAG.
+Our goal is to reduce the execution time of the computation DAG by $\tau$, which is equivalent to reducing the length of all critical paths by $\tau$ [^5] Let’s say there are two critical paths that run in parallel.
+They must be of equal length to both be critical paths.
+Here, if only one were shortened, the other will remain the sole critical path and the DAG will not execute faster.
+Since computations that are not on any critical path (i.e., non-critical computations) do not affect the length of the critical path, we remove them from the computation DAG.
 Finding Computations to Speed Up.
@@
 Figure 6 shows two examples of this.
-We have already 1 transformed an energy schedule into 2 a computation DAG where computations are represented by edges, and 3 removed all non-critical computations.
-Here, 4 shows two valid s-t cuts: Cut A and Cut B.
-5A speeds up the Backward computation cut by Cut A from $3\tau$ to $2\tau$, and the iteration time of the energy schedule was reduced by $\tau$.
-Similarly, 5B speeds up the Forward computation and Backward computation defined by Cut B from $5\tau$ to $4\tau$ and from $7\tau$ to $6\tau$ respectively, and the iteration time of the energy schedule was also reduced by $\tau$.
+4 shows two valid s-t cuts: Cut A and Cut B.
+5A speeds up the computation edge cut by Cut A from $3\tau$ to $2\tau$, and the iteration time of the energy schedule was reduced by $\tau$.
+Similarly, 5B speeds up the computation edges cut by Cut B from $5\tau$ to $4\tau$ and from $7\tau$ to $6\tau$, and the iteration time of the energy schedule was also reduced by $\tau$.
 Especially, in the second case, iteration time was only reduced because computations on two parallel critical paths were sped up together.
 Solving with Minimum Cut.
-We have shown with examples that any s-t cut represents a way to reduce the duration of the computation DAG by $\tau$.
-Then, a natural question is, which cut brings the smallest possible energy increase?
-We can precisely map the capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of the exponential function fit for that computation and defining it to be the edge’s flow capacity.
+We have shown that an s-t cut represents a way to speed up the entire DAG by $\tau$.
+But speeding up computations increases energy.
+Then, a natural question is, which cut brings the smallest energy increase?
+We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity.
 Then, our problem reduces to minimum cut, which we can solve with maximum flow.
-Appendix D has further details.
-Converting to GPU Frequencies.
-After finding the minimum capacity (minimum energy increase) cut, we will modify the durations of the computations involved in the cut, which results in a new energy schedule.
-Finally, we will convert the energy schedule into GPU frequencies that can be realized by the Perseus client.
-For each computation $i$, we convert its execution time $t$ to the slowest GPU frequency that will execute faster than $t$.
+After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining a new Pareto-optimal energy schedule.
+Appendix D provides details with mathematical expressions for flow capacities.
+Converting Back to GPU Frequencies.
+Finally, we convert the energy schedule into GPU frequencies that can be realized by the Perseus client.
+For each computation, we convert its planned execution time $t$ to the slowest GPU frequency that will execute faster than $t$.
 This is because when computations are tightly packed by our algorithm, while slightly speeding up a computation is acceptable, slowing down any computation on the critical path will directly slow down the entire DAG, increasing intrinsic energy bloat.
+API Description
+profiler.begin(type) Begin time and energy profiling for computation type.
+profiler.end(type) Record time and energy profiling results for computation type.
+controller.set_speed(type) Set the hardware’s execution speed as planned for computation type.
+server.set_straggler(id, delay, degree) Notify that a straggler is anticipated after delay seconds. A straggler returning to normal can be communicated by setting degree to 1.
+Table 2.
+The minimal set of Perseus client and server APIs that require implementation.
+One client process manages each accelerator.
+The type parameter should be either "forward" or "backward".
+On GPUs, the “speed” control knob is the SM frequency.
+The server API will issue an HTTP/RPC request to the server.
+The id parameter is the ID of an accelerator.
 Time Complexity Analysis.
 Our optimization algorithm has polynomial runtime.
-Let $N$ and $M$ respectively denote the number of stages and microbatches.
+Let $N$ and $M$ denote the number of stages and microbatches, respectively.
 Then, the computation DAG will have $O(NM)$ number of nodes and edges, and maximum flow with Edmonds-Karp runs in $O(N^{3}M^{3})$.
-While for general DAGs the total number of steps is known to be exponential to the size of the DAG [58], we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$.
+While for general DAGs the total number of steps is known to be exponential to the size of the DAG (skutella-thesis, ), we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$.
 See Appendix E for proof.
-In reality, commonly used number of stages ($N$) are 4 to 8 as too many increase pipeline bubble ratio [43, 14].
-The number of microbatches ($M$) is typically around $4N$ [26, 59], but recently with high data parallel degree, far less have been reported even for high-performance settings [14].
-As such, the algorithm runtime is negligible in practical scenarios (§6.5), given that large model training time is typically several weeks [48].
+In reality, commonly used number of stages ($N$) is 4 to 8 (at most tens) to reduce pipeline bubble ratio (megatronlm-sc21, ; mlperf_training_31, ).
+Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far less have been reported even for high-performance settings (mlperf_training_31, ).
+As such, algorithm runtime is practically negligible (§6.5), especially given that large model training easily takes weeks or months (patterson2021carbon, ).
 1
 2
-Input: DAG $\mathcal{G}$ with computations $i$
-Current energy schedule $p$
+3
+Input: DAG $\mathcal{G}$ of computations $i\in\mathcal{G}$
+Current Pareto-optimal energy schedule $p$
 Amount of iteration time to reduce $\tau$
-Output: Pareto-optimal schedule with reduced time $p^{\prime}$
-3
+Output: Pareto-optimal schedule w/ reduced time $p^{\prime}$
 4
-Function $\mathrm{ReduceTime}$($\mathcal{G}$, $p$, $\tau$):
+5
+6
 $\triangleright$ Construct edge-centric computation DAG ( 2)
-       $\mathcal{G}\leftarrow$ Split nodes into two and connect with edge $\mathcal{G}\leftarrow$ Crash bipartite node set to one node $\triangleright$ Remove non-critical computations ( 3)
-5       Annotate earliest & latest start times for $\forall i\in\mathcal{G}$ for $i\in\mathcal{G}$ do
-             $\triangleright$ Critical computations should have zero slack
+7 $\mathcal{G}\leftarrow$ Split nodes into two and connect with edge
+8
+$\triangleright$ Find and remove non-critical computations ( 3)
+9 Annotate earliest & latest start times for $\forall i\in\mathcal{G}$
+10 for $i\in\mathcal{G}$ do
 11       if $i$ has different earliest and latest start then
 12             Remove $i$ from $\mathcal{G}$
+13
 14       end if
-9
+15
 16 end for
+17
 $\triangleright$ Find set of computations to modify ( 4)
-       $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$) $\triangleright$ Modify computation durations ( 5)
-       $I\leftarrow\{i\mid i\textrm{ in }S-T\textrm{ cut}\}$ Modify duration of $\forall i\in I$ by $\tau$triangleright Assign frequencies from planned computation times
-11       $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned return $p^{\prime}$
-12
-13
-14
-Algorithm 2 ReduceTime: Reducing the execution time of a computation DAG by $\tau$ with minimal energy increase.
-4.4 Extensions
-In this section, we present extensions to our optimization algorithm useful for planning large model training.
-Tensor Parallelism.
-It is trivial to extend our algorithm to tensor parallelism, another essential ingredient of large model training.
-The observation is that tensor parallel techniques split operations in equal sizes.
-Therefore, GPUs that execute different portions of the same operation consume the same amount of time and energy.
-This allows Perseus to operate only on one tensor parallelism copy for each stage, decide the energy schedule for that copy, and replicate it to other copies in the same stage.
-We show that Perseus works well for hybrid parallelism in Section 6.4.
-Single-Choice Operations.
-Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latencies.
-For instance, especially on fast GPUs like A100, loading and copying input data into VRAM or communication over slower links can take considerable latency.
+18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$)
+19
+$\triangleright$ Modify computation durations ( 5)
+20 Modify duration of $\forall i$ in $S-T$ cut by $\tau$
+21
+$\triangleright$ Assign frequencies from planned computation times
+22 $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
+23
+24return $p^{\prime}$
+25
+26
+Algorithm 2 GetNextPareto: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
+4.4. Generalizations
+In this section, we present generalizations to our optimization algorithm useful for planning large model training.
+3D/Hybrid Parallelism.
+Operator parallelism techniques (e.g., data, tensor, or sequence parallelism) split operations in equal sizes, resulting in each GPU running the same computation.
+This allows Perseus to profile only one GPU per stage, decide the energy schedule for that GPU, and replicate it to all other GPUs in the same stage.
+We show that Perseus works well for 3D parallelism in Section 6.4.
+Constant-Time Operations.
+Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latency.
+For instance, loading and copying input data into VRAM or communication over slower links can take considerable latency.
 However, the time and energy consumption of these operations are not affected by the GPU’s frequency.
-Perseus can take the latency of such operations into account during planning by abstracting them as a node in the DAG with only one frequency choice.
+Perseus can take constant-time operations into account during planning by viewing them as a node with only one frequency choice.
 Other Pipeline Schedules.
-There are various schedules for pipeline parallel training, including GPipe [26], 1F1B [42], interleaved 1F1B [56], and early recomputation 1F1B [34].
+There are various schedules for pipeline parallel training, including GPipe (gpipe-neurips19, ), 1F1B (pipedreamflush-icml21, ), interleaved 1F1B (megatronlm-arxiv, ), and early recomputation 1F1B (merak-tpds23, ).
 As long as the computations on the schedule can be expressed as a DAG, Perseus can optimize its energy consumption without modification.
-If there is stage imbalance, we believe that any pipeline schedule will have intrinsic energy bloat.
-5 Implementation
-The Perseus server is implemented in 1,500 lines of Python [3]. The client is implemented in 300 lines of Python as a library that can be imported into training frameworks.
-As a reference, we have integrated the Perseus client with Merak [34], whose open source implementation takes the best of Megatron-LM [5] for high-performance tensor parallelism for transformer models and DeepSpeed [2] for its generic pipeline execution engine.
-Activation recomputation [12] is enabled by default to allow large batch sizes to fit in GPUs.
+As long as there is stage imbalance, any pipeline schedule will have intrinsic energy bloat.
+5. Implementation
+The Perseus server and client are implemented in Python.
+Perseus can optimize the energy consumption for any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented.
+As a reference, we have integrated the Perseus client with Merak (merak-tpds23, ), which marries high-performance tensor parallelism of Megatron-LM (megatronlm-github, ) and the generic pipeline execution engine of DeepSpeed (deepspeed-github, ).
+We provide an example of what is looks like to integrate the client with a training engine in Appendix F.
+Activation recomputation (checkpointing-arxiv16, ) is enabled to allow large batch sizes to fit in GPUs.
 Online Time/Energy Profiler.
-While accurate timing measurement is well-supported, modern GPUs (NVIDIA A40, A100, and H100) update their cumulative energy consumption counter [8] only once every 100 ms, which is similar to the timescale of one microbatch computation.
-This makes energy measurement noisy, and no ground truth is available.
-At the very moment the cumulative energy counter increases, its value should be accurate.
-Therefore, Perseus polls the energy counter in the background and records the timestamp of when the counter increased and its value.
-Then, collected (time, energy) points are interpolated with a linear line.
-The main training process only records the start and end timing of each computation, allowing Perseus to match timestamps with the interpolated energy curve to calculate each computation’s energy consumption.
-The GPU’s frequency is scanned from the highest to the lowest at iteration granularity.
-As an optimization, when computation energy increases for five consecutive frequencies, profiling is terminated.
+The client profiles the time and energy of each forward and backward computation in vivo at the beginning of the training job to take measurements representative of real training.
+While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs for measurement.
+The GPU’s frequency is profiled from the highest to the lowest at iteration granularity and stopped when energy consumption increases.
 Beyond that point, frequencies will consume more time and energy, making them suboptimal.
 Asynchronous Frequency Controller.
-The controller lives in a separate process that communicates with the main training process through a pipe, because setting the GPU’s frequency through NVML [8] should not block computation on the main process.
-While pipeline execution engine implementations differ widely, many have separate code blocks for forward and backward.
-Therefore, the controller exposes one method, set_frequency, which takes either "forward" or "backward" and sets the GPU’s frequency as planned.
-6 Evaluation
-We evaluate Perseus on five workloads and compare it against EnvPipe and Zeus. Our key findings are the following.
+The client-side controller spawns a separate process that asynchronously sets the GPU’s frequency through NVML (nvml, ) without blocking the main training process.
+Training frameworks can call set_speed at the beginning of forward or backward to set the GPU’s frequency as planned by the server.
+6. Evaluation
+We evaluate Perseus on five workloads and compare it against EnvPipe and Zeus. Our key findings are the following:
   • Perseus can effectively reduce intrinsic and extrinsic energy bloat. Training on real GPUs shows up to 28.5% energy savings using Perseus (§6.2).
-  • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings.
-We also observe that there is a tradeoff between energy savings and scale (§6.3).
+  • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings (§6.3).
   • Energy bloat reduction was possible because Perseus can enumerate efficient energy schedules on the “iteration time–energy” Pareto frontier (§6.4).
   • Perseus reduces energy bloat with low overhead (§6.5).
-6.1 Experimental Setup
+6.1. Experimental Setup
 Testbed.
 We run our evaluation workloads in a GPU cluster, where each node is equipped with an AMD EPYC 7513 CPU, 512 GB DRAM, and four NVIDIA A40-48G GPUs.
-For A100 results, we use the node provided by Chameleon Cloud [30], equipped with two Intel Xeon Platinum 8380 CPUs, 512 GB DRAM, and four NVIDIA A100-80G PCIe GPUs.
+For A100 results, we use a node provided by Chameleon Cloud (chameleoncloud, ), equipped with two Intel Xeon Platinum 8380 CPUs, 512 GB DRAM, and four NVIDIA A100-80G PCIe GPUs.
 Workloads and experiment parameters.
-We evaluate Perseus with various workloads spanning from GPT-3 [11], Bloom [66], BERT [15], T5 [51], to Wide-ResNet [69].
+We evaluate Perseus with various workloads spanning from GPT-3 (gpt3, ), Bloom (bloom-arxiv22, ), BERT (bert, ), T5 (t5-jmlr20, ), to Wide-ResNet (wideresnet-bmvc16, ).
 We use model variants with 1.3B to 6.7B parameters to run the models in our testbed, and scale them up to 176B parameters in large-scale emulation.
@@
 We use the minimum imbalance stage partitioning method described in Section 2.2 for all workloads.
-Appendix A lists complete model configurations, parameters, and stage partitioning details.
+Appendix B lists complete model configurations, parameters, and stage partitioning details.
+Metrics.
+We report energy reduction and slowdown of a training iteration (%), relative to using all maximum GPU frequencies.
+In most cases slowdown is close to zero, in which case energy and power reductions coincide.
+Reducing only extrinsic bloat is not possible, because Perseus reduces extrinsic bloat while keeping intrinsic bloat low as it slows down non-straggler pipelines.
+Therefore, we report (1) intrinsic bloat reduction without stragglers and (2) intrinsic + extrinsic bloat reduction with stragglers.
 Baselines.
-We mainly compare against two prior works:
-  • EnvPipe [13] reduces only intrinsic energy bloat while trying to minimize slowdown in pipeline execution.
-We compare Perseus’s amount of energy bloat reduction against that of EnvPipe (§6.2, §6.3).
-  • Zeus [68] characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
-6.2 Reducing Energy Bloat
-We start with overall energy bloat reduction – both intrinsic (§6.2.1) and extrinsic (§6.2.2) – achieved by Perseus and EnvPipe by running various models on our testbed.
-We compare Perseus against EnvPipe, which reduces energy consumption while trying to minimize slowdown.
-Both solutions use the same amount of GPU resources.
-6.2.1 Intrinsic Energy Bloat Reduction
+We mainly compare with two prior works:
+  • EnvPipe (envpipe-atc23, ) reduces only intrinsic energy bloat while trying to minimize slowdown. We compare Perseus’s energy bloat reduction with EnvPipe (§6.2, §6.3).
+  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
+6.2. Reducing Energy Bloat
+We start with overall energy bloat reduction – intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2) – achieved by Perseus and EnvPipe.
+All numbers were obtained by running on testbed GPUs, and all solutions use the same amount of GPU resources.
+6.2.1. Intrinsic Bloat Reduction Without Stragglers
 Model Energy Savings (%) Slowdown (%)
@@
 Wide-ResNet 1.5B 3.2 3.7 2.3 4.1
-[…]
+(a) Four stage pipeline parallelism on A100 GPUs
 Model Energy Savings (%) Slowdown (%)
@@
 Wide-ResNet 1.5B 20.4 16.5 0.2 0.5
-[…]
-Table 2: Intrinsic energy reduction and slowdown comparison between Perseus and EnvPipe.
+(b) Eight stage pipeline parallelism on A40 GPUs
+Table 3. [Experiment] Intrinsic energy bloat (without stragglers) reduction and iteration time slowdown.
 Table 3 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the “iteration time–energy” frontier) and that by EnvPipe.
 We make two observations regarding Perseus.
-First, models show varying amounts of energy savings because their stage imbalance vary (Table 1).
-For instance, unlike other models, Wide-ResNet 1.5B on A100 after minimum imbalance stage partitioning has nearly perfect stage balance, leaving little room for non-critical computations to slow down.
+First, models show varying amounts of energy savings because (1) their stage imbalance vary (Table 1), and (2) their forward and backward are composed on different computations, which affects time/energy sensitivity when changing the frequency.
 Second, A40 demonstrates more energy savings compared to A100.
@@
 Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency.
-However, we expect the upcoming NVIDIA H100 GPU to have better savings because its maximum frequency is 1755 MHz for the PCIe version and 1980 MHz for SXM [44].
+However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
 EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest, which is not always true.
 Additionally, it sometimes considerably degrades iteration time because it is not aware of single-choice operations inside the pipeline (§4.4) and can slow down some computations too much.
-6.2.2 Intrinsic + Extrinsic Energy Bloat Reduction
+6.2.2. Intrinsic + Extrinsic Bloat Reduction With Stragglers
 Model Method Energy Savings (%) given $T^{\prime}/T$
@@
 1.5B EnvPipe 4.9 4.8 4.7 4.5 4.4 4.3
-[…]
+(a) Four stage pipeline parallelism on A100 GPUs
 Model Method Energy Savings (%) given $T^{\prime}/T$
@@
 1.5B EnvPipe 16.4 16.2 15.8 15.4 15.0 14.6
-[…]
-Table 3: Energy savings of Perseus given varying degrees of straggler slowdown ($T^{\prime}/T$).
-EnvPipe cannot provide additional savings because it does not consider and cannot be applied to optimize extrinsic bloat.
-When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its time–energy frontier yields for longer iteration times.
-Table 3 shows the amount of energy savings given varying degrees of straggler slowdown.
-As described in Section 3.1, non-stragglers’ iteration time is set to be $\min(T^{*},T^{\prime})$, and energy is reduced from (1) slowing down the pipeline itself and (2) blocking on communication for a shorter amount of time, waiting for the straggler.
-Slowing down the non-stragglers beyond $T^{*}$ increases total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
-The percentage of energy savings increases for lower $T^{\prime}$ values and slowly wanes when $T^{\prime}>T^{*}$.
-This reduction is due to longer blocking time.
-That is, the absolute amount of energy savings in Joules is the largest at $T^{*}$, and constant afterward.
-However, energy consumption during blocking time ($T^{\prime}-T^{*}$) increases for larger $T^{\prime}$ values, lowering the percentage of energy savings for $T^{\prime}>T^{*}$.
+(b) Eight stage pipeline parallelism on A40 GPUs
+Table 4. [Experiment] Energy savings given varying straggler slowdown ($T^{\prime}/T$). Perseus can reduce extrinsic bloat while keeping intrinsic bloat low, whereas EnvPipe cannot.
+When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its “iteration time–energy” Pareto frontier yields for longer iteration times.
+Table 4 shows the amount of energy savings of a pipeline given varying degrees of straggler slowdown.
+As described in Section 3.1, a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$, and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time blocking on communication, waiting for the straggler.
+Slowing down the non-stragglers beyond $T^{*}$ will increase total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
+The percentage of savings first increases with the straggler’s iteration time, but then slowly wanes as it slows down beyond $T^{*}$.
+This is expected.
+The absolute amount of energy reduction in Joules is the largest when the non-straggler pipeline’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$.
+Thus, as the straggler slows down beyond $T^{*}$, the time and energy consumed while waiting for the straggler increases, lowering the percentage of energy savings.
 Finally, the point of maximum energy savings is different for each model.
-This is because different models have different $T^{*}$ values, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
-6.3 Large-Scale Emulation
+This is because each model has a different $T^{*}$ value, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
+6.2.3. How Much Potential Savings Was Realized?
+The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the “iteration time–energy” Pareto frontier).
+For intrinsic bloat without stragglers, Perseus on average realizes 74% and 89% the opportunity on A100 and A40, respectively, with negligible slowdown.
+This is possible because there are much more non-critical computations in the DAG that can be slowed down than critical ones.
+With stragglers, Perseus fully realizes potential savings when the straggler’s slowdown degrees are on average 1.1 and 1.15 on A100 and A40 respectively, which is not unrealistic considering slowdowns reported in literature (§2.3).
+6.3. Large-Scale Emulation
 Because we do not have access to a GPU cluster required to run huge models like GPT-3 175B, we use emulation grounded on fine-grained profiling for large-scale evaluation.
-Emulation Methodology.
-Total
+In general, trends in our emulation result match those obtained from real training in Section 6.2.
 # GPUs
@@
 # Microbatches
-per Pipelines
+Per Pipeline
 Global
@@
 8192 128 12
-Table 4: Experiment parameters for GPT-3 175B and Bloom 176B. Each pipeline has tensor parallel degree 8 and 8 pipeline stages.
-We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B in bfloat16 to construct the time and energy profile of each stage.
-Then, we run our optimization algorithm to obtain a theoretical “iteration time–energy” frontier, and use it to report emulated savings.
-We evaluate the changes in the amount of energy savings in a strong scaling setup presented in Table 4.
-We do not consider weak scaling, where per-pipeline batch is constant and increasing the global batch size proportionally, because varying the global batch size can affect the model quality [31, 21].
-We find that, compared to actual measurements of smaller scale models on A100, the emulator always underestimates the actual percentage of energy savings by 20.2% on average (see Appendix F for details).
-Meaning, the savings reported by our emulation can likely be considered an empirical lower bound for actual savings.
-We used A100 SXM GPUs for the A100 results in this section, which is more representative of large scale scenarios (Appendix G provides PCIe results).
+Table 5. Strong scaling parameters for large-scale emulation. A pipeline has tensor parallel degree 8 and 8 pipeline stages.
+Emulation Methodology.
+We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the “iteration time–energy” frontier.
+We perform strong scaling when varying the number of GPUs (Table 5) in order to keep the global batch size constant (largebatch1, ; largebatch2, ).
+We used A100 SXM GPUs for emulation, which we believe are more representative of large-scale training infrastructure.
+Emulator Fidelity.
+We compare the percentage of energy savings vs.
+running all maximum frequencies given by our emulator and real experiments for our A100 workloads and find that the emulator always underestimates energy savings.
+Specifically, savings on the leftmost and rightmost point of the Pareto frontier are underestimated by 18.6% and 21.7% on average, respectively.
+We believe this is due to our simplifying assumption that $P_{\mathrm{blocking}}$ is constant regardless of the GPU’s frequency.
+This means savings given by the emulator can be considered a lower bound for actual savings.
+Figure 7. [Emulation] Energy savings breakdown with straggler slowdown 1.2 and 1,024 GPUs.
 Results Summary.
-Figure 7: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1,024 GPUs.
-Figure 7 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1,024 GPUs as a representative.
-EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it reduces less than Perseus.
-In contrast, Perseus reduces per-iteration energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
-Intrinsic Bloat Reduction.
+Figure 7 breaks down the amount of energy bloat reduction for GPT-3 175B and Bloom 176B when slowdown degree is 1.2 on 1,024 GPUs.
+EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it is suboptimal.
+In contrast, Perseus reduces energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
 Model GPU Type
@@
 A40 6.97 4.49 3.12 2.41
-Table 5: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
-Table 5 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches.
+Table 6. [Emulation] Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B. Number of microbatches is varied following Table 5.
+Intrinsic Bloat Reduction Without Stragglers.
+Table 6 shows Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B.
+The number of microbatches was varied based on Table 5.
 For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases.
-This is fundamentally due to the ratio of microbatches in the pipeline’s warm-up and flush phase (beginning and end) vs.
+This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) vs.
 steady state phase (middle).
-Microbatches in the former phase are able to slow down until their minimum energy frequency, yielding large energy savings.
-However, microbatches in the latter (middle of the pipeline) cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings.
-When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction becomes more and more dominated by the average energy savings of steady state microbatches.
-Extrinsic Bloat Reduction.
-Figure 8: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines.
-When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$, whose location is denoted by a star.
-More blocking time after $T^{*}$ decreases overall energy savings.
+Most microbatches in the warm-up and flush phases can slow down until their minimum energy frequency, yielding large energy savings.
+However, microbatches in the pipeline’s steady state cannot slow down to their full potential when the amount of stage imbalance is not large, thereby yielding modest savings.
+When the number of microbatches in the pipeline increases, only the number of steady state microbatches increases, and energy reduction converges to the average energy savings of steady state microbatches.
+Figure 8.
+[Emulation] Perseus’s intrinsic + extrinsic energy bloat reduction with varying straggler slowdown ($T^{\prime}/T$).
+Number of pipelines is varied following Table 5.
+$\bigstar$ denotes $T^{*}/T$ for each pipeline.
 Please note the different Y-axes.
-We simulate stragglers in training large models with hybrid parallelism in various strong scaling configurations (Table 4), where a straggler is slower than others in varying degrees from 1.05 to 1.50.
-Other non-stragglers, after finishing their computation, must wait until the straggler finishes computation, where extrinsic energy bloat emerges.
-We measure the amount of energy bloat and Perseus’ extrinsic energy savings.
-Figure 8 shows the changes of extrinsic energy bloat reduction given varying degrees of straggler slowdown and various strong scale configurations.
-The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward, is consistent with what was observed in Section 6.2.2.
-An interesting observation here is that there is a tradeoff between scale and energy savings: more pipelines have less percentage of energy savings or less amount of energy savings per pipeline.
-It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline that cannot be optimized while all the other pipelines optimize their energy consumption.
-However, this holds only in weak-scaling configuration, i.e., per-pipeline batch size is constant (increasing the global batch size proportionally to the number of pipelines) and it is not the case in strong-scaling configuration, where the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed.
-With less number of microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration, which cannot be eliminated by intrinsic energy bloat reduction, increases, resulting in less energy savings.
-6.4 Iteration Time–Energy Frontier Comparison
-Figure 9: Iteration time–energy frontiers for GPT-3, achieved by Perseus and the two baselines derived from Zeus [68].
+Intrinsic + Extrinsic Bloat Reduction With Stragglers.
+We introduce stragglers in large-scale emulation in various strong scaling configurations (Table 5) and with varying straggler slowdown degree.
+Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus.
+The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
+An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline.
+It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption.
+However, this holds only in weak scaling configuration, where per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally to the number of pipelines).
+Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes.
+With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ).
+These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy saving percentages.
+However, as the absolute number of GPUs increase, even a small savings percentage gives huge absolute energy savings.
+6.4. Iteration Time–Energy Frontier Comparison
+Figure 9.
+[Experiment] Iteration time–energy frontiers for GPT-3, achieved by Perseus and the two baselines derived from Zeus (zeus-nsdi23, ).
 Perseus Pareto-dominates all other approaches.
 The dotted vertical line is the iteration time of running all GPUs at their maximum power limit, which is the default mode of operation.
-Please note the different Y-axes.
-The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm.
-Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus [68], which is an energy optimization framework for a single-GPU training with the training time–energy frontier.
-We implemented two Zeus-based baselines to make it work in parallelization configurations and generated “iteration time–energy” frontiers.
-  1. 1.
+Please note the different X- and Y-axes.
+The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm (§4).
+Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus (zeus-nsdi23, ).
+Since Zeus only produces the training time–energy frontier for single-GPU training jobs, we implemented two Zeus-based baselines for large model training scenarios:
   • ZeusGlobal: Scans one global power limit for all stages.
-  2. 2.
-    ZeusPerStage: Sets one power limit per stage that balances forward computation time.
-We run training and measurement under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) hybrid parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40.
-Figure 9 shows the frontiers of Perseus and Zeus for different sizes of GPT-3 under three parallelization configurations.
-Frontiers for other models are in Appendix H.
+  • ZeusPerStage: Finds a set of per-stage power limits that balances forward computation time.
+Figure 9 shows the frontiers of all solutions for different sizes of GPT-3 under three parallelization configurations: (a) four stage pipeline parallelism on A100; (b) eight stage pipeline parallelism on A40; and (c) 3D parallelism (data parallelism 2, tensor parallelism 2, pipeline parallelism 4) on A40.
+All results were obtained from running on testbed GPUs.
+Appendix G has frontiers for other models.
 Perseus Pareto-dominates Zeus.
-ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage equally, unable to reduce intrinsic energy bloat.
-While ZeusPerStage can balance the forward computation time of each stage, it is unaware of the critical path of the DAG, slowing down critical computations.
-In contrast, Perseus can precisely slow down non-critical computations, tightly packing computation.
-6.5 Overhead of Perseus
+ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, unable to reduce intrinsic energy bloat.
+ZeusPerStage can balance the forward computation time of each stage, but is unaware of the critical path of the DAG, slowing down critical computations.
+In contrast, Perseus can precisely slow down non-critical computations, effectively reducing energy bloat.
+6.5. Overhead of Perseus
 Profiling.
-Our online profiling method (§5) introduces extra training time by running some iterations at a lower frequency.
-For our A100 workloads, the average percentage of slowdown for such iterations was 8.2%, and the average extra training time introduced by profiling was 13 minutes.
-This is negligible compared to how long large model training can take.
+In the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage.
+For our A100 workloads, the initial profiling phase increased end-to-end training time by 13 minutes on average, which is negligible overhead for large model training.
 Algorithm Runtime.
-The average time spent on running Perseus’s optimization algorithm (§4) across the five A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes).
-For our largest scale emulation experiment (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds.
-While the runtime of the algorithm will increase with larger DAGs for larger models, we believe the overhead be justified because training time also increases with the scale of the training job.
-Finding the energy-optimal iteration time and energy schedule given the straggler’s iteration time for extrinsic energy bloat reduction is instant.
-7 Related Works
+The average runtime of the optimization algorithm (§4) across our A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes).
+For our largest scale emulation (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds, which is shorter because it is sufficient to optimize just one data parallel copy (§4.4).
+While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time also increases with the scale of the training job.
+Looking up the optimal energy schedule given $T^{\prime}$ is instant.
+7. Related Works
 Large Model Training.
-Many recent works focus on enabling and accelerating large model training using 3D parallelism (data, tensor, and pipeline parallelism).
-GPipe [26] and PipeDream [41] were the first to introduce pipeline parallelism and explicitly discussed the difficulty of perfectly balancing computation time across stages.
-Megatron-LM [56, 43] is a Transformer-based large model training framework that provides efficient manually designed execution plans.
-Later on, modern training frameworks that focus on large scale training [53, 59, 36, 37, 40, 34] were introduced to support various models at scale.
-DeepSpeed [53] introduced ZeRO redundancy optimizer that shards model states, which is recently widely used for large model training [52, 72].
-Alpa [73] and GSPMD[70] are automatic parallelization framework for general DNNs.
-Finally, some recent works have looked into fault-tolerant large model training as well [61, 27, 9].
-Unfortunately, energy consumption is not an optimization metric for any of the major large model training frameworks.
+Many recent works focus on enabling and accelerating large model training using 3D parallelism.
+GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) were the first to introduce pipeline parallelism.
+Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ) is a large model training framework that provides efficient 3D parallel execution plans tailored for Transformer models.
+3D parallelism is considered to be go-to solutions for modern large model development due to strong open-source implementations (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively simple implementation complexity.
+Finally, Alpa (alpa-osdi22, ) and GSPMD (spmd_socc22, ) provide automatic parallelization for general DNNs.
+However, energy consumption is not an optimization metric for any of the major large model training frameworks.
 DNN Training and Energy Consumption.
-A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training, including those that present observations and estimations [48, 60, 33, 16, 38] and those that propose optimization methods for training time, energy consumption, and carbon footprint [74, 63, 68, 67, 32, 13].
-Zeus [68] is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on simple single-GPU training.
-EnvPipe [13], on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown.
-However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings and an infinite loop during optimization.
-Perseus Pareto-dominates both Zeus and EnvPipe by viewing large model training as a computation DAG and introducing a principled optimization algorithm.
-Perseus is also the first to introduce the notion of extrinsic energy bloat in large model training and optimize both simultaneously.
-8 Conclusion
-We presented Perseus, an energy optimization system for large model training.
-Perseus builds on top of observation that there are fundamental computation imbalance at different levels in large model training that causes intrinsic and extrinsic energy bloat.
-We introduced a principled graph cut-based algorithm that simultaneously reduces both.
-Perseus advances the state-of-the-art of DNN training energy optimization by establishing a new time–energy Pareto frontier for large model training.
-The integration of Perseus into the training workflow has strong implications for the future of AI development.
-Importantly, reducing energy bloat leads to practically no latency and throughput degradation, which has the potential to greatly enhance the sustainability of distributed training in the midst of recent proliferation of LLMs and GenAI.
-Acknowledgements
-We thank Chameleon Cloud [30] for providing A100 nodes.
-Appendix A Workload Details
-A.1 Minimum Imbalance Pipeline Partitioning
+A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ) and proposed optimization methods (odpp, ; gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; depo, ; envpipe-atc23, ).
+Especially, Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on single-GPU training.
+EnvPipe (envpipe-atc23, ), on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown.
+However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings.
+Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm.
+Perseus is also the first to introduce the notion of energy bloat.
+8. Conclusion
+We presented Perseus, a software-based energy optimization system for large model training.
+Perseus builds on the observation that there are computation imbalances at different levels in large model training that causes intrinsic and extrinsic energy bloat, and introduces a principled graph cut-based algorithm that simultaneously reduces both.
+As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby reducing average power draw as well.
+Appendix A Visualizations for Intrinsic Energy Bloat
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Refer to captionRefer to caption
+Figure 10.
+Visualization of Perseus’s $T_{\min}$ solution for four stage pipeline workloads on NVIDIA A100 PCIe GPUs.
+For each workload, the left is running every computation at maximum frequency, and the right is Perseus’s energy schedule that reduces only intrinsic energy bloat without inflating iteration time.
+Note that these are not real workloads we run in Section 6; real workloads have far more microbatches.
+Figure 10 shows the timeline of running one training iteration of BERT 1.3B, T5 3B, Bloom 3B, and Wide-ResNet101 1.5B for maximum frequency and Perseus-optimized energy schedule, respectively.
+For visualization purposes, we set the number of microbatches to 6.
+Real evaluation workloads have more microbatches.
+Energy schedule found by Perseus successfully tunes down frequency for all models without slowing down the iteration time, tightly packing computations over time and reducing intrinsic energy bloat.
+Appendix B Workload Details
+B.1. Minimum Imbalance Pipeline Partitioning
 Model Size Imbalance Ratio Minimum Imbalance Ratio Partition
@@
 Wide-ResNet101 [74] 1.5B 1.09 1.25 [0, 8, 17, 26, 35] [0, 4, 8, 12, 16, 21, 26, 31, 35]
-[…]
+(a) NVIDIA A100 PCIe GPUs.
 Model Size Imbalance Ratio Minimum Imbalance Ratio Partition
@@
 Wide-ResNet101 [74] 1.5B 1.08 1.25 [0, 8, 17, 26, 35] [0, 4, 8, 12, 16, 21, 26, 31, 35]
-[…]
-Table 6: Imbalance ratio between the longest and the shortest stages for various models.
+(b) NVIDIA A40 GPUs.
+Table 7.
+Imbalance ratio between the longest and the shortest stages for various models.
 1.00 would mean perfect balance.
@@
 The vocabulary size of GPT-3 is 50k, Bloom 251k, and BERT 31k, which results in a very large linear layer for the last stage.
-This leads to varying amounts of imbalance and different minimum imbalance partitioning results for each models.
+This leads to varying amounts of imbalance and different minimum imbalance partitioning results for each model.
 T5.
@@
 Therefore, even with minimum imbalance partitioning, it is difficult to perfectly balance stages.
-A.2 Experiment Parameters
+B.2. Alternative Planning Methods
+3D parallelism is the go-to solution for large model training, and our minimum imbalance pipeline partitioning method is optimal for 3D parallelism because we implemented a brute force search.
+More advanced planning methods such as Alpa [77] exist, but for repetitive language model architectures like GPT-3, Alpa equally allocates Transformer layers to each stage [77], resulting in the same stage imbalance caused by the language modeling head.
+Furthermore, it is hard for tensor parallelism to divide operations infinitely; practically at most degree 8 (within one node) due to collective communication overhead.
+Thus, in the vast majority of models, there will always be some degree of imbalance between stages due to the minimum granularity of computation time.
+B.3. Experiment Parameters
 Model # Parameters Global Batch Size Microbatch Size # Microbatches
 gpt3-6.7b 6.7 B 1024 4 128
-Table 7: Experiment Parameters for 3D parallelism experiments on A40 GPUs.
+Table 8.
+Experiment Parameters for 3D parallelism experiments on A40 GPUs.
 Microbatch size is per-pipeline, and there are two data parallel copies of the same pipeline.
@@
 wide-resnet101 (width factor 8) 1.5 B 1536 32 48
-Table 8: Experiment Parameters for eight-stage pipeline parallelism experiments on A40 GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
+Table 9. Experiment Parameters for eight-stage pipeline parallelism experiments on A40 GPUs. Model variant names are as described in Torch Vision [49] or Huggingface Hub [68].
 Model # Parameters Global Batch Size Microbatch Size # Microbatches
@@
 wide-resnet101 (width factor 8) 1.5 B 1536 64 24
-Table 9: Experiment Parameters for Pipeline Parallelism Experiments on A100 PCIe GPUs. Model variant names are as described in Torch Vision [45] or Huggingface Hub [65].
-Tables 78, and 9 list model variant names and experiment parameters for our experiments.
+Table 10. Experiment Parameters for Pipeline Parallelism Experiments on A100 PCIe GPUs. Model variant names are as described in Torch Vision [49] or Huggingface Hub [68].
+Tables 89, and 10 list model variant names and experiment parameters for our experiments.
 Model names and configurations for GPT-3 were taken as is from the original model publication [9].
@@
 The unit time parameter $\tau$ was set to 1 ms for all experiments.
-Appendix B Pipeline Energy Minimization is NP-Hard
-The original Pipeline Energy Minimization problem is stated in formal terms:
-$\min_{F}\quad$ $\textrm{Energy}(F)$ (6)
+Appendix C Pipeline Energy Minimization is NP-hard
+The Pipeline Energy Minimization problem is stated again for convenience:
+(6) $\min_{F}$ $\textrm{Energy}(F)$
 s.t. $\textrm{Time}(F)\leq T^{\prime}$
@@
 In the following, we show that a simplification of pem is NP-hard by reduction from the 0/1 Knapsack problem, which makes pem NP-hard.
-B.1 One Stage Two Frequencies Simplification
+C.1. One Stage Two Frequencies Simplification
 A simplification of pem is considering the case where there is only one pipeline stage and two frequencies to choose from.
@@
 pem-1d returns true if and only if there exists a subset of operations $J\subseteq\{1,2,\ldots,n\}$ such that $\sum^{n}_{i=1}(X_{i}t_{i}(L)+(1-X_{i})t_{i}(H))\leq T^{\prime}$ and $\sum^{n}_{i=1}(X_{i}e_{i}(L)+(1-X_{i})e_{i}(H))\leq EC$.
-B.2 0/1 Knapsack Problem
+C.2. 0/1 Knapsack Problem
 Consider two length $n$ arrays containing positive integer weights $W=(w_{1},w_{2},\ldots,w_{n})$ and values $V=(v_{1},v_{2},\ldots,v_{n})$ where the $i$th item has weight $w_{i}\in\mathbb{Q}^{+}$ and value $v_{i}\in\mathbb{Q}^{+}$, and a knapsack with weight capacity $C\in\mathbb{Q}^{+}$.
@@
 It is well known that knapsack is NP-hard.
-B.3 NP-Hardness Proof
-Theorem 2.
+C.3. NP-hardness Proof
+Theorem C.1.
 pem-1d is NP-hard.
@@
 […]
-  1. 1.
+  1. (1)
     Construct $n$ computations and empty vectors $T_{L},T_{H},E_{L},E_{H}$.
-  2. 2.
+  2. (2)
     For $\forall i$, set $t_{i}(L)=w_{i}$ and append to $T_{L}$.
-  3. 3.
+  3. (3)
     For $\forall i$, set $t_{i}(H)=0$ and append to $T_{H}$.
-  4. 4.
+  4. (4)
     For $\forall i$, set $e_{i}(L)=-v_{i}$ and append to $E_{L}$.
-  5. 5.
+  5. (5)
     For $\forall i$, set $e_{i}(H)=0$ and append to $E_{H}$.
-  6. 6.
+  6. (6)
     Set $T^{\prime}$ = $C$ and $EC$ = $-P$.
-  7. 7.
+  7. (7)
     Output $(T=(T_{L},T_{H}),E=(E_{L},E_{H}),T^{\prime},EC)$
@@
 Therefore, the function $f$ takes $O(n)$ time, which is polynomial time w.r.t the input size.
-Appendix C Visualizations for Intrinsic Energy Bloat
-Figure 10: Visualization of Perseus’s solution for four stage pipeline workloads on NVIDIA A100 PCIe GPUs.
-For each workload, the left is running every computation at maximum frequency, and the right is Perseus’s energy schedule that reduces only intrinsic energy bloat without inflating iteration time.
-Note that these are not real workloads we run in Section 6; real workloads have far more microbatches.
-Figure 10 shows the timeline of running one training iteration of BERT 1.3B, T5 3B, Bloom 3B, and Wide-ResNet101 1.5B for maximum frequency and Perseus-optimized energy schedule, respectively.
-For visualization purposes, we set the number of microbatches to 6.
-Real evaluation workloads have more microbatches.
-Energy schedule found by Perseus successfully tunes down frequency for all models without slowing down the iteration time, tightly packing computations over time and reducing intrinsic energy bloat.
-Appendix D Full Details of ReduceTime
+Appendix D Full Details of GetNextPareto
 For the sake of presentation, we made a simplifying statement in the body of the paper that computations only speed up by $\tau$.
@@
 In the following, we describe the procedure of annotating edges with flow capacities and then solving the problem with maximum flow.
-D.1 Generating Capacity DAG
+D.1. Generating Capacity DAG
 On the computation DAG, we first remove all the computations that are not on any of the critical paths and construct the Critical DAG.
@@
 This can be described as the following problem:
-$\min_{I^{+},\leavevmode\nobreak\ I^{-}}\quad\sum_{i\in I^{+}}{e_{i}^{+}}-\sum_{i\in I^{-}}{e_{i}^{-}},$ (7)
-where $e_{i}(t_{i})$ is the exponential function fit to Pareto-optimal (time, energy) measurments for computation $i$, $e_{i}^{+}=e_{i}(t_{i}-\tau)-e_{i}(t_{i})$ is the extra amount of energy needed to speed up $i$ by $\tau$ and $e_{i}^{-}=e_{i}(t_{i})-e_{i}(t_{i}+\tau)$ is the energy saved from slowing down $i$ by $\tau$.
+(7) $\min_{I^{+},~{}I^{-}}\quad\sum_{i\in I^{+}}{e_{i}^{+}}-\sum_{i\in I^{-}}{e_{i}^{-}},$
+where $e_{i}(t_{i})$ is the exponential function fit to Pareto-optimal (time, energy) measurements for computation $i$, $e_{i}^{+}=e_{i}(t_{i}-\tau)-e_{i}(t_{i})$ is the extra amount of energy needed to speed up $i$ by $\tau$ and $e_{i}^{-}=e_{i}(t_{i})-e_{i}(t_{i}+\tau)$ is the energy saved from slowing down $i$ by $\tau$.
 An important fact that leads us to the solution is that (i) the problem of finding the set of edges to modify such that energy increase is minimized (i.e., solving Equation 7) coincides with (ii) finding the minimum cut of a DAG whose lower and upper bound flow capacities are defined as
-$(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$ (8)
+(8) $(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$
 This equivalence was discovered by Phillips and Dessouky [53, 27].
@@
 Finally, after the minimum cut has been identified from the Capacity DAG, edges in $I^{+}$ are sped up by $\tau$ and those in $I^{-}$ are slowed down by $\tau$, ultimately reducing the length of every critical path exactly by $\tau$ with the smallest possible energy increase.
-D.2 Max Flow Algorithm on the Capacity DAG
+D.2. Max Flow Algorithm on the Capacity DAG
 A characteristic of our Capacity DAG that precludes the direct application of well-known max flow algorithms is that some edges also have flow lower bounds, asserting that at least a certain amount of flow must pass through in the edge.
@@
 2
+3
+4
 Input: Directed Acyclic Graph $G=(V,E)$
@@
 Output: A maximum feasible flow of $G$ if it exists
-3
-4
-Construct a new graph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding new source and sink nodes $s^{\prime}$ and $t^{\prime}$, edges from $s^{\prime}$ to each node in $V$, edges from each node in $V$ to $t^{\prime}$, and an edge from $t$ to $s$triangleright Define the capacity $c^{\prime}(e)$ of each edge $e\in E^{\prime}$
-5: for $v\in V$ do
-6       $c^{\prime}(s^{\prime}v)\leftarrow\sum_{u\in V}l(uv)$ $c^{\prime}(vt^{\prime})\leftarrow\sum_{w\in V}l(vw)$
+5
+6
+7
+8Construct a new graph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding new source and sink nodes $s^{\prime}$ and $t^{\prime}$, edges from $s^{\prime}$ to each node in $V$, edges from each node in $V$ to $t^{\prime}$, and an edge from $t$ to $s$
+9
+$\triangleright$ Define the capacity $c^{\prime}(e)$ of each edge $e\in E^{\prime}$
+10: for $v\in V$ do
+11       $c^{\prime}(s^{\prime}v)\leftarrow\sum_{u\in V}l(uv)$
+12       $c^{\prime}(vt^{\prime})\leftarrow\sum_{w\in V}l(vw)$
+13
 14 end for
-8for $uv\in E$ do
+15for $uv\in E$ do
 16       $c^{\prime}(uv)\leftarrow u(uv)-l(uv)$
+17
 18 end for
-$c^{\prime}(ts)\leftarrow\infty$triangleright Find the max flow on G’
-$f^{\prime}\leftarrow$ EdmondsKarp ($G^{\prime}$, $c^{\prime}$) $\triangleright$ $G$ has a feasible flow if and only if $G^{\prime}$ has a saturating flow
+[…]
+20
+$\triangleright$ Find the max flow on G’
+21 $f^{\prime}\leftarrow$ EdmondsKarp ($G^{\prime}$, $c^{\prime}$)
+22
+$\triangleright$ $G$ has a feasible flow if and only if $G^{\prime}$ has a saturating flow
 23 if FlowValue ($f^{\prime}$) $\neq\sum_{v\in V}{c^{\prime}(s^{\prime}v)}$ then
 24       return nil
+25
 26 end if
+27
 $\triangleright$ Convert $f^{\prime}$ to a feasible flow $f$ in $G$
@@
 29       $f(uv)\leftarrow f^{\prime}(uv)+l(uv)$
+30
 31 end for
+32
 $\triangleright$ Construct residual graph and improve $f$ to max flow
 33 for $uv\in E$ do
-18       $c(uv)\leftarrow u(uv)-f(uv)$ $c(vu)\leftarrow f(vu)-l(vu)$
+34       $c(uv)\leftarrow u(uv)-f(uv)$
+35       $c(vu)\leftarrow f(vu)-l(vu)$
+36
 37 end for
-20return EdmondsKarp ($G$, $c$)
-21
+38return EdmondsKarp ($G$, $c$)
+39
+40
+41
 Algorithm 3 Maximum Flow with Lower Bounds
@@
 The assumptions are valid for all pipeline schedules known to the authors, including GPipe [29] and 1F1B [46].
-Theorem 3.
+Theorem E.1.
 For DAGs that represent pipeline schedules, the number of iterations needed is $O(N+M)$.
@@
 Therefore, the total number of pipeline computations in the critical path of the entire DAG is $O(N+M)$.
-$t_{\max}$ and $t_{\min}$ can be constructed by multiplying the number of computations with the average execution time of an computation.
+$t_{\max}$ and $t_{\min}$ can be constructed by multiplying the number of computations with the average execution time of a computation.
 Computations are executed with frequencies $f_{\min}$ and $f_{\max}$, respectively, and thus the multipliers of $N$ and $M$ do not cancel out when $t_{\max}-t_{\min}$ is evaluated.
@@
 ∎
-Appendix F Underestimations of the Emulator
-Model Leftmost point Rightmost point
-gpt3-xl 20.5% 23.6%
-bert-huge-uncased 26.7% 35.3%
-t5-3b 19.1% 19.1%
-bloom-3b 9.3% 10.7%
-wide-resnet101 (width factor 8) 17.5% 19.8%
-Table 10: By how much the emulator underestimates the percentage of energy savings for four-stage pipeline workloads on A100 PCIe.
-Leftmost point represents intrinsic energy bloat, whereas the rightmost point extrinsic.
-Our emulator, used to present energy saving results for large scale workloads (GPT-3 175B and Bloom 176B), always underestimates energy savings compared to actual measurements, and argue that the emulated results are empirically a lower bound of actual savings.
-The characteristics of time and energy changes depends more on the type of the GPU than the parallelization setting as different GPU generations show drastically different sensitivity in terms of time and energy when frequency is varied.
-Therefore, we compare results from the A100 PCIe GPU.
-Table 10 shows the percentage of energy savings underestimation for two points on the “iteration time–energy” frontier:
-  • Leftmost point: The point with the shortest iteration time, corresponding to how much the emulator underestimates the amount of possible intrinsic energy bloat reduction.
-  • Rightmost point: The point with the longest iteration time, corresponding to how much the emulator underestimates the amount of possible energy savings brought about by slowing down the entire pipeline in the case of a straggler.
-We believe that underestimation is happening due to our simplifying assumption that $P_{\mathrm{blocking}}$, the power consumption of the GPU while blocking on communication, is constant regardless of the GPU’s frequency.
-In reality, $P_{\mathrm{blocking}}$ would be an increasing function of frequency.
-Therefore, when the emulator assumes a constant $P_{\mathrm{blocking}}$ when calculating the energy consumption of a pipeline that is overall running on a reduced frequency, it will be overestimating its total energy consumption, leading to underestimations.
-Appendix G Large-Scale Emulation
-Section 6.3 used A100 SXM GPUs, which we momentarily had access to, in order to profile model layers and run large scale emulation.
-This is because we believe that SXM models are more representative of real large scale GPU clusters that train large models.
-Nonetheless, in this section, we provide the same set of results from A100 PCIe GPUs, as well, along with the same A40 GPUs results.
-All other experiment methodology and parameters are identical.
-Results Summary.
-Figure 11: Energy savings breakdown of large models emulation with straggler slowdown ($T^{\prime}/T$) 1.20 and 1024 GPUs.
-Figure 11 shows the amount of energy bloat reduction for GPT-3 175B and Bloom 176B large models when slowdown degree is 1.2 on emulated 1024 GPUs as a representative.
-Intrinsic Bloat Reduction.
-Model GPU Type
-Energy Savings (%)
-given # Microbatches
-12 24 48 96
-GPT-3 175B A100 9.02 9.77 10.17 10.39
-A40 11.81 10.22 9.34 8.88
-Bloom 176B A100 6.39 4.99 4.21 3.80
-A40 6.97 4.49 3.12 2.41
-Table 11: Perseus’s intrinsic energy bloat reduction of each non-straggler pipeline for GPT-3 175B and Bloom 176B with various number of microbatches. All scenarios run with 8 pipeline stages.
-Table 11 presents the changes of intrinsic energy bloat saving of a single pipeline in the case of GPT-3 175B and Bloom 176B with respect to various number of microbatches.
-GPT-3 175B on A100 is unique and has the opposite trend compared to others: intrinsic energy bloat reduction increases with the number of microbatches.
-The opposite trends stem from the amount of stage imbalance on each GPU, and the amount of slowdown the minimum-energy frequency incurs.
-What is unique about GPT-3 175B on A100 is that the amount of imbalance is large whereas the minimum-energy frequency incurs little slowdown for forward and backward computations.
-Therefore, only for GPT-3 175B on A100, microbatches in the middle of the pipeline are able to fully slow down to their potential.
-Thus, when such microbatches are added to the pipeline, the percentage of intrinsic energy bloat reduction increases.
-Extrinsic Energy Bloat.
-Figure 12: Perseus’s extrinsic energy savings during a single iteration by slowing down non-straggler pipelines.
-When $T^{\prime}>T^{*}$, slowing non-stragglers down starts increasing energy consumption, hence non-stragglers do not slow down further than $T^{*}$.
-More blocking time after $T^{*}$ decreases overall energy savings.
-Please note the different Y-axes.
-Figure 12 presents extrinsic energy bloat results in the strong scaling configuration, where a straggler is slower than others in varying degrees of slowdown from 1.05 to 1.50.
-Appendix H Iteration Time–Energy Frontiers
+{listing}{minted}
+[]python # In the framework’s pipeline execution engine: from perseus.client import profiler, controller
+def train_step(model, dataloader): …
+for instuction in pipeline_schedule: if isinstance(instruction, Forward): controller.set_speed(”forward”) profiler.begin(”forward”) # Run forward on microbatch profiler.end(”forward”) elif isinstance(instruction, Backward): controller.set_speed(”backward”) profiler.begin(”backward”) # Run backward on microbatch profiler.end(”backward”) … Perseus client API integration example.
+Appendix F Perseus Client Integration Example
+Listing E shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
+A typical structure of a pipeline execution engine is to have instructions for each distinct operations in the pipeline, including not only forward and backward executions, but also P2P and collective communications, and implement a handler for each instruction.
+Therefore, framework developers can wrap such handlers with the Perseus client APIs to mark their beginning and end.
+Appendix G Iteration Time–Energy Frontiers
 (a) BERT
@@
 (d) Wide-ResNet
-Figure 13: Eight stage pipeline parallelism on A40.
+Figure 11. Eight stage pipeline parallelism on A40.
 Figure 11 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
@@
 (d) Wide-ResNet
-Figure 14: 4 stage pipeline parallelism on A100 PCIe.
+Figure 12. 4 stage pipeline parallelism on A100 PCIe.
 Figure 12 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs.
@@
 Thus, the noise in data loading time becomes more noticeable in A100 PCIe.
-; bottom:-0.4ex;">Exml[LOGO] body>

### v3

@@
 Training large AI models on numerous GPUs consumes a massive amount of energy, making power delivery one of the largest limiting factors in building and operating datacenters for AI workloads.
-However, we observe that not all energy consumed during training directly contributes to end-to-end throughput, and a significant portion can be removed without slowing down training, which we call energy bloat.
+However, we observe that not all energy consumed during training directly contributes to end-to-end throughput; a significant portion can be removed without slowing down training.
+We call this portion energy bloat.
 In this work, we identify two independent sources of energy bloat in large model training and propose Perseus, a training system that mitigates both.
-To do this, Perseus obtains the “iteration time–energy” Pareto frontier of any large model training job using an efficient graph cut-based algorithm and schedules the energy consumption of computations across time to remove both types of energy bloat.
-Evaluation on large models including GPT-3 and Bloom shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification, enabling energy reduction – and therefore cost savings – otherwise unattainable before.
+To do this, Perseus obtains the time–energy tradeoff frontier of a large model training job using an efficient graph cut-based algorithm, and schedules computation energy consumption across time to reduce both types of energy bloat.
+Evaluation on large models, including GPT-3 and Bloom, shows that Perseus reduces the energy consumption of large model training by up to 30% without any throughput loss or hardware modification.
+[^1] Perseus is open-source as part of Zeus (zeus-nsdi23, ) at https://ml.energy/zeus.
+Energy-efficiency, datacenter power management, straggler, distributed training, large model training
+journalyear: 2024copyright: acmlicensedconference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695970isbn: 979-8-4007-1251-7/24/11ccs: Computer systems organizationccs: Computing methodologies Machine learningccs: Software and its engineering Power management
 1. Introduction
@@
 For instance, training GPT-3 (gpt3, ) reportedly consumed 1.3 GWh (patterson2021carbon, ).
-Then, this was recently dwarfed by a 200B model trained by Amazon consuming about 11.9 GWh (hamilton2024constraint, ), which is enough to power more than 1,000 average US households for a year (us-household, ).
-Such energy-intensive large model training workloads not only inflate datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ).
-Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ; oobleck-sosp23, ), energy optimization still remains an open challenge (greenai-cacm20, ; patterson2021carbon, ).
-While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute.
-As such, a recent line of work demonstrated that software plays a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; enemy-hpca23, ; envpipe-atc23, ).
-In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing the average power draw.
-To that end, we identify energy bloat – the portion of energy consumption that can be removed without slowdown – in software systems for large model training.
-We find two independent sources of energy bloat: intrinsic and extrinsic, and propose a single optimization framework that minimizes both.
+Then, this was dwarfed by Amazon’s training of a 200B model, which consumed about 11.9 GWh (hamilton2024constraint, )—enough to power more than 1,000 average US households for a year (us-household, ).
+Such energy-intensive large model training not only inflates datacenter operational expenses, but also made power delivery a primary challenge in building datacenters today (cbre2023, ; mckinsey2023, ; cbre2024, ; aienergy-joule24, ; openai-keynote-hotchips24, ).
+Despite recent works on accelerating large model training (megatronlm-sc21, ; alpa-osdi22, ; merak-tpds23, ), energy optimization remains an open challenge (greenai-cacm20, ; patterson2021carbon, ).
+While energy optimization is well-studied in the hardware community (eyeriss-jssc16, ; vivienne17efficient, ; puma-asplos19, ; kws-jssc24, ), the power bottleneck of recent datacenters (cbre2023, ; mckinsey2023, ; cbre2024, ; dvfs-boosting-asplos24, ; openai-keynote-hotchips24, ) shows that efficiency gains from hardware advancement alone are not sufficient to sustain the growing demand for AI compute.
+In light of this, recent works show that software can play a significant role in energy optimization by capturing application characteristics that general-purpose hardware cannot (e.g., no need to finish computation before the deadline), bringing hardware-agnostic energy-efficiency gains (gpoeo-tpds21, ; zeus-nsdi23, ; chase-ccai23, ; envpipe-atc23, ; crosslayer-energy-eecs24, ).
+In this paper, we seek a software method that reduces the energy consumption of large model training without slowdown, thereby also reducing average power draw.
+To that end, we identify energy bloat, the portion of energy consumption that can be removed without slowdown in software systems for large model training.
+We find two independent sources of energy bloat—intrinsic and extrinsic—and propose a single optimization framework that minimizes both.
 Intrinsic energy bloat comes from computation imbalance when a large model is distributed across multiple GPUs with pipeline parallelism (§2.2).
 Balancing the amount of computation in each pipeline stage is an important problem for distributed execution planning (pipedream-sosp19, ; dapple-ppopp21, ; autopipe-icml21, ; alpa-osdi22, ), but perfectly balancing every stage is not always possible because layers in a DNN are coarse-grained tensor operations with varying amounts of computation.
-When all stages do not have the same computation time, those not on the critical path of computation run needlessly fast – i.e., they consume energy that does not contribute to the overall training throughput.
+When stages have unequal computation times, those not on the critical path of computation run needlessly fast—that is, they consume energy that does not contribute to the overall training throughput.
 Such intrinsic energy bloat opens up the opportunity to precisely slow down each non-critical computation in the pipeline such that the length of the critical path does not change.
-Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion to scale out training to massive datasets, and one or more pipelines run slower than the others (§2.3).
-Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ), hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ), etc., and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ).
-All pipelines running faster than the straggler pipeline are needlessly fast and wasting energy that does not affect the overall training throughput.
-Thus, extrinsic energy bloat allows us to slow down entire pipelines (while still keeping their intrinsic energy bloat low) without delaying gradient synchronization.
+Extrinsic energy bloat, in contrast, arises when multiple pipelines run in parallel in a synchronous fashion, and one or more pipelines run slower than the rest (§2.3).
+Root causes behind such slowdowns are varied, including power/thermal throttling (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-asplos24, ), I/O bottlenecks in the storage/network (datastalls-vldb21, ; recsys-dsi-isca22, ; antdt-icde24, ), and hardware/software failures (bamboo-nsdi23, ; oobleck-sosp23, ; recycle-sosp24, ), and the likelihood of their presence increases with the scale and duration of training (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ).
+All pipelines running faster than the straggler pipeline are needlessly fast, wasting energy that does not affect the overall training throughput.
+Thus, we can slow down entire pipelines without delaying gradient synchronization.
 In this work, we propose Perseus, which formulates a unified optimization framework to remove both intrinsic and extrinsic energy bloat from large model training (§3).
-At its core, Perseus efficiently pre-characterizes the entire “iteration time–energy” Pareto frontier, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers.
+At its core, Perseus efficiently pre-characterizes the entire time–energy tradeoff frontier of a training iteration, allowing it to minimize intrinsic bloat under normal operation and to mitigate extrinsic bloat arising from stragglers.
 Existing works fall short on both fronts.
 EnvPipe (envpipe-atc23, ) is limited to intrinsic bloat reduction with a point solution that leads to suboptimal energy reduction.
-Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its “iteration time–energy” frontier suboptimal for large models.
-We show that characterizing the “iteration time–energy” Pareto frontier is NP-hard not only to solve but also to approximate within any constant factor.
+Zeus (zeus-nsdi23, ), in contrast, ignores intrinsic bloat as it only considers single-GPU training, which also renders its time–energy frontier suboptimal for large models.
+We show that characterizing the optimal time–energy Pareto frontier is NP-hard not only to solve, but also to approximate within any constant factor.
 Given this impasse, we propose an efficient algorithm that optimally solves a relaxed problem instead (§4).
-At its core, Perseus represents one training iteration as a directed acyclic graph (DAG).
-Each node in the DAG represents either a forward or backward computation in a particular stage, and nodes are connected with edges that represent computation dependencies.
-Perseus’s optimization algorithm then efficiently generates all energy schedules, defined as the planned time and energy consumption of each DAG node (computation), that are on the “iteration time–energy” Pareto frontier.
-In doing so, instead of finding each point on the frontier individually, Perseus employs a graph cut-based algorithm that iteratively crawls up the frontier from the bottom, generating the next Pareto-optimal energy schedule by carefully tweaking its neighbor energy schedule on the frontier.
-Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the Pareto frontier.
-Perseus is split into a server and a client library (§5).
-The client library can be integrated into any large model training framework as well as any accelerator that has an energy measurement and execution speed control interface, giving the server an abstract view of training computations and their time and energy consumption.
-The server produces Pareto-optimal energy schedules, which is realized by the client.
-Evaluation on large models including GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and scaled-up Wide-ResNet (wideresnet-bmvc16, ) shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing both energy consumption and average power draw (§6).
+To do so, Perseus represents one training iteration as a directed acyclic graph (DAG) of forward and backward computations in each pipeline stage.
+Then, Perseus efficiently generates all energy schedules, defined as the planned time and energy consumption of each computation, that are on the time–energy frontier using a graph cut-based algorithm that iteratively crawls up the frontier from the bottom.
+Minimizing intrinsic and/or extrinsic energy bloat is then as simple as choosing the appropriate energy schedule from the pre-characterized time–energy frontier.
+Perseus consists of a client library and a server (§5).
+The client library integrates with a large model training framework and accelerator to measure computation energy consumption and control accelerator speed.
+The server produces optimized energy schedules, using the abstract computation DAG and time/energy measurements provided by the client.
+Evaluation on large models (GPT-3 (gpt3, ), BERT (bert, ), T5 (t5-jmlr20, ), Bloom (bloom-arxiv22, ), and a scaled-up version of Wide-ResNet (wideresnet-bmvc16, )), shows that Perseus is able to reduce per-iteration energy consumption by up to 30% with negligible or no slowdown, reducing energy consumption and average power draw (§6).
 Overall, we make the following contributions in this paper:
   • We identify intrinsic and extrinsic energy bloat in large model training, fundamentally caused by computation time imbalance at different levels.
-  • We propose Perseus, a software-only energy optimization system that removes energy bloat through a unified optimization framework and a graph cut-based algorithm.
+  • We propose Perseus, a software-only energy optimization system that reduces energy bloat through a unified optimization framework and a graph cut-based algorithm.
   • We evaluate Perseus on a diverse set of large model workloads and show that it significantly reduces energy bloat, bringing hardware-agnostic energy savings.
 2. Motivation
+Figure 1.
+One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs, drawn to scale.
+For example, F5 and B5 in the S2 row denote forward and backward for the fifth microbatch on Stage 2.
+The critical path is traced with a blue line.
+Colors show power consumption.
+Other models are visualized in Appendix A.
 First, we provide necessary background regarding large model training (§2.1).
@@
 2.1. Large Model Training
-Large model training today is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ; llama3-blog, ; deepspeed-github, ; megatronlm-github, ; gpt-neox-github, ).
-Especially, pipeline parallelism partitions a large model and its training batch into multiple stages and microbatches respectively, and pipelines forward and backward computations through stages.
-Then, such pipelines are replicated multiple times to perform data parallel training.
-Since all pipelines must synchronize gradients at the end of each iteration, computations for the next iteration can only begin when all pipelines have finished and exchanged gradients.
+Large model training is mostly dominated by 3D (data, tensor, and pipeline) parallelism (megatronlm-sc21, ; megatron-turing-arxiv22, ; bloom-arxiv22, ; merak-tpds23, ; falcon-arxiv23, ).
+Especially, pipeline parallelism partitions a large model into multiple stages and its training batch into microbatches, and pipelines forward and backward computations through the stages.
+Then, such pipelines are replicated to perform data parallel training.
+Pipelines can only move on to the next iteration after every pipeline has finished and synchronized gradients.
 2.2. Intrinsic Energy Bloat
-Figure 1.
-One training iteration of GPT-3 1.3B with 4 pipeline stages and 6 microbatches on NVIDIA A100 GPUs drawn to scale.
-For example, F5 and B5 in the S2 row denote Forward and Backward for the fifth microbatch on Stage 2.
-The critical path of computation is traced with a blue line.
-Colors show power consumption.
-Especially, the blue background indicates that GPUs consume power when blocking on communication.
-Other models are visualized in Appendix A.
+We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a).
+In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage.
+Such gaps exist because the computation time of each pipeline stage is not perfectly balanced.
+Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying sizes.
 Model # Parameters Imbalance Ratio
@@
 Wide-ResNet101 (wideresnet-bmvc16, ) 1.5B 1.09 1.25
-Table 1. Latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
-We profile GPT-3 1.3B on NVIDIA A100 GPUs and visualize the timeline of one training iteration in Figure 1(a).
-In addition to the familiar bubbles in the 1F1B schedule (megatronlm-sc21, ), we observe gaps in between forward and backward computations, where the GPU is simply blocking on communication with an adjacent stage.
-Such gaps exist because the computation time of each pipeline stage is not perfectly balanced.
-Partitioning stages in a balanced manner is an important problem in distributed execution planning (gpipe-neurips19, ; pipedream-sosp19, ; dapple-ppopp21, ; alpa-osdi22, ), but perfect balancing is difficult because DNNs are essentially a sequence of coarse-grained tensor operations with varying size.
+Table 1. Forward latency ratio of the longest to the shortest stage on A100 GPUs. 1.00 would mean perfect balance.
 To understand the amount of possible pipeline stage imbalance, we exhaustively searched for the pipeline partition with the smallest imbalance ratio, defined as the ratio of the longest stage forward computation latency to the shortest.
@@
 See Appendix B for partitioning details and sources of imbalance.
-Given stage imbalance, as we have seen in Figure 1(a), not all forward and backward computations are on the critical path of computation.
-This means that non-critical computations running at their maximum speed is not contributing to faster iteration time, and thus simply wasting energy.
+Given stage imbalance, not all forward and backward computations are on the critical path of computation (Figure 1(a)).
+This means that non-critical computations running at their maximum speed are not contributing to faster iteration time, and thus simply wasting energy.
 We call this intrinsic energy bloat, which can be reduced by precisely slowing down each non-critical computation without lengthening the critical path (Figure 1(b)).
@@
 Numerous replicas of the same pipeline run in a data parallel fashion in large model training.
-Because every pipeline must synchronize gradients at the end, if a pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)).
-Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time is wasteful.
-We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the pipeline.
-To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time and precisely slow down computations in the non-straggler keeping that iteration time (Figure 2(b)).
+Because every pipeline must synchronize gradients at the end, if one pipeline runs slower, all other pipelines must wait until the straggler pipeline finishes (Figure 2(a)).
+Since the straggler pipeline determines end-to-end iteration time, all other pipelines running at their fastest possible iteration time are wasteful.
+We call this extrinsic energy bloat, because unlike intrinsic energy bloat, its cause is extrinsic to the training pipeline.
+To reduce extrinsic bloat while keeping intrinsic bloat low, one can determine the energy-optimal iteration time for non-straggler pipelines and precisely slow down computations so that non-straggler pipelines attain that iteration time (Figure 2(b)).
 Figure 2.
-Two data parallel pipelines run, and the first one became a straggler (e.g., due to the second GPU temporarily overheating).
-The non-straggler pipeline wastes energy by running as fast as possible.
-Such extrinsic energy bloat can be removed by precisely slowing down non-straggler pipelines.
+Among two data parallel pipelines, the first one becomes a straggler.
+The non-straggler pipeline causes extrinsic energy bloat by running as fast as possible (a), which can be reduced by precisely slowing it down (b).
 Stragglers arise from numerous sources.
-Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-arxiv23, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ), acting like a persistent straggler pipeline.
-Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ) deploy heterogeneous pipelines, introducing non-uniform iteration times.
+Thermal or power throttling in a datacenter can result in 10–50% slowdown (A100powerthermal-energies21, ; gpupower-cal23, ; thunderbolt-osdi20, ; mvpp-asplos20, ; polca-asplos24, ), and I/O bottlenecks in the storage or network can be longer than GPU computation by up to 4×4$\times$4 × (datastalls-vldb21, ; recsys-dsi-isca22, ; antdt-icde24, ), acting like a persistent straggler pipeline.
+Recent failure-resilient training frameworks (bamboo-nsdi23, ; oobleck-sosp23, ; recycle-sosp24, ) deploy heterogeneous pipelines, introducing non-uniform iteration times.
 With increasing job and infrastructure scale, the probability of encountering stragglers increases (largescalefailures-sc17, ; philly-atc19, ; mlaas-nsdi22, ; megascale-nsdi24, ; tpuv4-nsdi24, ).
-Stragglers are often times known to and anticipated by the training infrastructure, because they were created by the infrastructure itself (e.g., throttling, non-compute bottlenecks, fault-tolerant planning).
-Therefore, in this work, we focus on how to plan time and energy across time and allow quick adaptation assuming that such information is available.
+In this work, we focus on stragglers that are known to and anticipated by the training infrastructure, generally because they were created by the infrastructure itself (e.g., power and thermal throttling, non-compute bottlenecks, fault-tolerant planning).
+Such stragglers also tend to persist beyond typical training iteration times.
+Therefore, Perseus focuses on planning time and energy consumption across time and allowing quick adaptation, assuming that information about stragglers is available.
 2.4. Potential Benefits of Reducing Energy Bloat
-To gauge potential energy savings, we measure the energy savings of slowing down every computation in the pipeline to their minimum-energy frequencies.
-This will slow down the pipeline’s iteration time, but can act as an upper bound for energy savings.
-For our experimental workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively.
-We show in Section 6.2 that Perseus can realize most of the potential savings with negligible slowdown, even through our problem is NP-hard.
+To gauge potential energy savings, we measure the energy savings achieved by slowing down every computation in the pipeline to their minimum-energy frequencies.
+This will slow down iteration time, but can act as an upper bound for energy savings.
+For our workloads in Section 6.2, this gives on average 16% and 27% energy reduction on A100 and A40 GPUs, respectively.
+Section 6.2 shows that Perseus can realize most of the potential savings with negligible slowdown.
 Figure 3.
-Three cases of where the straggler pipeline’s iteration time $T^{\prime}$ can be.
-$T_{\min}$ and $T^{*}$ are the shortest and longest iteration time on the “iteration time–energy” Pareto frontier.
+Three cases showing where the straggler pipeline’s iteration time $T^{\prime}$ can be.
+$T_{\min}$ and $T^{*}$ are the shortest and longest iteration times on the time–energy frontier.
 The black dot is when all computations run at the maximum speed, which wastes energy.
 The green dot is the energy-optimal iteration time of the non-straggler pipeline.
-Solid orange dots make up the “iteration time–energy” Pareto frontier, and the orange dotted line shows that energy increases beyond $T^{*}$.
+Solid orange dots make up the frontier, and the orange dotted line shows that iteration energy increases beyond $T^{*}$.
 3. Perseus Overview
-Perseus is an energy optimization system for large model training.
-In this section, we present our unified optimization framework that aims to remove both types of energy bloat (§3.1) and walk through the workflow of Perseus (§3.2).
+We first present Perseus’s unified optimization framework that aims to remove both types of energy bloat (§3.1), and then walk through the workflow of Perseus (§3.2).
 3.1. Unified Optimization Framework
@@
 Our goal is to minimize a pipeline’s energy consumption by controlling the execution speed of each computation in the pipeline.
-In doing so, we can slow down a pipeline’s iteration time at most until the straggler’s iteration time $T^{\prime}$:
+In doing so, we can slow down a pipeline’s iteration time up to the straggler’s iteration time $T^{\prime}$:
 (1) $\min_{F}$ $\quad\textrm{Energy}(F)$
 s.t. $\quad\textrm{Time}(F)\leq T^{\prime}$
-where $F$ is the set of GPU frequencies [^2] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms (AMD GPUs via AMDSMI (amdsmi, )), which is much shorter than typical large model computation latencies.
+where $F$ is an assignment of GPU frequencies [^3] The SM frequency of NVIDIA GPUs can be set via NVML (nvml, ) in around 10 ms, which is much shorter than typical large model computation latencies.
 Locking the GPU’s frequency provides deterministic computation latency (nvml, ; clockwork-osdi20, ), making it suitable for tightly planning and packing execution over time.
-to run each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively.
-Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ on the “iteration time–energy” 2D plane.
-However, we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the “iteration time–energy” Pareto frontier.
-Now, let us assume we have a fully characterized “iteration time–energy” Pareto frontier with $T_{\min}$ and $T^{*}$ bookending the frontier (described in Section 4).
-$T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption.
-Figure 3 depicts the three cases where the straggler’s iteration time $T^{\prime}$ can possibly be w.r.t.
-the Pareto frontier:
+to each forward and backward computation in the pipeline, and $\textrm{Time}(F)$ and $\textrm{Energy}(F)$ are the iteration time and energy consumption of the pipeline when executed with $F$, respectively.
+Changing $F$ will lead to different values of $\textrm{Time}(F)$ and $\textrm{Energy}(F)$, but we are only interested in ($\textrm{Time}(F)$, $\textrm{Energy}(F)$) points that are on the time–energy tradeoff frontier.
+Now, let us assume we have a fully characterized time–energy frontier, bookended by $T_{\min}$ and $T^{*}$ (Section 4 is dedicated to describing how).
+$T_{\min}$ is the shortest iteration time on the frontier, which is the same as the iteration time of running every computation at the maximum speed, and $T^{*}$ is the iteration time with minimum energy consumption, which is when each computation runs at the frequency that consumes the least amount of energy for that computation.
+[^4] This is typically not the lowest frequency, because computations running with very low frequencies incur more latency increase than power reduction, resulting in higher energy consumption.
+Figure 3 shows the three possible cases regarding where the straggler’s iteration time $T^{\prime}$ can be:
   1. (1)
-    When there are no stragglers (Figure 3(a)), we simply operate on the point on the Pareto frontier with iteration time $T_{\min}$, which reduces intrinsic energy bloat.
+    Figure 3(a): When there are no stragglers, we simply select the point on the frontier with iteration time $T_{\min}$, which reduces only intrinsic energy bloat.
   2. (2)
-    When a moderately slow straggler is detected (Figure 3(b)), we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
+    Figure 3(b): When a moderately slow straggler is detected, we additionally reduce extrinsic energy bloat while keeping intrinsic bloat low by slowing down all non-straggler pipelines until $T^{\prime}$, using up all the slack time.
   3. (3)
-    Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier (Figure 3(c)).
+    Figure 3(c): Finally, the straggler’s iteration time may go beyond the minimum-energy point $T^{*}$ on the frontier.
 In this case, we only slow down non-stragglers until $T^{*}$, because going past $T^{*}$ will instead increase energy.
@@
 (2) $T_{\textrm{opt}}=\min(T^{*},T^{\prime}).$
-Therefore, if we pre-characterize the entire “iteration time–energy” Pareto frontier by finding all Pareto-optimal $F$s, Perseus will be able to react instantly to emerging stragglers by looking up $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$.
-Finally, we note that unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal.
-Being too fast or too slow can both waste energy.
+Therefore, when a straggler emerges (i.e., $T_{\min}<T^{\prime}$), Perseus can compute $T_{\textrm{opt}}$ using Equation 2 and quickly look up the frequency plan $F_{\textrm{opt}}$ that leads to iteration time $T_{\textrm{opt}}$ using the pre-characterized time–energy frontier.
+Finally, we note that, unlike other problem settings that do not consider energy consumption, fully utilizing all the slack time created by the straggler is not always energy-optimal; being too fast or too slow can both waste energy.
 3.2. Perseus Architecture
@@
 System Components.
-Figure 4 shows the high-level architecture of Perseus.
-Perseus is split into a framework- and hardware-agnostic server and a framework-integrated and device-specific client.
-The server pre-characterizes the “iteration time–energy” Pareto frontier (§4) and caches them for fast lookup.
-The client profiles pipeline computations online during training and realizes the energy schedule by changing the GPU’s frequency during runtime (§5).
+Perseus’s architecture is shown in Figure 4.
+Perseus consists of a framework- and accelerator-agnostic server and a framework-integrated and accelerator-specific client.
+The server is a cluster-wide singleton.
+For various training jobs, the server pre-characterizes the time–energy frontier of one iteration (§4) and caches energy schedules for fast lookup.
+The client profiles pipeline computations online during training and realizes energy schedules by setting the GPU’s frequency during runtime (§5).
 Figure 4. Perseus architecture and workflow.
-Training Lifecycle Using Perseus.
-In Perseus, a training job is primarily specified by its computation DAG for one training iteration.
-When the job begins execution, 1 the Perseus client invokes its Online Time/Energy Profiler (§5) to measure the time and energy of each forward and backward computation on each frequency while training is running.
-Upon receiving profile results, 2 the server characterizes the “iteration time–energy” Pareto frontier (§4), and Pareto-optimal energy schedules are saved in a lookup table indexed by $T^{\prime}$.
-Then, 3 the energy schedule corresponding to the shortest iteration time is deployed to the client.
+Training Lifecycle.
+For the Perseus server, a training job is primarily specified by its computation DAG for one training iteration.
+When the job begins execution, 1 the Perseus client invokes its online time–energy profiler (§5) to measure the time and energy of each forward and backward computation on each supported frequency.
+Profiling is done in vivo during the initial tens of training iterations.
+Upon receiving profiling results, 2 the server begins asynchronously characterizing the time–energy frontier (§4) while training continues.
+When characterization finishes, energy schedules on the frontier are saved in a lookup table indexed by $T^{\prime}$.
+Then, 3 the energy schedule corresponding to $T_{\min}$ is deployed to the client.
 Energy schedules are realized by the client’s asynchronous frequency controller, integrated into the training framework (§5).
 During training, 4 the training infrastructure (e.g., datacenter rack power/temperature manager) notifies the Perseus server of a straggler and its degree of slowdown.
-The server then 5 quickly reacts to this by looking up the Pareto-optimal energy schedule corresponding to the anticipated straggler iteration time, and deploys it to the client.
-4. Characterizing the Pareto Frontier
-In this section, we describe our algorithm to efficiently obtain the “iteration time–energy” Pareto frontier for a training pipeline in detail.
-We first formulate the problem, show that it is NP-hard, and provide a relaxation (§4.1).
+The server then 5 quickly reacts to this by looking up the energy schedule corresponding to the anticipated straggler iteration time (the one with iteration time $T_{\textrm{opt}}$), and deploys it to the client.
+4. Characterizing the Time–Energy Frontier
+In this section, we describe our algorithm to efficiently obtain the time–energy tradeoff frontier for a training pipeline in detail.
+We first formulate the problem, show that it is NP-hard, and describe a relaxed version (§4.1).
 Then, we provide an overview of our algorithm (§4.2) and describe the core subroutine in our algorithm (§4.3).
@@
 where $P_{\textrm{blocking}}$ is the power consumption of the GPU when it is blocking on communication, $N$ is the number of pipeline stages, and $t_{i}(f_{i})$ and $e_{i}(f_{i})$ are the time and energy consumption of computation $i$ with frequency $f_{i}$, respectively.
-As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the “iteration time–energy” Pareto frontier ($\textrm{Time}(F)$ vs.
+[^5] An assumption here is that $P_{\textrm{blocking}}$ is constant, as a GPU blocking on communication is busy-looping inside a NCCL kernel without heavy computation utilization.
+As derived in Section 3.1, given straggler iteration time $T^{\prime}$, we draw a vertical line on the time–energy tradeoff frontier ($\textrm{Time}(F)$ vs.
 1+ 2) at $T_{\textrm{opt}}$ and find $F_{\textrm{opt}}$ where the two lines intersect.
-Equation 3 shows that the “iteration time–energy” Pareto frontier of a pipeline only depends on the straggler’s iteration time $T^{\prime}$ in 2, which is only a constant upwards shift of the frontier.
-Therefore, if we characterize the Pareto frontier of $\textrm{Time}(F)$ vs.
-1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any value of $T^{\prime}$.
+Equation 3 shows that the time–energy frontier of a pipeline depends on the straggler’s iteration time $T^{\prime}$ only in the second term 2, which is merely an upward shift of the frontier.
+Therefore, if we characterize the tradeoff frontier of $\textrm{Time}(F)$ vs.
+1, that frontier can be used to find $T_{\textrm{opt}}$ and $F_{\textrm{opt}}$ for any straggler iteration time $T^{\prime}$.
 Thus, we define
 (4) $\textrm{Energy}(F)=\sum_{i}{\left(e_{i}(f_{i})-P_{\textrm{blocking}}\cdot t_{i}(f_{i})\right)}$
-and characterize the Pareto frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
-Finding the Pareto Frontier.
-Finding one point on the Pareto frontier with iteration time $T$ is equivalent to solving the following optimization problem:
+and characterize the frontier of $\textrm{Time}(F)$ vs. $\textrm{Energy}(F)$.
+Finding the Time–Energy Frontier.
+Finding one point on the Pareto-optimal tradeoff frontier with iteration time $T$ is equivalent to solving the following optimization problem:
 (5) $\min_{F}$ $\quad\textrm{Energy}(F)$
@@
 Reduction from Knapsack. Details in Appendix C. ∎
-The complete Pareto frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable. Therefore, we seek an appropriate relaxation of the problem.
-One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is because the possible choices of GPU frequencies are discrete.
-However, if choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ).
+The complete Pareto-optimal tradeoff frontier can be obtained by solving PEM for all $T\in[T_{\min},T^{*}]$, which is clearly intractable.
+Therefore, we seek an appropriate relaxation of the problem that will yield a nearly Pareto-optimal frontier.
+One of the reasons PEM is NP-hard is because it is a discrete optimization problem where the possible choices of computation time and energy are discrete, which is in turn because GPUs only support discrete frequencies (e.g., in 15 MHz steps and nothing in the middle).
+However, if frequency choices were continuous, the problem is exactly and efficiently solvable (skutella1998, ).
 This is akin to integer linear programs becoming tractable when relaxed to linear programs.
 The transform from the original problem to the relaxed version is done by fitting a continuous exponential function ($a\cdot e^{bt}+c$) to Pareto-optimal computation time and energy measurements for each forward and backward computation.
-We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realizes most of the opportunity for savings.
+We choose the exponential function due to its inherent flexibility and natural fit to data (more details in Appendix D).
+We show in Section 6.2 that this relaxation produces high-quality approximate solutions that realize most of the opportunity for savings.
 Solving the relaxed problem returns the time and energy consumption planned for each computation in the pipeline, or the energy schedule.
-Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$ (Section 4.3).
-4.2. Iteratively Discovering the Pareto Frontier
-Figure 5. Starting from the Pareto-optimal energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up the Pareto frontier.
+Then, this is transformed back to a feasible solution of the original problem, which is the set of GPU frequencies $F$.
+4.2. Iteratively Discovering the Frontier
+Figure 5. Starting from the energy schedule that consumes the minimum energy, we iteratively reduce its iteration time to trace up and iteratively discover the tradeoff frontier.
+Now, we first describe our iterative strategy of characterizing the frontier, and dive deeper into one iteration in Section 4.3.
 Although our relaxed problem is no longer NP-hard, solving it for each $T^{\prime}\in[T_{\min},T^{*}]$ from scratch is inefficient.
-Instead, what if we can tweak an energy schedule that is already Pareto-optimal to generate its neighbor energy schedule on the Pareto frontier?
-Then, we can start from one end of the Pareto frontier and trace along the frontier to the other end, discovering every Pareto-optimal energy schedule.
+Instead, what if we can tweak an existing schedule already on the frontier to generate its neighbor energy schedule on the frontier?
+Then, we can start from one end of the frontier and trace along to the other end, discovering fine-grained optimized energy schedules.
 […]
-We note that starting from the energy schedule that consumes the maximum energy (i.e., the black dot) would be incorrect.
-That energy schedule is not Pareto-optimal because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
+We note that tracing down from the energy schedule that consumes the maximum energy (i.e., Figure 5 black dot) would be incorrect.
+That schedule is far from optimized because, although it will execute with the least amount of time, stage imbalance leaves room for energy reduction (§2.2).
 1
@@
 Iteration time with all max frequencies $T_{\min}$
-Output: Set of Pareto-optimal energy schedules $\mathcal{P}$
+Output: Set of optimized schedules $\mathcal{S}$
 4
@@
 $\triangleright$ Begin with the minimum energy schedule
-7 $p\leftarrow$ Minimum energy for all computations
-8 $\mathcal{P}\leftarrow\{p\}$
+7 $s\leftarrow$ Minimum energy for all computations
+8 $\mathcal{S}\leftarrow\{s\}$
 9
-10while $\mathrm{IterationTime}(\mathcal{G},p)>T_{\min}$ do
+10while $\mathrm{IterationTime}(\mathcal{G},s)>T_{\min}$ do
        $\triangleright$ Reduce time by $\tau$ with minimal energy increase (§4.3)
-11       $p\leftarrow$ GetNextPareto($\mathcal{G}$, $p$, $\tau$)
-12       $\mathcal{P}\leftarrow\mathcal{P}\cup\{p\}$
+11       $s\leftarrow$ GetNextSchedule($\mathcal{G}$, $s$, $\tau$)
+12       $\mathcal{S}\leftarrow\mathcal{S}\cup\{s\}$
 13
@@
 15
-$\triangleright$ Every schedule in $\mathcal{P}$ is Pareto-optimal
-16 return $\mathcal{P}$
+16return $\mathcal{S}$
 17
 18
-Algorithm 1 Iteratively Discovering the Pareto Frontier.
+Algorithm 1 Iteratively discovering the frontier.
 Algorithm 1 provides an overview of our optimization process.
@@
 Starting from there, the iteration time of the schedule is iteratively reduced by unit time $\tau$ while incurring minimal energy increase (line 1; Section 4.3).
-This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process is Pareto-optimal.
-4.3. Finding the Neighbor Pareto-Optimal Schedule
+This is repeated until the total iteration time of the schedule can no longer be reduced, and every energy schedule encountered in the process forms our frontier.
+4.3. Finding the Neighbor Energy Schedule
 Figure 6.
@@
 Given a 1F1B pipeline schedule with 2 stages and 3 microbatches ( 1), it is first transformed to an equivalent representation of computation DAG ( 2).
-Then the Critical DAG ( 3) is obtained by considering every and only the computations on the critical path.
+Then the Critical DAG ( 3) is obtained by considering only the computations on the critical path.
 Our key observation is that any valid s-t cut on the Critical DAG will reduce the iteration time by unit time $\tau$.
-Cut A and Cut B are two examples of valid s-t cut ( 4).
+Cut A and Cut B are two examples of valid s-t cuts ( 4).
 Either reducing the one computation associated with Cut A ( 5A ) or reducing the two computations associated with Cut B ( 5B ) reduces the iteration time by $\tau$.
-In this section, we describe our core subroutine GetNextPareto (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
+In this section, we describe our core subroutine GetNextSchedule (Algorithm 1, line 1). Figure 6 provides visualizations of the process. The entire procedure is given in Algorithm 2.
 Node- and Edge-Centric Computation DAGs.
@@
 Solving with Minimum Cut.
-We have shown that an s-t cut represents a way to speed up the entire DAG by $\tau$.
+We have seen that any s-t cut represents a way to speed up the entire DAG by $\tau$.
 But speeding up computations increases energy.
 Then, a natural question is, which cut brings the smallest energy increase?
-We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity.
-Then, our problem reduces to minimum cut, which we can solve with maximum flow.
-After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining a new Pareto-optimal energy schedule.
-Appendix D provides details with mathematical expressions for flow capacities.
+We can precisely map the flow capacity of an s-t cut to the amount of energy increase from speeding up cut edges.
+That is, by finding the amount of energy increase each computation will incur with the slope of its exponential function (§4.1) and defining it to be the edge’s flow capacity, we can reduce our problem to minimum cut, which we can solve with maximum flow.
+After finding the minimum cut, we modify the durations of the computations involved in the cut, obtaining the neighbor energy schedule.
+Appendix E provides details with mathematical expressions for flow capacities.
 Converting Back to GPU Frequencies.
@@
 On GPUs, the “speed” control knob is the SM frequency.
-The server API will issue an HTTP/RPC request to the server.
-The id parameter is the ID of an accelerator.
+set_straggler is invoked by the infrastructure with the id of an accelerator to notify the server via HTTP/RPC.
 Time Complexity Analysis.
@@
 While for general DAGs the total number of steps is known to be exponential to the size of the DAG (skutella-thesis, ), we prove that for DAGs that represent pipeline computations, the number of steps is $O(N+M)$, yielding a final polynomial time complexity of $O((N+M)N^{3}M^{3})$.
-See Appendix E for proof.
+See Appendix F for proof.
 In reality, commonly used number of stages ($N$) is 4 to 8 (at most tens) to reduce pipeline bubble ratio (megatronlm-sc21, ; mlperf_training_31, ).
-Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far less have been reported even for high-performance settings (mlperf_training_31, ).
+Number of microbatches ($M$) is typically around $4N$ (gpipe-neurips19, ; megatron-turing-arxiv22, ), but recently with high data parallel degree, far fewer have been reported even for high-performance settings (mlperf_training_31, ).
 As such, algorithm runtime is practically negligible (§6.5), especially given that large model training easily takes weeks or months (patterson2021carbon, ).
@@
 Input: DAG $\mathcal{G}$ of computations $i\in\mathcal{G}$
-Current Pareto-optimal energy schedule $p$
+Current energy schedule $s$
 Amount of iteration time to reduce $\tau$
-Output: Pareto-optimal schedule w/ reduced time $p^{\prime}$
+Output: Neighbor schedule with reduced time $s^{\prime}$
 4
@@
 $\triangleright$ Find set of computations to modify ( 4)
-18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $p$)
+18 $S$, $T\leftarrow$ FindMinCut($\mathcal{G}$, $s$)
 19
@@
 $\triangleright$ Assign frequencies from planned computation times
-22 $p^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
+22 $s^{\prime}\leftarrow$ $\min f_{i}$ that runs no slower than planned
 23
-24return $p^{\prime}$
+24return $s^{\prime}$
 25
 26
-Algorithm 2 GetNextPareto: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
+Algorithm 2 GetNextSchedule: Reducing the execution time of the DAG by $\tau$ with minimal energy increase.
 4.4. Generalizations
@@
 Constant-Time Operations.
-Apart from computation and blocking on communication, there are other operations in the training pipeline that may take non-trivial latency.
+There are operations in the training pipeline that may take non-trivial latency, other than computation and blocking on communication.
 For instance, loading and copying input data into VRAM or communication over slower links can take considerable latency.
@@
 The Perseus server and client are implemented in Python.
-Perseus can optimize the energy consumption for any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented.
+Perseus can optimize any training infrastructure, framework, and accelerator as long as the APIs in Table 2 can be implemented, and the accelerator supports multiple execution speeds that trade off computation time and energy.
 As a reference, we have integrated the Perseus client with Merak (merak-tpds23, ), which marries high-performance tensor parallelism of Megatron-LM (megatronlm-github, ) and the generic pipeline execution engine of DeepSpeed (deepspeed-github, ).
-We provide an example of what is looks like to integrate the client with a training engine in Appendix F.
+While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs.
+We provide an example of what is looks like to integrate the client with a training engine in Appendix G.
 Activation recomputation (checkpointing-arxiv16, ) is enabled to allow large batch sizes to fit in GPUs.
-Online Time/Energy Profiler.
-The client profiles the time and energy of each forward and backward computation in vivo at the beginning of the training job to take measurements representative of real training.
-While training engine implementations differ widely, many have separate code blocks for forward and backward, allowing them to be wrapped with the profiler APIs for measurement.
-The GPU’s frequency is profiled from the highest to the lowest at iteration granularity and stopped when energy consumption increases.
-Beyond that point, frequencies will consume more time and energy, making them suboptimal.
+Profiler.
+Accurate profiling is important to our optimization algorithm; inaccurate latency profiles (especially underestimations) may slow down the end-to-end latency of the DAG, whereas inaccurate energy profiles can lead the algorithm to incorrectly select computations to speed up.
+Fortunately, the latency of a fixed set of GPU computations, especially with the GPU’s frequency locked, is known to be very stable (nvml, ; clockwork-osdi20, ).
+Furthermore, to ensure that profiling results are representative of real training, the Perseus client profiles the time and energy of each forward and backward computation at the beginning of the training job in vivo.
+Each supported GPU frequency is profiled one by one from the highest to the lowest for about five iterations (more if one iteration has less microbatches).
+After a certain frequency, lower frequencies result in both more time and energy consumed, making them strictly suboptimal compared to higher frequencies.
+Profiling is terminated at that point.
+Finally, we profile $P_{\textrm{blocking}}$ using two GPUs.
+One GPU blocks on P2P communication and the other sleeps, and we measure the power consumption of the blocking GPU.
+It is sufficient to profile $P_{\textrm{blocking}}$ once per GPU model.
 Asynchronous Frequency Controller.
@@
   • In emulated large-scale training scenarios, Perseus significantly outperforms the baselines by consistently providing up to 30% energy savings (§6.3).
-  • Energy bloat reduction was possible because Perseus can enumerate efficient energy schedules on the “iteration time–energy” Pareto frontier (§6.4).
+  • Energy bloat reduction is possible because Perseus can enumerate efficient energy schedules on the time–energy frontier (§6.4).
   • Perseus reduces energy bloat with low overhead (§6.5).
@@
 Metrics.
-We report energy reduction and slowdown of a training iteration (%), relative to using all maximum GPU frequencies.
-In most cases slowdown is close to zero, in which case energy and power reductions coincide.
+We report GPU energy reduction and slowdown of a training iteration (%) relative to using all maximum GPU frequencies.
+In most cases slowdown is close to zero, in which case energy and average power reductions coincide.
 Reducing only extrinsic bloat is not possible, because Perseus reduces extrinsic bloat while keeping intrinsic bloat low as it slows down non-straggler pipelines.
@@
   • EnvPipe (envpipe-atc23, ) reduces only intrinsic energy bloat while trying to minimize slowdown. We compare Perseus’s energy bloat reduction with EnvPipe (§6.2, §6.3).
-  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s “iteration time–energy” Pareto frontier against that of Zeus (§6.4).
+  • Zeus (zeus-nsdi23, ) characterizes the time–energy tradeoff of single GPU training. We compare Perseus’s time–energy frontier against that of Zeus (§6.4).
 6.2. Reducing Energy Bloat
-We start with overall energy bloat reduction – intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2) – achieved by Perseus and EnvPipe.
-All numbers were obtained by running on testbed GPUs, and all solutions use the same amount of GPU resources.
+We start with overall energy bloat reduction—intrinsic bloat without stragglers (§6.2.1) and intrinsic + extrinsic bloat with stragglers (§6.2.2)—achieved by Perseus and EnvPipe.
+All numbers were obtained by running on testbed GPUs.
+All solutions use the same amount of GPU hardware resources.
 6.2.1. Intrinsic Bloat Reduction Without Stragglers
@@
 Table 3. [Experiment] Intrinsic energy bloat (without stragglers) reduction and iteration time slowdown.
-Table 3 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the “iteration time–energy” frontier) and that by EnvPipe.
+Table 3 compares the energy savings achieved by Perseus’s minimum iteration time energy schedule (leftmost point of the time–energy frontier) and that by EnvPipe.
 We make two observations regarding Perseus.
-First, models show varying amounts of energy savings because (1) their stage imbalance vary (Table 1), and (2) their forward and backward are composed on different computations, which affects time/energy sensitivity when changing the frequency.
+First, models show varying amounts of energy savings because (1) their stage imbalances vary (Table 1), and (2) their forward and backward are composed of different computations, which affects time/energy sensitivity when changing the frequency.
+For instance, unlike other models, Wide-ResNet 1.5B on A100 after minimum imbalance stage partitioning has nearly perfect stage balance, resulting in low intrinsic energy bloat.
+However, as will be seen in Section 6.2.2, such models tend to achieve greater extrinsic bloat savings because most of their computations run at a high frequency, and slowing them down due to stragglers leads to higher energy reduction.
 Second, A40 demonstrates more energy savings compared to A100.
@@
 Thus, tuning down the GPU’s frequency yields a relatively smaller change in computation time and energy compared to those at the maximum frequency.
-However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
-EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest, which is not always true.
+However, we expect the more recent GPUs to have better percentage savings due to higher maximum frequency (e.g., 1980 MHz for H100 SXM (H100whitepaper, )) and better absolute savings due to high TDP (e.g., 1,200 W for each GPU on GB200 (blackwell-news, )).
+EnvPipe in general provides lower energy savings, primarily due to its assumption that the final stage of a pipeline is always the heaviest.
+This is only correct with a probability of $1/N$, where $N$ is the number of pipeline stages.
 Additionally, it sometimes considerably degrades iteration time because it is not aware of single-choice operations inside the pipeline (§4.4) and can slow down some computations too much.
@@
 Table 4. [Experiment] Energy savings given varying straggler slowdown ($T^{\prime}/T$). Perseus can reduce extrinsic bloat while keeping intrinsic bloat low, whereas EnvPipe cannot.
-When stragglers create extrinsic energy bloat, the amount of energy savings for a non-straggler pipeline depends on how much energy reduction its “iteration time–energy” Pareto frontier yields for longer iteration times.
-Table 4 shows the amount of energy savings of a pipeline given varying degrees of straggler slowdown.
-As described in Section 3.1, a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$, and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time blocking on communication, waiting for the straggler.
-Slowing down the non-stragglers beyond $T^{*}$ will increase total energy consumption; hence, Perseus does not slow down pipelines further than $T^{*}$.
-The percentage of savings first increases with the straggler’s iteration time, but then slowly wanes as it slows down beyond $T^{*}$.
+When stragglers create extrinsic energy bloat, the amount of energy savings depends on how much energy reduction the time–energy frontier yields for longer iteration times.
+Table 4 shows the amount of energy savings of a non-straggler pipeline given varying straggler slowdowns.
+For a given slowdown factor ($T^{\prime}/T$), a non-straggler pipeline’s iteration time is set to be $T_{\textrm{opt}}=\min(T^{*},T^{\prime})$3.1), and energy reduction comes from (1) slowing down the pipeline itself and (2) reducing the time (and energy) blocking on communication, waiting for the straggler.
+The percentage of savings initially increases with the straggler’s iteration time, but then gradually decreases as it slows down beyond $T^{*}$.
 This is expected.
-The absolute amount of energy reduction in Joules is the largest when the non-straggler pipeline’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$.
-Thus, as the straggler slows down beyond $T^{*}$, the time and energy consumed while waiting for the straggler increases, lowering the percentage of energy savings.
+The absolute amount of energy reduction in Joules is the largest when the straggler’s iteration time is $T^{*}$ and constant afterward, because Perseus does not slow down non-straggler pipelines beyond $T^{*}$3.1).
+Thus, as the straggler slows down beyond $T^{*}$, additional time and energy is consumed while waiting for the straggler, lowering the percentage of energy savings.
 Finally, the point of maximum energy savings is different for each model.
 This is because each model has a different $T^{*}$ value, which is determined by how much each stage’s computation slows down on the minimum-energy frequency.
-6.2.3. How Much Potential Savings Was Realized?
-The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the “iteration time–energy” Pareto frontier).
-For intrinsic bloat without stragglers, Perseus on average realizes 74% and 89% the opportunity on A100 and A40, respectively, with negligible slowdown.
+6.2.3. How Much Potential Saving Was Realized?
+The largest possible savings under our problem setting occurs when running every computation at their minimum-energy frequencies (i.e., the $T^{*}$ point on the time–energy frontier).
+For intrinsic bloat without stragglers, Perseus realizes on average 74% and 89% of the potential savings on A100 and A40, respectively, with negligible slowdown.
 This is possible because there are much more non-critical computations in the DAG that can be slowed down than critical ones.
@@
 Emulation Methodology.
-We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the “iteration time–energy” frontier.
+We profile the time and energy consumption of each layer (e.g., Transformer decoder) in GPT-3 175B and Bloom 176B and run our optimization algorithm to obtain the time–energy frontier.
 We perform strong scaling when varying the number of GPUs (Table 5) in order to keep the global batch size constant (largebatch1, ; largebatch2, ).
@@
 running all maximum frequencies given by our emulator and real experiments for our A100 workloads and find that the emulator always underestimates energy savings.
-Specifically, savings on the leftmost and rightmost point of the Pareto frontier are underestimated by 18.6% and 21.7% on average, respectively.
+Specifically, savings on the leftmost and rightmost point of the frontier are underestimated by 18.6% and 21.7% on average, respectively.
 We believe this is due to our simplifying assumption that $P_{\mathrm{blocking}}$ is constant regardless of the GPU’s frequency.
-This means savings given by the emulator can be considered a lower bound for actual savings.
+This means the savings given by the emulator can be considered a lower bound for actual savings.
 Figure 7. [Emulation] Energy savings breakdown with straggler slowdown 1.2 and 1,024 GPUs.
-Results Summary.
+Result Summary.
 Figure 7 breaks down the amount of energy bloat reduction for GPT-3 175B and Bloom 176B when slowdown degree is 1.2 on 1,024 GPUs.
-EnvPipe can only reduce intrinsic bloat as it does not provide an “iteration time–energy” frontier; even for intrinsic bloat, it is suboptimal.
+EnvPipe can only reduce intrinsic bloat as it does not provide a time–energy frontier; even for intrinsic bloat, it is suboptimal.
 In contrast, Perseus reduces energy consumption by up to 30% by reducing both intrinsic and extrinsic energy bloat.
@@
 Table 6 shows Perseus’s intrinsic energy bloat reduction without stragglers for GPT-3 175B and Bloom 176B.
-The number of microbatches was varied based on Table 5.
-For all models, as more and more microbatches are added to the pipeline, the amount of intrinsic bloat decreases.
-This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) vs.
-steady state phase (middle).
+The number of microbatches is varied based on Table 5.
+For all models, as more microbatches are added to the pipeline, the amount of intrinsic bloat decreases.
+This is fundamentally due to the ratio of microbatches in the 1F1B’s warm-up and flush phase (beginning and end) versus the steady state phase (middle).
 Most microbatches in the warm-up and flush phases can slow down until their minimum energy frequency, yielding large energy savings.
@@
 Please note the different Y-axes.
-Intrinsic + Extrinsic Bloat Reduction With Stragglers.
-We introduce stragglers in large-scale emulation in various strong scaling configurations (Table 5) and with varying straggler slowdown degree.
-Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus.
-The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
-An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline.
-It may seem intuitive to assume that more pipelines brings more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption.
-However, this holds only in weak scaling configuration, where per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally to the number of pipelines).
-Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes.
-With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ).
-These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy saving percentages.
-However, as the absolute number of GPUs increase, even a small savings percentage gives huge absolute energy savings.
-6.4. Iteration Time–Energy Frontier Comparison
 Figure 9.
@@
 Please note the different X- and Y-axes.
-The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the “iteration time–energy” frontier obtained using Perseus’s optimization algorithm (§4).
+Intrinsic + Extrinsic Bloat Reduction With Stragglers.
+We introduce stragglers of varying slowdowns in large-scale emulation.
+Figure 8 reports the amount of intrinsic + extrinsic energy bloat reduction achieved by Perseus.
+The trend where energy saving increases until $T^{\prime}<T^{*}$ and wanes afterward is consistent with Section 6.2.2.
+An interesting observation here is that there is a tradeoff between scale and energy savings: configurations with more pipelines have less percentage of energy savings or less amount of energy savings per pipeline.
+It may seem intuitive to assume that more pipelines bring more energy savings, as there is only one straggler pipeline and all the other pipelines can reduce their energy consumption.
+However, this holds only in weak scaling scenarios, where the per-pipeline batch size is held constant (i.e., increasing the global batch size proportionally with the number of pipelines).
+Instead, in the more realistic strong scaling configuration (i.e., the global batch size is constant and per-pipeline batch size is decreased as more pipelines deployed), each pipeline’s number of microbatches changes.
+With fewer microbatches, the ratio of pipeline bubble (time that GPUs are idle) at the beginning and end of each pipeline iteration increases (megatronlm-sc21, ).
+These bubbles cannot be perfectly eliminated by intrinsic energy bloat reduction, resulting in a smaller energy savings percentages.
+However, as the absolute number of GPUs increases, even a small savings percentage is expected to yield huge absolute energy savings.
+6.4. Iteration Time–Energy Frontier Comparison
+The energy bloat reductions in Sections 6.2 and 6.3 were made possible by the time–energy frontier obtained using Perseus’s optimization algorithm (§4).
 Here, we further examine the frontier with different parallelization configurations and models and compare against Zeus (zeus-nsdi23, ).
@@
 All results were obtained from running on testbed GPUs.
-Appendix G has frontiers for other models.
-Perseus Pareto-dominates Zeus.
-ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, unable to reduce intrinsic energy bloat.
+Appendix H has frontiers for other models.
+Perseus Pareto-dominates both baselines derived from Zeus.
+ZeusGlobal is unaware of pipeline stage imbalances and slows down every stage, and therefore is unable to intrinsic energy bloat.
 ZeusPerStage can balance the forward computation time of each stage, but is unaware of the critical path of the DAG, slowing down critical computations.
@@
 Profiling.
-In the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage.
+At the beginning of training, the client-side profiler (§5) profiles forward and backward computations in each stage.
 For our A100 workloads, the initial profiling phase increased end-to-end training time by 13 minutes on average, which is negligible overhead for large model training.
@@
 The average runtime of the optimization algorithm (§4) across our A100 workloads was 6.5 minutes, with the longest being Bloom 3B (15.7 minutes).
-For our largest scale emulation (GPT-3 175B on A100 with 8,192 GPUs), the algorithm ran for 87 seconds, which is shorter because it is sufficient to optimize just one data parallel copy (§4.4).
-While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time also increases with the scale of the training job.
-Looking up the optimal energy schedule given $T^{\prime}$ is instant.
-7. Related Works
+For our largest-scale emulation workload (GPT-3 175B on A100 with 8,192 GPUs), the algorithm took 87 seconds, which is short because it is sufficient to optimize just one data parallel pipeline (§4.4).
+While runtime will increase with larger DAGs for larger models, we believe the overhead is justified because training time is likely to also increase with the scale of the training job.
+Looking up the optimal energy schedule given the straggler’s iteration time $T^{\prime}$ is instantaneous.
+7. Related Work
 Large Model Training.
 Many recent works focus on enabling and accelerating large model training using 3D parallelism.
-GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) were the first to introduce pipeline parallelism.
-Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ) is a large model training framework that provides efficient 3D parallel execution plans tailored for Transformer models.
-3D parallelism is considered to be go-to solutions for modern large model development due to strong open-source implementations (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively simple implementation complexity.
-Finally, Alpa (alpa-osdi22, ) and GSPMD (spmd_socc22, ) provide automatic parallelization for general DNNs.
+GPipe (gpipe-neurips19, ) and PipeDream (pipedream-sosp19, ) first introduced pipeline parallelism.
+3D parallelism, especially popularized for Transformer-based models by Megatron-LM (megatronlm-arxiv, ; megatronlm-sc21, ), is considered to be the go-to solution for modern large model training due to strong open-source projects (megatronlm-github, ; deepspeed-github, ; gpt-neox-github, ) and relatively low implementation complexity.
+Extending this, Alpa (alpa-osdi22, ), GSPMD (spmd_socc22, ), and nnScaler (nnscaler-osdi24, ) provide automatic parallelization for general DNNs.
 However, energy consumption is not an optimization metric for any of the major large model training frameworks.
-DNN Training and Energy Consumption.
-A recent line of work has highlighted the enormous amount of energy consumption and carbon emission of DNN training (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ) and proposed optimization methods (odpp, ; gpoeo, ; zeus-nsdi23, ; chase-ccai23, ; depo, ; envpipe-atc23, ).
-Especially, Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but focuses on single-GPU training.
-EnvPipe (envpipe-atc23, ), on the other hand, aims to reduce the energy consumption of large model training with minimum slowdown.
-However, its heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings.
-Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm.
-Perseus is also the first to introduce the notion of energy bloat.
+Several works utilize computation idle times within large model training pipelines to insert additional useful work (bamboo-nsdi23, ; pipefisher-mlsys23, ; hydro-osdi23, ).
+In contrast, Perseus’s approach is to slow down preceding computations to fill the idle time and reduce power and energy consumption.
+The rationale is twofold.
+First, the amount of idle time–especially in the steady state of the pipeline–is typically short.
+Inserting extra work may slow down the entire pipeline or require pipeline re-partitioning to make it feasible.
+Second, GPUs, especially those in the earlier stage of the pipeline, are likely already constrained by memory capacity, leaving little headroom for extra computation.
+We note that even with extra computation inserted, some idle time is likely to remain, leaving room for Perseus to provide energy savings.
+Deep Learning and Energy Consumption.
+A line of work measures or estimates the large amount of energy consumption and carbon emission of deep learning workloads (patterson2021carbon, ; energy-nlp-policy, ; quantify-carbon, ; dodge-faact22, ; bloomcarbon-arxiv23, ).
+In terms of optimization, some works determine the GPU’s execution speed for a single fixed sequence of GPU computations (odpp-ccgrid20, ; datadriven-ccgrid20, ; gpoeo-tpds21, ; chase-ccai23, ; depo-future23, ; cofris-igsc23, ), falling short when the time–energy frontier of a complex large model computation DAG needs to be characterized.
+Zeus (zeus-nsdi23, ) is a recent work that observes the tradeoff between GPU computation time and energy consumption, but still focuses on single-GPU training.
+EnvPipe (envpipe-atc23, ), on the other hand, aims to find a point solution that reduces the energy consumption of large model training with minimum slowdown.
+However, EnvPipe’s heuristic assumes that the last pipeline stage is always the bottleneck, leading to suboptimal savings.
+In terms of optimizing a single training job, Perseus is a superset of both Zeus and EnvPipe, achieved by viewing large model training as a computation DAG and introducing a principled optimization algorithm.
+Big Data and Energy Consumption.
+There are existing works that change the frequency of CPUs to improve the energy efficiency of big data workloads, including those that only execute computation-intensive phases (e.g., map and reduce) with the maximum frequency (mapreduce-dvfs-igcc11, ; hadoop-dvfs-fgcs16, ), those that choose the lowest frequency that can meet a given deadline (hadoop-dvfs-js17, ; express-icac17, ), or those that choose frequencies that balance the completion time of parallel tasks (yet simply minimizing idle time for smaller tasks within one stage) (spark-dvfs-js21, ).
+Perseus provides a principled approach to setting the job’s end-to-end completion time (§3.1) and task execution speed (§4), and can be directly applied to big data workloads by generating a DAG of CPU computations and optimizing the CPU’s frequencies.
 8. Conclusion
 We presented Perseus, a software-based energy optimization system for large model training.
-Perseus builds on the observation that there are computation imbalances at different levels in large model training that causes intrinsic and extrinsic energy bloat, and introduces a principled graph cut-based algorithm that simultaneously reduces both.
-As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby reducing average power draw as well.
+Perseus builds on the observation that there are computation imbalances at different levels in large model training that cause intrinsic and extrinsic energy bloat, and simultaneously reduces both with a principled graph cut-based algorithm.
+As a result, Perseus advances the state-of-the-art of DNN training energy optimization by delivering energy savings with negligible slowdown, thereby also reducing average power draw and enhancing the sustainability of large model training.
+Acknowledgements
+We would like to thank the SOSP reviewers, our shepherd Eddie Kohler, and SymbioticLab members for their insightful feedback.
+This work is in part supported by NSF grants CNS-2104243, CNS-2106184, and CCF-2327011, NWO VICI grant 639.023.812, grants from VMware and the Mozilla Foundation, and a gift from Salesforce.
+We also thank Chameleon Cloud for providing A100 nodes as well as CloudLab.
+Jae-Won Chung is additionally supported by the Kwanjeong Educational Foundation.
 Appendix A Visualizations for Intrinsic Energy Bloat
@@
 Therefore, even with minimum imbalance partitioning, it is difficult to perfectly balance stages.
-B.2. Alternative Planning Methods
+B.2. Does Imbalance Decrease with Larger Models?
+Specifically for Transformer [77]-based models with homogeneous intermediate Transformer layers, the degree of imbalance will decrease if the number of pipeline stages is held constant.
+This is because the relative size of the embedding layer and language model head will decrease as intermediate Transformer layers increase in number and size.
+However, as the model grow larger, the number of pipeline stages will have to be increased simultaneously.
+This increases imbalance, because the relative amount of computation in one layer with respect to the amount in one pipeline stage becomes larger.
+As such, there is no simple relationship between model size and the amount of imbalance.
+B.3. Alternative Planning Methods
 3D parallelism is the go-to solution for large model training, and our minimum imbalance pipeline partitioning method is optimal for 3D parallelism because we implemented a brute force search.
@@
 Thus, in the vast majority of models, there will always be some degree of imbalance between stages due to the minimum granularity of computation time.
-B.3. Experiment Parameters
+B.4. Experiment Parameters
 Model # Parameters Global Batch Size Microbatch Size # Microbatches
@@
 Therefore, the function $f$ takes $O(n)$ time, which is polynomial time w.r.t the input size.
-Appendix D Full Details of GetNextPareto
+Appendix D Continuous Relaxation
+Figure 11. Pareto-optimal (time, energy) choices for forward and backward computations in each stage of GPT-3 0.3B on NVIDIA A40 GPUs.
+The Pipeline Energy Minimization problem is NP-Hard, as we have proved in Appendix C.
+Thus, we perform continuous relaxation for the problem by fitting an exponential function to each Pareto-optimal (time, energy) points for each forward and backward computation in each stage.
+Furthermore, the exponential function captures the nature of diminishing returns well, in that the amount of energy consumption needed to reduce computation time by a constant amount increases multiplicatively.
+Figure 11 visualizes Pareto-optimal (time, energy) measurements for GPT-3 0.3B on A40 GPUs, showing that the exponential function is a natural fit for data.
+This was our consistent observation across different GPUs and models in our evaluation.
+Appendix E Full Details of GetNextSchedule
 For the sake of presentation, we made a simplifying statement in the body of the paper that computations only speed up by $\tau$.
@@
 In the following, we describe the procedure of annotating edges with flow capacities and then solving the problem with maximum flow.
-D.1. Generating Capacity DAG
+E.1. Generating Capacity DAG
 On the computation DAG, we first remove all the computations that are not on any of the critical paths and construct the Critical DAG.
@@
 (8) $(l_{i},u_{i})=\begin{cases}(0,e_{i}^{+})&\textrm{if }t_{i}\textrm{ is longest possible (slowest)}\\ (e_{i}^{-},\infty)&\textrm{if }t_{i}\textrm{ is shortest possible (fastest)}\\ (e_{i}^{-},e_{i}^{+})&\textrm{otherwise.}\end{cases}$
-This equivalence was discovered by Phillips and Dessouky [53, 27].
+This equivalence was given by the theoretical works of Phillips and Dessouky [64, 31].
 Thus, we construct the Capacity DAG from the Critical DAG by annotating its edges with flow capacities given by Equation 8.
@@
 Finally, after the minimum cut has been identified from the Capacity DAG, edges in $I^{+}$ are sped up by $\tau$ and those in $I^{-}$ are slowed down by $\tau$, ultimately reducing the length of every critical path exactly by $\tau$ with the smallest possible energy increase.
-D.2. Max Flow Algorithm on the Capacity DAG
+E.2. Max Flow Algorithm on the Capacity DAG
 A characteristic of our Capacity DAG that precludes the direct application of well-known max flow algorithms is that some edges also have flow lower bounds, asserting that at least a certain amount of flow must pass through in the edge.
@@
 Algorithm 3 Maximum Flow with Lower Bounds
-Appendix E Proof for Polynomial Runtime
-Perseus enumerates the entire “iteration time–energy” Pareto frontier one by one, and the runtime of one iteration is polynomial time with respect to the number of stages $N$ and the number of microbatches $M$.
+Appendix F Proof for Polynomial Runtime
+Perseus enumerates the entire time–energy frontier one by one, and the runtime of one iteration is polynomial time with respect to the number of stages $N$ and the number of microbatches $M$.
 Thus, determining whether the entire algorithm runs in polynomial time reduces to whether the worst case number of iterations is polynomial with respect to $N$ and $M$.
-While for general DAGs the maximum number of points on the Pareto frontier can be exponential with respect to the size of the DAG [59], here we prove that under mild assumptions for DAGs that represent pipeline schedules, the number of iterations is $O(N+M)$.
+While for general DAGs the maximum number of points on the frontier can be exponential with respect to the size of the DAG [71], here we prove that under mild assumptions for DAGs that represent pipeline schedules, the number of iterations is $O(N+M)$.
 The assumptions are valid for all pipeline schedules known to the authors, including GPipe [34] and 1F1B [56].
-Theorem E.1.
+Theorem F.1.
 For DAGs that represent pipeline schedules, the number of iterations needed is $O(N+M)$.
@@
 for instuction in pipeline_schedule: if isinstance(instruction, Forward): controller.set_speed(”forward”) profiler.begin(”forward”) # Run forward on microbatch profiler.end(”forward”) elif isinstance(instruction, Backward): controller.set_speed(”backward”) profiler.begin(”backward”) # Run backward on microbatch profiler.end(”backward”) … Perseus client API integration example.
-Appendix F Perseus Client Integration Example
-Listing E shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
+Appendix G Perseus Client Integration Example
+Listing F shows an example integration of Perseus’s Client API (first three rows in Table 2) with a hypothetical (but typical) training framework’s pipeline execution engine.
 A typical structure of a pipeline execution engine is to have instructions for each distinct operations in the pipeline, including not only forward and backward executions, but also P2P and collective communications, and implement a handler for each instruction.
 Therefore, framework developers can wrap such handlers with the Perseus client APIs to mark their beginning and end.
-Appendix G Iteration Time–Energy Frontiers
+Appendix H Time–Energy Frontiers
 (a) BERT
@@
 Figure 12. Eight stage pipeline parallelism on A40.
-Figure 11 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
+Figure 12 shows the time–energy frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads ran with eight stage pipeline parallelism, measured in NVIDIA A40 GPUs.
 T5 shows an interesting frontier due to the hardware topology of our A40 machine setup: Each node has four GPUs and NVLink connects GPUs 0 and 1, and 2 and 3; GPUs 1 and 2 must communicate through the NUMA interconnect; Finally, different nodes are connected with Infiniband only adjacent to GPUs 0 and 1 (data to and from GPUs 2 and 3 must also go through the NUMA interconnect).
@@
 Figure 13. 4 stage pipeline parallelism on A100 PCIe.
-Figure 12 shows the “iteration time–energy” frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs.
+Figure 13 shows the time–energy frontiers achieved by Perseus and the two baseline approaches for the rest of the workloads, measured with four stage pipeline parallelism in NVIDIA A100 PCIe GPUs.
 Wide-ResNet has a noisy frontier because the variability in microbatch loading time introduces noise in the end-to-end iteration time when computations are tightly packed by Perseus.